### v1

# Security Weaknesses of Copilot Generated Code in GitHub

Abstract.
Modern code generation tools use AI models, particularly Large Language Models (LLMs), to generate functional and complete code. While such tools are becoming popular and widely available for developers, using these tools is often accompanied by security challenges, leading to insecure code merging into the code base. Therefore, it is important to assess the quality of the generated code, especially in terms of its security. Researchers have recently explored various aspects of code generation tools, including security. However, many open questions about the security of the generated code require further investigation, especially the security issues of automatically generated code in the wild. To this end, we conducted an empirical study by analyzing the security weaknesses in code snippets generated by GitHub Copilot that are found as part of publicly available projects hosted on GitHub. The goal is to investigate the types of security issues and their scale in real-world scenarios (rather than crafted scenarios). To this end, we identified 435 code snippets generated by GitHub Copilot from publicly available projects. We then conducted extensive security analysis to identify Common Weakness Enumeration (CWE) instances in these code snippets. The results show that (1) 35.8% of Copilot generated code snippets contain CWEs, and those issues are spread across multiple languages, (2) the security weaknesses are diverse and related to 42 different CWEs, in which CWE-78: OS Command Injection, CWE-330: Use of Insufficiently Random Values, and CWE-703: Improper Check or Handling of Exceptional Conditions occurred the most frequently, and (3) among the 42 CWEs identified, 11 of those belong to the currently recognized 2022 CWE Top-25. Our findings confirm that developers should be careful when adding code generated by Copilot (and similar AI code generation tools) and should also run appropriate security checks as they accept the suggested code. It also shows that practitioners should cultivate corresponding security awareness and skills.
Code Generation, Security Weaknesses, CWEs, GitHub Copilot
$c$s: Software and its engineering Software development techniquesccs: Security and privacy Software security engineering
1. Introduction
Code generation tools aim to automatically generate functional code based on prompts, which can include text descriptions (comments), code (such as function signatures, expressions, variable names, etc.), or a combination of text and code (Sarkar et al., 2022). After writing an initial code or comment, developers can rely on code generation tools to complete the remaining code. This approach can save development time and accelerate the software development process. Automated code generation tools have always been an active research discussion topic (Siddiq and Santos, 2022; Li et al., 2022b). Some of the earliest work can be traced back to the 1960s when Waldinger and Lee proposed a program synthesizer called PROW, which automatically generated LISP programs based on specifications provided by users in the form of predicate calculus (Waldinger and Lee, 1969). As computer languages continued to evolve, more and more programming languages began to support meta-programming, making automated code generation technology more efficient and flexible. In recent years, the rapid development of artificial intelligence technologies, particularly in the form of machine learning and deep learning models, has accelerated the development of code generation technologies.
Recent advancements in code generation came with the emergence of Large Language Models (LLMs). LLMs are deep learning models trained on a large code/text corpus with powerful language understanding capabilities that can be used for tasks such as natural language generation, text classification, and question-answering systems (Chen et al., 2021). Compared to previous deep learning methods, the latest developments in LLMs, such as Generative Pre-trained Transformer (GPT) models, have opened up new opportunities to address the limitations of existing automated code generation technology (Mozannar et al., 2022). Currently, code generation tools based on LLMs have also been widely applied, such as Codex by OpenAI (OpenAI, [n. d.]), AlphaCode by DeepMind (Li et al., 2022a), and CodeWhisperer by Amazon (Becker et al., 2022).
These models are trained on billions of public open-source lines of code, which includes public code with unsafe coding patterns (He and Vechev, 2023). Therefore, code generation tools based on such models can pose security risks, and the code they generate may also have security weaknesses. For example, GitHub Copilot may produce some insecure code, as its underlying model Codex is pre-trained on untrusted data from GitHub (Brown et al., 2020), which is known to contain buggy programs (Rokon et al., 2020). In addition, the code with vulnerabilities generated by these code generation tools may continue to be used to train the model, thus further generating code with vulnerabilities, leading to a vicious cycle. Previous research has studied code generation tools, with more focus on the correctness of the results (Lertbanjongngam et al., 2022; Wong et al., 2022; Dakhel et al., 2022; Pudari and Ernst, 2023), and relatively less attention has been paid to security aspects (Pearce et al., 2022; Siddiq et al., 2022; Perry et al., 2022). To the best of our knowledge, potential security weaknesses in practical scenarios have not been fully considered and addressed in previous work, and GitHub Copilot clarifies that “the users of Copilot are responsible for ensuring the security and quality of their code (GitHub, [n. d.])”. GitHub also provides tools such as CodeQL to help developers scan the security issues in their code.
To this end, we conducted an empirical study on the security weaknesses of generated code by GitHub Copilot, which is available on GitHub. We chose Copilot as our research subject because it is a commercial instance of AI-assisted programming and has gained much attention and popularity among developers since its launch in 2021. The security weaknesses of code generated by Copilot have also gained attention in the research and practice community. Furthermore, thousands of developers in the GitHub community have shared their experiences of using Copilot in real-world systems (Dakhel et al., 2022). We collected code generated by Copilot that has been used in projects on GitHub and analyzed the security of the generated code through the lens of a real-world production environment. Then, we used static analysis tools to perform security analysis on the collected code snippets and classified the security weaknesses in the code snippets using the Common Weakness Enumeration (CWE).
Our findings show that: (1) 35.8% of Copilot generated code snippets have security weaknesses, and security weaknesses arise regardless of the programming language used; (2) the security weaknesses are diverse and related to 42 different CWEs, in whichCWE-78: OS Command Injection, CWE-330: Use of Insufficiently Random Values and CWE-703: Improper Check or Handling of Exceptional Conditions are the most frequently occurred; and (3) among the 42 CWEs identified, 11 CWEs belong to the currently recognized 2022 CWE Top-25.
The contributions of this work: (1) We curated a dataset of code snippets generated by Copilot that has been used in projects on GitHub (a curated data (Fu et al., 2023) is made available online for future research in this area.) and conducted security checks on them, which can to some extent reflect the frequency of security weaknesses encountered by developers when using Copilot to generate code in actual coding; (2) We extensively checked all possible CWEs in the code snippets and analyzed them. This can help developers understand the common CWEs caused by using Copilot to generate code in actual coding and how to accept the code suggestions provided by Copilot safely.
The rest of this paper is structured as follows: Section 2 presents the related work. Section 3 presents the research questions and the research design of this study. Section 4 presents our study results, which are further discussed in Section 5. The potential threats to validity are clarified in Section 6. Section 7 concludes this work with future work directions.
2. Related Work
2.1. AI-assisted Code Generation Tools
With the rise of code generation tools integrated with IDEs, many studies have evaluated these code generation systems based on transformer models to better understand their effectiveness in real-world scenarios. Previous research mainly focused on whether the code generated by these tools can meet users’ functional requirements. Yetistiren et al. (Yetistiren et al., 2022) evaluated the effectiveness, correctness, and efficiency of the code generated by GitHub Copilot, and the results showed that GitHub Copilot could generate valid code with a success rate of 91.5%, making it a promising tool. Sobania et al. (Sobania et al., 2022) evaluated the correctness of the code generated by GitHub Copilot and compared the tool with an automatic program generator with a Genetic Programming (GP) architecture. They concluded there was no significant difference between the two methods on benchmark problems. Nguyen and Nadi (Nguyen and Nadi, 2022) conducted an empirical study using 33 LeetCode problems and created queries for Copilot in four different programming languages. They evaluated the correctness and comprehensibility of the code suggested by Copilot by running tests provided by LeetCode. They found that Copilot’s suggestions have lower complexity. Burak et al. (Yetiştiren et al., 2023) evaluated the code quality of AI-assisted code generation tools (GitHub Copilot, Amazon CodeWhisperer, and ChatGPT). They compared the improvements between the latest and older versions of Copilot and CodeWhisperer and found that the quality of generated code had improved.
In recent years, researchers have also started to focus on the experience of developers when using AI-assisted code generation tools and how the tools can improve productivity by observing their behavior. Vaithilingam et al. (Vaithilingam et al., 2022) studied how programmers use and perceive Copilot, and they found that while Copilot may not necessarily improve task completion time or success rate, it often provides a useful starting point. They also noted that participants faced difficulties in understanding, editing, and debugging the code snippets generated by Copilot. Barke et al. (Barke et al., 2022) presented the first theoretical analysis of how programmers interact with Copilot based on the observations of 20 participants. Sila et al. (Lertbanjongngam et al., 2022) conducted an empirical study on AlphaCode, identifying similarities and performance differences between code generated by code generation tools and code written by human developers. They argued that software developers should check the generated code for potentially problematic code that could introduce performance weaknesses.
These works above conducted relatively extensive evaluations of code generation tools in terms of correctness, effectiveness, and robustness. However, there is still room for improvement regarding its security, as detailed in the following section.
2.2. Security of Code Generation Techniques and LLMs
Code security is an issue that cannot be ignored in the software development process. Recent work has primarily focused on evaluating the security of the code generation tools and the security of the LLMs that these tools are based on.
Pearce et al. (Pearce et al., 2022) first evaluated the security of GitHub Copilot in generating programs by identifying known weaknesses in the suggested code. The authors prompted Copilot to generate code for 89 cybersecurity scenarios and evaluated the weaknesses in the generated code. They found that 40% of the suggestions in the relevant context contained security-related bugs (i.e., CWE classification from MITRE (The MITRE Corporation, [n. d.])). Siddiq et al. (Siddiq et al., 2022) conducted a large-scale empirical study on code smells in the training set of a transformer-based Python code generation model and investigated the impact of these harmful patterns on the generated code. They observed that Copilot introduces 18 code smells, including non-standard coding patterns and two security smells (i.e., code patterns that often lead to security defects). Khoury et al. (Khoury et al., 2023) studied the security of the source code generated by the ChatGPT chatbot based on LLMs, and they found that ChatGPT was aware of potential weaknesses but still frequently generated some non-robust code.
Several researchers also compared the situation where code generation tools produce insecure code with that of human developers. Sandoval et al. (Sandoval et al., 2022) conducted a security-driven user study, and their results showed that the rate at which AI-assisted user programming produced critical security errors was no more than 10% of the control group, indicating that the use of LLMs does not introduce new security risks. Asare et al. (Asare et al., 2023) conducted a comparative empirical analysis of these tools and language models from a security perspective and investigated whether Copilot is as bad as humans in generating insecure code. They found that while Copilot performs differently across vulnerability types, it is not as bad as human developers when it comes to introducing vulnerabilities in code. In addition, researchers have also constructed datasets to test the security of these tools. Tony et al. (Tony et al., 2023) proposed LLMSecEval, a dataset containing 150 natural language prompts that can be used to evaluate the security performance of LLMs. Siddiq et al. (Siddiq and Santos, 2022) provided a dataset, SecurityEval, for testing whether a code generation model has weaknesses. The dataset contains 130 Python code samples.
Unlike the works above, we studied the security weaknesses exhibited by code generation tools in a real-world production environment (i.e., GitHub). We collected code snippets from GitHub generated by developers using Copilot in daily production as a source of research data, whereas in the Pearce et al. (Pearce et al., 2022) study, the research data came from code generated by the authors using Copilot based on the natural language prompts related to high-risk network security weaknesses. In addition to this, Pearce et al. configured CodeQL only to examine CWEs targeted by security weaknesses associated with the prompted scenarios. In contrast, we used various static analysis tools to examine all types of CWEs and analyze them extensively. Our research results may help developers understand what common CWEs are prone to result from using Copilot to generate code in real coding.
2.3. Security Static Analysis
Vulnerabilities detection is critical to improve software security and ensure quality  (Iannone et al., 2022). There are two used methods for vulnerability detection in source code: via static and dynamic code analysis. Dynamic analysis techniques are more sound and precise but lack coverage (Ernst, 2003). On the other hand, static analysis is less precise but offers greater coverage and allows to analyze programs without the need to execute them (Sui et al., 2020). Static analysis has been widely used to find security issues in code, given it is cheaper to run and can conduct whole program analyses without the need to execute the program (Dunlap et al., 2023). OWASP (OWASP, [n. d.]) provides a list of commonly used static analysis tools. This includes tools like CodeQL: a general-purpose automatic scanning tool, FindBugs: a tool for Java programs, ESLint: a tool for JavaScript programs, Bandit: a tool for Python programs, and GoSec: a tool for Go programs. Such tools have been widely used in previous security analysis research  (Tómasdóttir et al., 2018; Pearce et al., 2022; Siddiq et al., 2022).
Kaur et al. (Kaur and Nayyar, 2020) compared static analysis tools for vulnerability detection in scanning C/C++ and Java source code. Tomasdottir et al. (Tómasdóttir et al., 2018) conducted an empirical study on ESLint, the most commonly used JavaScript static analysis tool among developers. Pearce et al. (Pearce et al., 2022) used CodeQL for security weakness scanning of generated Python and C++ code. Siddiq et al. (Siddiq and Santos, 2022) used Bandit to check Python code generated using a test dataset.
These static analysis tools support different analysis algorithms and techniques. By using multiple tools for analysis, potential weaknesses in the code can be discovered from different perspectives and levels, avoiding omissions and improving the accuracy of the analysis. Our study first used CodeQL to scan the collected code snippets. CodeQL is an open-source tool that supports multiple languages, including Java, JavaScript, C++, C#, and Python. It can find weaknesses in a codebase based on a set of known weaknesses/rules. In addition, to obtain more comprehensive scan results, we supplemented the scan of code in different languages with static analysis tools (i.e., Cppcheck and Bandi) tailored to specific languages.
3. Research Design
In this section, we describe our research design in detail. In Section 3.1, we first define our Research Questions (RQs), followed by the process of collecting and filtering the code snippets generated by Copilot in Section 3.2. We then explain the security analysis performed on the identified snippets and the process of filtering the raw results generated by static analysis tools in Section 3.3.
3.1. Research Goal and Questions
This empirical study aims to understand the potential security weaknesses in Copilot generated code. We first collected code snippets generated by Copilot from GitHub projects as the data source for our research. It should be noted that it is not possible to access all the code generated by Copilot in GitHub projects, as there is no direct way to identify if part of a file was generated by Copilot (i.e., source files do not contain any signatures to indicate if Copilot generates the code). However, we can identify many code snippets by searching through the repository description and the comments provided in the code (details provided in Section 3.2.2). We then analyzed the identified snippets for security weaknesses. We aim to help developers and researchers understand common security weaknesses when using Copilot without focusing on whether the code generation aspects are used correctly.
We conducted this empirical study following the guidelines proposed in (Easterbrook et al., 2008). The RQs, their rationale, and the research process of this study (see Fig. 1) are detailed in the subsections below.
Figure 1. Overview of the research process
RQ1. How secure is the code generated by Copilot in GitHub Projects?
Rationale: Copilot may produce code suggestions that developers accept but these suggestions may include security weaknesses that could potentially make the program vulnerable. The answer to RQ1 helps to understand the frequency of security weaknesses developers encounter when using Copilot in production.
RQ2. What security weaknesses are present in the code snippets generated by Copilot?
Rationale: Copilot generated code may contain security weaknesses (Pearce et al., 2022), and developers should conduct a rigorous security review before accepting the code generated by Copilot. As clarified in the documentation of GitHub Copilot “the users of Copilot are responsible for ensuring the security and quality of their code (GitHub, [n. d.])”. The answer to RQ2 can help developers better understand possible security weaknesses in the code generated by Copilot, thereby enabling them to prevent and fix these weaknesses more effectively.
RQ3. How many security weaknesses belong to the MITRE CWE Top-25?
Rationale: The MITRE list contains the top 25 most dangerous security weaknesses. The answer to RQ3 can help developers understand whether the code generated by Copilot contains widely recognized types of security weaknesses and Copilot’s ability to handle these recent and common weaknesses.
3.2. Data Collection and Filtering
We chose GitHub as the primary data source for answering our RQs. GitHub is widely used by developers. As the world’s largest code hosting platform, GitHub contains millions of public code repositories and offers access to a large number of code resources, allowing us to cover multiple programming languages and project types in our study. We used code snippets generated by Copilot on GitHub as our research object to analyze the relevant security weaknesses of Copilot. Our scripts and dataset are provided online in our replication package (Fu et al., 2023).
3.2.1. Code Snippets Collection
Step 1. To collect code snippets generated by Copilot, we first conducted a pilot search to formulate our search keywords. First, we used “GitHub Copilot” and “Copilot” as our search keywords. As expected, we found that the term “Copilot” not only refers to the code generation tool launched by GitHub but also to tools in the aviation or telemetry fields. Therefore, using the keyword “Copilot” solely may return irrelevant content that may not be related to the use of the tool. On the other hand, using “GitHub Copilot” as a search keyword can exclude content unrelated to the code generation tool Copilot and narrow the search scope, which is what we have used to locate the code snippets.
However, even with this basic search keyword, we still need to carefully filter the search results to ensure they are truly related to GitHub Copilot. Although using “GitHub Copilot” as a search keyword increases the relevance of the results to Copilot, these results are not necessarily the code snippets generated by Copilot. It should be noted that many code snippets containing the “GitHub Copilot” keyword in the search results display GitHub Copilot as text. Developers may use them to describe their experience using Copilot to generate code or showcase information related to Copilot. These code snippets are not what we need because they do not directly relate to the code generated by Copilot. Our target is code generated by Copilot, not code snippets containing the keyword “Copilot”.
Our observations from the pilot search showed that using keywords such as “by GitHub Copilot”, “use GitHub Copilot”, and “with GitHub Copilot” can improve the accuracy of the search results. These keywords enable us to focus more on the code generated using Copilot rather than code snippets that contain other content related to Copilot. In addition, since our goal is to use automated analysis tools to perform security scans on the collected code snippets, we further limited the types of code snippets during the search to Python, JavaScript, Java, C++, C#, and Go. These are the mainstream languages supported by Copilot, and also the languages supported by CodeQL. We collected the Code parts from these search results. Considering that some projects declare using GitHub Copilot generated code in their README files or project description provided in GitHub, we decided to retain the results from the Repository label in the search results. Fig. 2 shows an example of our search process.
Table 1 reports the search terms we used and the number of search results obtained from GitHub. In this step, we collected a total of 8,004 results, of which 7,749 were from the Code label, and 255 were from the Repository label. The same search result may contain multiple keywords, meaning there are duplicate projects in the collected data. After removing duplicate projects, we obtained a total of 4249 search results, of which 4081 were from the Code label, and 168 were from the Repository label. Table 2 shows the number of different language types of search results we obtained from the Code label.
Figure 2. Example of the search process
Table 1. Search results based on different terms used
# Search Term # Code # Repositories
1
“By GitHub Copilot”
2549
54
2
“Use GitHub Copilot”
1822
77
3
“With GitHub Copilot”
3378
127
Total 7749 255
Table 2. Search Results from GitHub
# Language # Code results
L1
Python
784
L2
JavaScript
863
L3
Java
641
L4
C++
437
L5
C#
386
L6
Go
970
Total 4081
3.2.2. Filtering Code Snippets
Step 2. After obtaining the results from the keyword searches, we further filtered them by not only considering the accuracy of the keywords but also investigating the project’s documentation, code comments, and other metadata in the search results to determine whether they were generated by GitHub Copilot. Additionally, since we wanted to obtain code snippets used in real-world projects, we excluded search results used to solve simple algorithmic problems on platforms, such as LeetCode, which generally involve simple code and may not involve security weaknesses.
We begin by explaining the terminology used in data filtering: the search results under the Repository label are the projects that contain code files, and the search results under the Code label are individual code files. Those code files contain Copilot generated code snippets. In filtering the projects, we followed three rules: (1) for search results under the Repository label, we identified projects that are fully generated by Copilot, as declared in the projects description or the associated README file(s). We retained code files for Python, JavaScript, Java, C++, C#, and Go, which are the main languages supported by Copilot. (2) For search results under the Code label, we retained code files with comments showing the code generated by Copilot. (3) As we mentioned above, we then excluded code files used to solve simple algorithm problems. We provide examples for the three rules in Figs.  3,  4 and  5. As shown in the example in Fig. 3 for the Repository label, we kept all the Python files. In the next example in Fig. 4, we kept the entire file where the Copilot generated code snippet was located. In Fig. 5, the code snippet was removed as it was determined the code just solved a simple algorithmic problem. Meanwhile, for the code files retained under the Repository label, we consider the entire code file as code generated by Copilot. In other words, we assume that all code in the file is generated by Copilot because it was stated in the README file that it was all generated by Copilot. For code files retained under the Code label, we know that the files contain code snippets, perhaps even just a few lines of code, generated by Copilot. Instead of identifying the specific Copilot generated code in this step, we combine the warning messages from the security scan and the code comments in the file to determine whether Copilot generates the code snippet with the security problem (this process is explained further in Section 3.3.2).
After completing the pilot data labeling, the first author checked the rest of the search results, and obtained a total of 465 code snippets. After removing duplicate results, we finally obtained 435 different code snippets. Among them, 249 are from the Repository label, and 186 are from the Code label. Table 3 shows the types and numbers of code snippets obtained.
Figure 3. Example of rule 1: project is fully written by Copilot
Figure 4. Example of rule 2: files with comments showing the code generated by Copilot
Figure 5. Example of rule 3: files used to solve simple algorithm problems
Table 3. Code snippets from GitHub
# Language # Code Snippets: Repository # Code Snippets: Code Total
L1 Python 132 119 251
L2 JavaScript 51 28 79
L3 Java 25 18 43
L4 C++ 14 12 26
L6 Go 19 1 20
L5 C# 8 8 16
Total 249 186 435
3.3. Data Pre-processing and Analysis
3.3.1. Data Pre-process
Step 3. CodeQL is a scalable static security analysis tool that is widely used in practice, and enables users to analyze code and detect relevant weaknesses using predefined queries and test suites and supports for multiple languages (including Java, JavaScript, C++, C#, Go, and Python (GitHub, 2021a)). Before using CodeQL to scan the identified code snippets for security weaknesses, we needed to create a CodeQL database for the source code. For interpreted languages like Python and JavaScript, the source code can be directly analyzed, while for compiled languages such as Java, the source code will need to be compiled first and then imported into the CodeQL database. Therefore, we first compiled the code snippets of all compiled languages (i.e., C#, Java, C++, and Go). We removed any code snippets that could not be compiled. For successfully compiled files, we generated the CodeQL database required for queries. At the same time, for interpreted languages Python and JavaScript files, we stored 20 files in each database to improve efficiency, because if we generate a database for an exceptionally large number of files, this would increase the database compilation and scanning time, which is much longer than partitioning them into small databases. In total, we obtained 80 code databases available for CodeQL scanning. Table 4 shows the types and numbers of files in each database.
Table 4. Databases for CodeQL scanning
# Language # Databases: Repository # Databases: Code
L1 Python 7 6
L2 JavaScript 3 2
L3 Java 8 18
L4 C++ 5 12
L5 C# 3 7
L6 Go 13 1
Total 39 46
3.3.2. Data Analysis
Step 4. We used well-known automated static analysis tools listed by OWASP (OWASP, [n. d.]) to scan the collected code snippets. Since different static analysis tools may use different algorithms and rules to detect security weaknesses, using multiple tools can increase our chances of discovering security issues in the code. To improve the coverage and accuracy of the results, we used two static analysis tools for security checks on each code snippet (i.e., CodeQL plus a dedicated tool for the specific language).
We first used CodeQL to analyze the code in our dataset. The default query suite for the standard CodeQL query package is codeql-suites/<lang>-code-scanning.qls. There are several useful query suites in the codeql-suite directory of each package. For example, the codeql/cpp-queries package contains the following query suites (GitHub, 2021b):
  • Cpp-code-scanning.qls, which is the standard code scanning query for C++. It covers various features and syntax of C++ and aims to discover some common weaknesses in the code.
  • Cpp-security-extended.qls, which includes some more advanced queries than cpp-code-scanning.qls and can detect more security weaknesses.
  • Cpp-security-and-quality.qls, which combines queries related to security and quality, covering various aspects of C++ development from basic code structure and naming conventions to advanced security and performance weaknesses. It aims to help developers improve the security and quality of their code.
In this study, we scanned code snippets using the <language>-security-and-quality.qls test suite related to security weaknesses. These test suites check for multiple security properties and cover many CWEs. For example, the python-security-and-quality.qls test suite for Python provides 168 security checks, the JavaScript test suite provides 203 security checks, and the C++ test suite provides 163 security checks. As the query reports only provide the name and description of the security issues, we manually matched the results in the query reports with the corresponding CWE IDs.
We then selected other popular static security analysis tools for files in each program languages we analyzed. We used the following popular security analysis tools: Bandit for Python, ESLint for JavaScript, Cppcheck for C ++, Findbugs for Java, Roslyn for C# and Gosec for Go. In cases where we could not directly obtain the CWE ID related to the security issue from the scan results, we manually mapped the security attributes to the corresponding CWE for later analysis. We explain the specific correspondences in detail in Section 4.2.
Step 5. We scanned code snippets from the Repository and Code labels, and we filtered the scan results before analyzing them. We first removed the scan results that were repeatedly prompted by two of the tools, then removed the results that were unrelated to the security issue, and finally confirmed that the Copilot generated code indeed caused the results related to the security issue. As we explained in Section 3.2.2, we considered the code snippet from the Repository label to be the entire code file. Therefore, we kept the entire scan results from the Repository label and counted all the security issues they suggested. For the code snippet obtained from the Code label, we started by scanning the code file. If the static analysis tool found a security issue in the code, we located the code snippet in the file according to the line number of the security issue indicated by the scanning result. We determined whether it was generated by Copilot based on the comments before and after the code snippet. If Copilot indeed generated the code snippet with a security issue, we kept the scan result for our subsequent statistics. We further analyzed the filtered scan results in Step 6, detailed in Section 4 according to the specific RQs. We provide our full dataset (including code snippets, full scan results, and filtered results) in our replication package (Fu et al., 2023).
4. Results
We present the results of three RQs formulated in Section 3.1 below. For each RQ, we first explain how we analyzed the collected code snippets to answer the RQ. We then provide a detailed presentation of the final results for each RQ.
4.1. RQ1: How secure is the code generated by Copilot?
Approach. To answer this RQ, we collected 435 code snippets generated by Copilot from GitHub projects. These snippets cover six common programming languages. We used two static analysis tools (CodeQL + another language-dedicated tool) to scan and analyze the code snippets and then combine the results obtained from the two tools. The aim is to achieve a better coverage of security issues. Therefore, as long as one of the tools detected the presence of a security issue, the code snippet was considered vulnerable.
In the analysis results obtained from the CodeQL tool, three types of warnings were used to describe the detected weaknesses: Recommendation, which provides suggestions for improving code quality; Warning, which alerts to potential weaknesses that could cause code to run abnormally or unsafely; and Error, which is the highest level of warning and alert to inform that the error could cause code to fail to compile or run incorrectly. Since our research primarily focused on security weaknesses, we only counted code snippets that had warnings and errors, and we ignored the other code quality recommendations. For the scanning results from the Code label, we also needed to identify whether the security issues obtained from the scan were from Copilot generated code snippets based on the comment that appears before the method. We provide an example of the scan results filtration in Fig. 6. In Step 1, we first went to the corresponding file to locate the specific code snippet based on the start and end lines of the scan results that suggested a security issue. In Step 2, we located the code at Line 103 and found no comment indicating that Copilot generated it. In Step 3, we also found that the code snippet generated by Copilot in the file was located at Lines 226 to 239, and we then determined that the code snippet generated by Copilot did not cause this security issue and discarded this scan result from further analysis. Finally, we aggregated the filtered results obtained using multiple analysis tools to calculate the number of code snippets with security issues detected.
Figure 6. Example of filtering scan results from the Code label that are generated by Copilot
Results. Table 5 shows the numbers of code snippets for different types and the numbers and percentages of code snippets with security weaknesses. From the statistical results, we found that out of the 435 code snippets generated by Copilot, 35.8% of them have security weaknesses, regardless of the programming language. There is a higher proportion of security weaknesses in Python and JavaScript code, which are the most popular Copilot languages developers use (Stackscale, 2021). Out of the 251 Python code snippets, 39.4% have security weaknesses. Among the 79 JavaScript code snippets we collected, 27.8% have security weaknesses. Among all programming languages, C++ code snippets have the highest proportion of security weaknesses, reaching 46.1%. Go also has a relatively high proportion of security weaknesses, at 45.0%. In comparison, the proportion of files with security issues is lower for C# and Java code, at 25% and 23.2%, respectively.
Table 5. The number and percentage of code snippets with security weaknesses generated by Copilot
Language # Snippets
# Snippets containing
security weaknesses
Python 251 99 39.4%
JavaScript 79 22 27.8%
Java 43 10 23.2%
C++ 26 12 46.1%
Go 20 9 45.0%
C# 16 4 25%
Total 435 156 35.8%
4.2. RQ2: What security weaknesses are present in the code snippets generated by Copilot?
Approach. To answer RQ2, we processed the results of the scans conducted for RQ1, eliminating duplicate security issues detected at the same code snippet location. In total, we identified 600 security weaknesses across 435 code snippets. Table 6 shows the number of security weaknesses found in code files of different programming languages.
Table 6. The number of security weaknesses in code snippets generated by Copilot
Language
# Snippets containing
security weaknesses
# Total security
weaknesses
Python 99 352
JavaScript 22 93
Java 10 56
C++ 12 70
Go 9 18
C# 4 11
Total 156 600
For each code snippet, we used CWEs to classify the security issues identified by the static analysis. Each CWE has a unique ID and a set of related descriptions, including its potential impact and how to detect and fix the CWE (The MITRE Corporation, [n. d.]). Some static analysis tools we used, such as Bandit and Gosec, provide a CWE ID corresponding to the detected security issues in their scan results. For other scan tools that do not directly give a CWE ID, such as Codeql, ESLint, and FindBugs, we manually associated the provided security issue information with a CWE ID, which is detailed below. The QL queries used in CodeQL often point to a specific CWE, and the scanning process typically displays the CWE-ID associated with the QL query. Although only the name of the QL query is shown in the scan results, we can manually correlate the name field with the CWE ID. For example, the result shows “Hard-coded credential in API call”, and we know that Hardcodedcredential.ql query belongs to CWE-798, this security issue can be mapped to the same CWE (798). In addition, while FindBugs and ESLint also identify security issues, their scan results only provide descriptions of the security issues. We manually associated the descriptions of the security issues in the results with relevant CWE descriptions to determine the specific CWE category to which these security issues belong. Initially, two authors independently matched each description of the security issue with a CWE ID. In case of disagreement, a discussion was initiated between the two authors, and one other author (a security expert) was then involved to provide his assessment. This process continued until all the descriptions of the security issues in the results were matched with CWE IDs. Table 7 shows the list of manually matched CWE IDs and warning messages from CodeQL, ESLint, and FindBugs results. In the final stage, we performed a statistical analysis of CWE weaknesses in 156 code snippets that contained security weaknesses.
Results. Table 8 shows the distribution of CWEs in the code snippets, including the number of code snippets that contain a certain CWE (Related Snippets) and the total number of occurrences (Frequency) of the CWE in the code snippets (we put those CWEs whose Frequency = 1 in “Others”). Note that one related code snippets may contain multiple instances of specific CWE. In total, we found 600 CWEs in 435 code snippets. These security weaknesses were related to 42 types of CWE, indicating that developers face a variety of security weaknesses when using Copilot. CWE-78: OS Command Injection is the most frequently occurred CWE, as it was detected in 15 code snippets (representing 14% of the security weaknesses), followed by CWE-330: Use of Insufficiently Random Values, CWE-703: Improper Check or Handling of Exceptional Conditions , CWE-400: Uncontrolled Resource Consumption and CWE-502: Deserialization of Untrusted Data. Some CWEs appeared less frequently, such as CWE-95: Eval Injection, and CWE-22: Improper Limitation of a Pathname to a Restricted Directory.
Additionally, many CWEs occur with a probability of less than 1%, for example, CWE-176: Improper Handling of Unicode Encoding, CWE-312: Cleartext Storage of Sensitive Information, and CWE-326: Inadequate Encryption Strength. This indicates that the types of security issues are closely related to the specific scenarios in which developers use Copilot and make the security issues become apparent, emphasizing the importance of maintaining vigilance and caution when programming.
Table 7. The warning messages of the scan results by CodeQL, ESLint, and FindBugs manually matched with corresponding CWE IDs
Tool Warning Message CWE-ID
CodeQL
Reflected server-side cross-site scripting
CWE-79
CodeQL
Flask app is run in debug mode
CWE-215
CodeQL
Clear-text logging of sensitive information
CWE-532
CodeQL
Clear-text storage of sensitive information
CWE-312
CodeQL
Information exposure through an exception
CWE-209
CodeQL
Request without certificate validation
CWE-295
CodeQL
Assignment to constant
CWE-682
CodeQL
Log injection
CWE-117
CodeQL
Identical operands
CWE-570
CodeQL
Incomplete string escaping or encoding
CWE-176
CodeQL
DOM text reinterpreted as HTML
CWE-79
CodeQL
Arbitrary file write during archive extraction (“Zip Slip”)
CWE-22
CodeQL
Hard-coded credential in API call
CWE-798
CodeQL
Hard-coded credentials
CWE-798
CodeQL
Uncontrolled data used in path expression
CWE-22
CodeQL
Resource not released in destructor
CWE-416
CodeQL
Missing Dispose call on local IDisposable
CWE-690
CodeQL
Use of the return value of a procedure
CWE-252
CodeQL
Dereferenced variable may be null
CWE-476
ESLint
Generic Object Injection Sink
CWE-502
ESLint
Function Call Object Injection Sink
CWE-20
ESLint
Unsafe Regular Expression
CWE-20
ESLint
Variable Assigned to Object Injection Sink
CWE-95
FindBugs
M S Dm: Hardcoded constant database password
CWE-798
FindBugs
H I Dm: Found reliance on default encoding
CWE-116
FindBugs
M D ICAST: Integral division result
CWE-682
FindBugs
H C IL: There is an apparent infinite loop
CWE-835
FindBugs
M B RV:Exceptional return value of java.io.File.mkdirs()ignored
CWE-252
FindBugs
M D NP: Possible null pointer dereference
CWE-476
FindBugs
M B ODR: DatabaseOperate.readDatabase() may fail to close Connection
CWE-404
Table 8. Distribution of CWEs in code snippets
CWE-ID # Related Code Snippets Frequency of Specific CWE Percentage
CWE-78
15
84
14.0%
CWE-330
34
81
13.5%
CWE-703
20
78
13.0%
CWE-398
11
60
10.0%
CWE-502
15
56
9.3%
CWE-400
22
50
8.3%
CWE-20
10
19
3.1%
CWE-252
2
13
2.1%
CWE-259
6
13
2.1%
CWE-404
3
13
2.1%
CWE-451
3
13
2.1%
CWE-682
2
13
2.1%
CWE-116
3
11
1.8%
CWE-690
3
9
1.5%
CWE-798
5
9
1.5%
CWE-561
4
8
1.3%
CWE-95
4
8
1.3%
CWE-22
4
6
1.0%
CWE-327
5
5
¡1%
CWE-532
3
5
¡1%
CWE-563
4
4
¡1%
CWE-605
4
4
¡1%
CWE-89
3
4
¡1%
CWE-295
1
3
¡1%
CWE-476
1
3
¡1%
CWE-775
2
3
¡1%
CWE-117
1
2
¡1%
CWE-209
1
2
¡1%
CWE-215
2
2
¡1%
CWE-416
1
2
¡1%
CWE-570
2
2
¡1%
CWE-664
1
2
¡1%
CWE-676
2
2
¡1%
CWE-79
1
2
¡1%
CWE-94
2
2
¡1%
Others
=1
¡1%
42 Types Total: 600
4.3. RQ3: How many security weaknesses belong to the CWE Top-25?
Approach. The code in our collected dataset was generated between June 2021 and June 2023. To compare whether the security issues in Copilot generated code are widespread in this period, we chose MITRE 2022 CWE Top-25 list (The MITRE Corporation, 2022) as our baseline. Then, we compared the CWEs obtained in RQ2 with the CWE Top-25.
Results. The distribution of CWEs found compared to the MITER list is shown in Table 9. The results show that the CWE weaknesses present in the code generated by Copilot belong to eleven CWE types included in the MITER CWE Top-25 list. This means these are present issues and are currently among the most common and serious security weaknesses in practice. It is worth noting that the 237 security issues present in the code snippet correspond to these 11 CWEs, while another 31 CWEs cover the remaining 363 security issues. This indicates that the CWE Top-25 weaknesses are also prevalent in the code generated by Copilot. Therefore, developers using Copilot must pay close attention to these weaknesses and take appropriate measures to prevent them before they are integrated into their codebase. At the same time, we can see that CWE-78: OS Command Injection is the most frequently occurring weakness from the Top-25 security weaknesses, ranking sixth in the Top-25 list and first in our RQ2 results. Although CWE-400: Uncontrolled Resource Consumption is ranked towards the end of the Top-25 list, it is one of the weaknesses with a high occurrence frequency. Some CWEs with a higher ranking in the Top-25 list do not appear frequently in Copilot generated code, such as CWE-79: Cross-site Scripting and CWE-89: SQL Injection.
Table 9. The CWEs that belong to the 2022 CWE Top-25 list
CWE-ID Description # Related Snippets Frequency
CWE-78
OS Command Injection
15
84
CWE-502
Deserialization of Untrusted Data
15
56
CWE-400
Uncontrolled Resource Consumption
22
50
CWE-20
Improper Input Validation
10
19
CWE-798
Use of Hard-coded Credentials
5
9
CWE-22
Improper Limitation of a Pathname to a Restricted Directory
4
6
CWE-89
SQL Injection
3
4
CWE-476
NULL Pointer Dereference
1
3
CWE-94
Code Injection
2
2
CWE-416
Use After Free
1
2
CWE-79
Cross-site Scripting
1
2
Total 237
5. Discussion
In this section, we explain the study results in Section 5.1 and then discuss their implications in Section 5.2.
5.1. Interpretation of Results
RQ1: How secure is the code generated by Copilot?
Among the 435 code snippets generated by Copilot, we found that 35.8% of these code snippets contain security weaknesses. Those weaknesses appear in all six top-used programming languages supported by Copilot. Furthermore, when it comes to the occurrence of security issues in code snippets of different programming languages, it is important to analyze them in conjunction with the popularity of the languages (mend.io, 2023). In code snippets written in languages like Python and JavaScript, which are frequently used with Copilot, there may be a slightly higher number of security issues. However, overall the proportion of security issues across these six languages ranges from 25% to 45%, showing no significant difference.
Besides, we also found that CWE-502: Deserialization of Untrusted Data and CWE-400: Uncontrolled Resource Consumption problems mainly appeared in code snippets written in Python and JavaScript. This could be attributed to certain features that made their code more flexible, such as dynamic typing and dynamic interpretation. Therefore, developers should pay special attention to the security of their JavaScript and Python-generated code, taking appropriate measures to validate input data and manage resources effectively to minimize security risks. The results of RQ1 suggest that in practical production, although Copilot can help developers write code faster and increase productivity, additional security assessments and fixes are also required to ensure that the generated code does not introduce potential security risks.
RQ2: What security weaknesses are present in the code snippets generated by GitHub Copilot?
After conducting a security evaluation of 425 code snippets generated by Copilot, a total of 600 security weaknesses were identified, involving 42 CWE types, which is around 10% of the CWEs (439 CWE types) in software development (The MITRE Corporation, 2023). This may be due to the reason that Copilot generates code in different programming languages and application scenarios, and a wide variety of application scenarios may lead to various types of security issues. In addition, since the Copilot base model (Codex) is trained on publicly available data that potentially contain various types of security weaknesses, this can lead to the presence of multiple CWEs in the generated code by Copilot. This set of 42 CWE types covers many types of security issues, and Table 10 shows the types of security issues that these 42 CWEs are relevant to.
Table 10. The CWEs and Types of Security issues
Type of Security Issue Relevant CWEs
Web security issue
CWE-79, CWE-94, CWE-690, CWE-732
Access control issue
CWE-252, CWE-259, CWE-327, CWE-338
Input validation and representation issue
CWE-20, CWE-89, CWE-116
Command injection issue
CWE-78, CWE-563, CWE-835
SQL injection issue
CWE-89, CWE-95
File handling issue
CWE-22, CWE-570
Insecure storage
CWE-502, CWE-775
Improper error handling
CWE-398, CWE-400
Encryption issue
CWE-327
Memory management issue
CWE-416
Buffer errors
CWE-476
Insecure random number
CWE-330
Incorrect type conversion
CWE-703
The diversity of security weaknesses indicates that developers using Copilot face various security risks. These risks are diverse, covering different development environments and application scenarios. At the same time, it also reflects the inevitability of security weaknesses in Copilot generated code. Developers need to have corresponding security awareness and skills and take appropriate security measures to avoid these risks in a timely and targeted manner. In addition, we can see that developers often encounter CWE-78: OS Command Injection, CWE-330: Use of Insufficiently Random Values, andCWE-703: Improper Check or Handling of Exceptional Conditions, CWE-502: Deserialization of Untrusted Data, and CWE-400:Uncontrolled Resource Consumption, which appear in multiple code snippets and have a high frequency of occurrence. This can remind developers to take timely and targeted security measures to mitigate these risks. For example, developers should perform adequate validation of user inputs. In addition to this, it is also necessary to restrict the program’s permissions so that they only access essential resources. The results of RQ2 reveal the security weaknesses that developers may encounter in an actual production environment and their frequency of occurrence, which can help developers be aware of security aspects of code generated by Copilot and take appropriate measures to address the security weaknesses in an informed manner.
RQ3:How many security weaknesses belong to the CWE Top-25?
As shown in Table 9, eleven of the CWEs in Copilot generated code can be found in the 2022 CWE Top-25 list, covering more than 237 security issues (39.5% of 600 identfied CWEs) in our dataset. This indicates that the commonly acknowledged top 25 weaknesses in software development, which are considered the most prevalent and dangerous, are also prevalent in the code generated by Copilot. Therefore, developers need to pay special attention to these frequently occurring weaknesses and take corresponding measures to avoid and fix them. We also observed that some vulnerabilities from the CWE top-25 list were not detected in our scans, indicating that Copilot may sanitize and prevent specific weaknesses from being suggested to developers. GitHub is gradually enhancing the security of Copilot and its underlying model (Codex) (GitHub, 2023). We also identified 31 security weaknesses in the code that do not belong to the CWE Top-25 list. Although these less common security weaknesses may not be as widespread as CWE Top-25, attackers can still exploit them. For example, we only detected one instance of CWE-732: Incorrect Permission Assignment for Critical Resource in our dataset. This security weakness is not commonly found in code and only occurs when specific users have certain permissions. However, it can lead to significant security risks when it does occur. Developers should also be aware of these less common security weaknesses to fully protect their code from attacks.
5.2. Implications
Code Snippets with Security Weaknesses: In practical production, practitioners often use Copilot to generate code in six languages: Python, JavaScript, Java, C++, Go, and C#. These languages all inevitably produce security weaknesses. We conjecture that practitioners using Copilot will likely encounter security weaknesses, regardless of the programming language used, and security checks are mandatory. When using Copilot, practitioners should conduct their own assessment of the generated code with the support of security analysis tools. They should exercise extreme caution when attempting to rely entirely on Copilot’s behavior, especially for the most commonly used languages with Copilot: Python and JavaScript.
Types of Security Weaknesses in Copilot Generated Code: Practitioners using Copilot may encounter a variety of security weaknesses. Our results show that these weaknesses are related to over 40 CWEs. This finding indicates that there are diverse security scenarios in production, and practitioners must have the corresponding security awareness and skills and adopt multiple security prevention measures to address security risks so that they do not simply accept vulnerable code suggestions. At the same time, our study reveals the frequency of related CWEs. When using Copilot to generate code, practitioners should pay particular attention to specific weaknesses, such as CWE-78: OS Command Injection, CWE-330: Use of Insufficiently Random Values, CWE-502: Deserialization of Untrusted Data, CWE-703: Improper Check or Handling of Exceptional Conditions, and CWE-400: Uncontrolled Resource Consumption. Our findings can assist practitioners in proactively preventing and addressing security issues in a targeted manner.
The CWEs in Copilot Generated Code from the Top-25 CWE List: Common security weaknesses in software development are also prevalent in code generated by Copilot. As a good practice, developers can use the CWE Top-25 list as a guide to understand which security weaknesses are most common and dangerous in the generated code and take appropriate measures to improve the code security. Additionally, the CWE Top-25 provides a standardized approach for security assessment, and developers can also use it to conduct security audits of the code generated by Copilot. Developers should also follow best practices and use code analysis tools (static, dynamic, or hybrid) to check the suggested code by Copilot (or any code generation tools) before integrating any code suggestions. Such tools can safeguard the code and help in discovering weaknesses early.
6. Threats to Validity
The validity threats are discussed according to the guidelines in (Runeson and Höst, 2009). Note that we did not consider internal validity threats since we did not investigate any relationships between variables and results.
Construct Validity is the degree to which a measurement can explain the theoretical structure and characteristics of the measurement, reflecting the extent to which the studied operational measures truly represent the researcher’s ideas and the content investigated based on the research questions. This study has three threats to the construct validity: (1) Using the keyword-based search – We used a keyword-based search to collect relevant code snippets from GitHub. The results obtained through the keyword-based search may not cover all code snippets generated by Copilot on GitHub. We tried to mitigate this threat by constantly and iteratively refining the keywords and using synonyms. (2) Manual data filtering – We manually screened the results obtained from the keyword-based search by analyzing the comments, tags, and other metadata of the code snippets to determine whether they were generated by Copilot. Since this process was manually done, it may have been influenced by personal bias. At the same time, we assumed that all code files contained in projects declared to be written by Copilot in markdown files were generated by Copilot, and we included them in the research data. This has an impact on the construct validity, as we could not have excluded the possibility of human-written code files in these projects. (3) Manual association of CWEs – We manually associated some warning messages prompted by static analysis tools with CWEs. Some warning messages could be associated with multiple CWEs, while we only focused on assigning one most suitable CWE for each warning message. Personal bias may occur with this step, impacting the final association, and we mitigated the bias by two authors conducting the association with the assessment by a security expert.
External Validity refers to the extent to which research results can be generalized and the degree to which people outside the investigated cases are interested in the research results. It indicates whether the research results are representative and can be validated in similar contexts. Our dataset consists of Copilot generated code snippets collected from open-source projects on GitHub. During the filtering process, we excluded code that utilized Copilot to solve algorithmic problems, aiming to ensure that the collected data genuinely reflected real-world production environments. Due to the reason that the data from GitHub is not diversified enough, there are many code snippets from Game projects. The peculiarity of the data source may make the dataset incomplete, thereby threatening the external validity of the results. However, GitHub is one of the largest code hosting platforms in the world, with hundreds of millions of public code repositories, and is popular among developers and the technology community. The code snippets obtained from GitHub are diverse, which mitigates this threat. Furthermore, we acknowledge the need to collect more diverse code snippets from different platforms to increase the generalizability of the results. We will consider adopting more diversified ways or platforms to collect code. Additionally, due to the limitations of static analysis tools themselves, these tools could not scan all CWEs, and there is a degree of false-positives in the scanning results (as the case with static analysis tools (Sui et al., 2020; Kang et al., 2022)). Although we attempted to use two static analysis tools to increase the coverage of weaknesses, these tools may have limited abilities in analyzing some CWEs with specific error rates, which may impact the completeness and correctness of the results.
Reliability refers to the extent to which a specific research method can produce consistent results. We used multiple automated static analysis tools to analyze the Copilot generated code snippets to improve security weaknesses detection. Developers have widely used these automated tools. The querying mechanism of these tools ensures that the scan results remain consistent when used multiple times. In addition, we performed two rounds of scanning with two tools for security checks on each code snippet, intending to complement the results of one tool with the other one. By implementing these measures, we believe that our research results are reliable and these threats to reliability are mitigated.
7. Conclusions
Automatic code generation and recommendation has been an active research area due to the advancement of AI and specifically LLMs. AI code generation tools such as Copilot can greatly improve the development efficiency of programmers, but they can also introduce vulnerabilities and security risks. In this paper, we present the results of an empirical study to analyze security weaknesses in Copilot generated code found in public GitHub projects. We identified 435 code snippets generated by Copilot from GitHub projects and analyzed those snippets for security weaknesses using static analysis tools. This study aims to help developers understand the security risks of weaknesses introduced in the code generated by Copilot (and potentially similar code generation tools). Our results show: (1) 35.8% of the 435 Copilot generated code snippets contain security weaknesses, spreading across six programming languages. (2) The detected security weaknesses are diverse in nature and are associated with 42 different CWEs. The CWEs that occurred most frequently are CWE-78: OS Command Injection, CWE-330: Use of Insufficiently Random Values, and CWE-703: Improper Check or Handling of Exceptional Conditions (3) Among these CWEs, 11 appear in the MITRE CWE Top-25 list, demonstrating their commonality, current and severity. These are: CWE-78: OS Command Injection, CWE-502: Deserialization of Untrusted Data, CWE-400: Uncontrolled Resource Consumption, CWE-89: SQL Injection, CWE-20: Improper Input Validation, CWE-22: Improper Limitation of a Pathname to a Restricted Directory, CWE-94: Code Injection, CWE-476: NULL Pointer Dereference, CWE-798: Use of Hard-coded Credentials, CWE-79: Cross-site Scripting, and CWE-416: Use After Free.
In the future, we plan to: (1) collect additional code snippets from other open source repositories and industrial projects, and code snippets generated by newer releases of Copilot; (2) analyze and summarize the application scenarios of these code snippets, studying how practitioners use Copilot and fix the issues in development; and (3) compare the results with other emerging Generative AI code generation tools such as CodeWhisperer, aiXcoder, and Code Llama.

### v2

@@
 Abstract.
-Modern code generation tools use AI models, particularly Large Language Models (LLMs), to generate functional and complete code.
-While such tools are becoming popular and widely available for developers, using these tools is often accompanied by security challenges, leading to insecure code merging into the code base.
-Therefore, it is important to assess the quality of the generated code, especially in terms of its security.
-Researchers have recently explored various aspects of code generation tools, including security.
-However, many open questions about the security of the generated code require further investigation, especially the security issues of automatically generated code in the wild.
-To this end, we conducted an empirical study by analyzing the security weaknesses in code snippets generated by GitHub Copilot that are found as part of publicly available projects hosted on GitHub.
-The goal is to investigate the types of security issues and their scale in real-world scenarios (rather than crafted scenarios).
-To this end, we identified 435 code snippets generated by GitHub Copilot from publicly available projects.
-We then conducted extensive security analysis to identify Common Weakness Enumeration (CWE) instances in these code snippets.
-The results show that (1) 35.8% of Copilot generated code snippets contain CWEs, and those issues are spread across multiple languages, (2) the security weaknesses are diverse and related to 42 different CWEs, in which CWE-78: OS Command Injection, CWE-330: Use of Insufficiently Random Values, and CWE-703: Improper Check or Handling of Exceptional Conditions occurred the most frequently, and (3) among the 42 CWEs identified, 11 of those belong to the currently recognized 2022 CWE Top-25.
-Our findings confirm that developers should be careful when adding code generated by Copilot (and similar AI code generation tools) and should also run appropriate security checks as they accept the suggested code.
+Modern code generation tools, utilizing AI models like Large Language Models (LLMs), have gained popularity for producing functional code.
+However, their usage presents security challenges, often resulting in insecure code merging into the code base.
+Evaluating the quality of generated code, especially its security, is crucial.
+While prior research explored various aspects of code generation, the focus on security has been limited, mostly examining code produced in controlled environments rather than real-world scenarios.
+To address this gap, we conducted an empirical study, analyzing code snippets generated by GitHub Copilot from GitHub projects.
+Our analysis identified 452 snippets generated by Copilot, revealing a high likelihood of security issues, with 32.8% of Python and 24.5% of JavaScript snippets affected.
+These issues span 38 different Common Weakness Enumeration (CWE) categories, including significant ones like CWE-330: Use of Insufficiently Random Values, CWE-78: OS Command Injection, and CWE-94: Improper Control of Generation of Code.
+Notably, eight CWEs are among the 2023 CWE Top-25, highlighting their severity.
+Our findings confirm that developers should be careful when adding code generated by Copilot and should also run appropriate security checks as they accept the suggested code.
 It also shows that practitioners should cultivate corresponding security awareness and skills.
-Code Generation, Security Weaknesses, CWEs, GitHub Copilot
-$c$s: Software and its engineering Software development techniquesccs: Security and privacy Software security engineering
+Code Generation, Security Weakness, CWE, GitHub Copilot, GitHub Project
+journal: TOSEMccs: Software and its engineering Software development techniquesccs: Security and privacy Software security engineering
 1. Introduction
@@
 This approach can save development time and accelerate the software development process.
-Automated code generation tools have always been an active research discussion topic (Siddiq and Santos, 2022; Li et al., 2022b).
-Some of the earliest work can be traced back to the 1960s when Waldinger and Lee proposed a program synthesizer called PROW, which automatically generated LISP programs based on specifications provided by users in the form of predicate calculus (Waldinger and Lee, 1969).
+Automated code generation tools have always been a topic of active research (Siddiq and Santos, 2022; Li et al., 2022b).
+Some of the earliest work can be traced back to the 1960s, when Waldinger and Lee proposed a program synthesizer called PROW, which automatically generated LISP programs based on specifications provided by users in the form of a predicate calculus (Waldinger and Lee, 1969).
 As computer languages continued to evolve, more and more programming languages began to support meta-programming, making automated code generation technology more efficient and flexible.
-In recent years, the rapid development of artificial intelligence technologies, particularly in the form of machine learning and deep learning models, has accelerated the development of code generation technologies.
+In recent years, the rapid development of artificial intelligence technologies, particularly machine learning and deep learning models, has accelerated the development of code generation technologies.
 Recent advancements in code generation came with the emergence of Large Language Models (LLMs).
-LLMs are deep learning models trained on a large code/text corpus with powerful language understanding capabilities that can be used for tasks such as natural language generation, text classification, and question-answering systems (Chen et al., 2021).
+LLMs are deep learning models trained on a large code/text corpus with powerful language understanding capabilities that can be used for tasks such as natural language generation, text classification, and question-answer systems (Chen et al., 2021).
 Compared to previous deep learning methods, the latest developments in LLMs, such as Generative Pre-trained Transformer (GPT) models, have opened up new opportunities to address the limitations of existing automated code generation technology (Mozannar et al., 2022).
-Currently, code generation tools based on LLMs have also been widely applied, such as Codex by OpenAI (OpenAI, [n.
+Currently, LLM-based code generation tools have also been widely applied, such as Codex by OpenAI (OpenAI, [n.
 d.]), AlphaCode by DeepMind (Li et al., 2022a), and CodeWhisperer by Amazon (Becker et al., 2022).
 These models are trained on billions of public open-source lines of code, which includes public code with unsafe coding patterns (He and Vechev, 2023).
-Therefore, code generation tools based on such models can pose security risks, and the code they generate may also have security weaknesses.
-For example, GitHub Copilot may produce some insecure code, as its underlying model Codex is pre-trained on untrusted data from GitHub (Brown et al., 2020), which is known to contain buggy programs (Rokon et al., 2020).
-In addition, the code with vulnerabilities generated by these code generation tools may continue to be used to train the model, thus further generating code with vulnerabilities, leading to a vicious cycle.
+Therefore, code generation tools based on such models can pose code quality issues (Liu et al., 2024), and the code generated by these tools may also suffer from security weaknesses (She et al., 2023).
+For example, GitHub Copilot may produce some insecure code, as its underlying Codex model is pre-trained on untrusted data from GitHub (Brown et al., 2020), which is known to contain buggy programs (Rokon et al., 2020).
+According to the developer security company Snyk, GitHub Copilot may also replicate existing security issues in code to suggest insecure code when the user’s existing codebase contains security issues (Paul Krill, 2024).
+In addition, the code with vulnerabilities generated by these code-generation tools may continue to be used to train the model, thus further generating code with vulnerabilities, leading to a vicious cycle.
 Previous research has studied code generation tools, with more focus on the correctness of the results (Lertbanjongngam et al., 2022; Wong et al., 2022; Dakhel et al., 2022; Pudari and Ernst, 2023), and relatively less attention has been paid to security aspects (Pearce et al., 2022; Siddiq et al., 2022; Perry et al., 2022).
-To the best of our knowledge, potential security weaknesses in practical scenarios have not been fully considered and addressed in previous work, and GitHub Copilot clarifies that “the users of Copilot are responsible for ensuring the security and quality of their code (GitHub, [n.
-d.])”.
-GitHub also provides tools such as CodeQL to help developers scan the security issues in their code.
-To this end, we conducted an empirical study on the security weaknesses of generated code by GitHub Copilot, which is available on GitHub.
+To the best of our knowledge, potential security weaknesses in practical scenarios have not been fully considered and addressed in previous work, and GitHub Copilot clarifies that “the users of Copilot are responsible for ensuring the security and quality of their code” (GitHub, 2023b).
+Code generation algorithms of Copilot are incentivized to suggest code to be accepted rather than other code qualities, e.g., easy to read and understand, which has an adverse impact for the code quality generated by Copilot (Harding and Kloster, 2024).
+GitHub also provides tools such as CodeQL to help developers scan for security issues in their code.
+To this end, we conducted an empirical study on the security weaknesses of the generated code by GitHub Copilot, which is available on GitHub.
 We chose Copilot as our research subject because it is a commercial instance of AI-assisted programming and has gained much attention and popularity among developers since its launch in 2021.
@@
 Furthermore, thousands of developers in the GitHub community have shared their experiences of using Copilot in real-world systems (Dakhel et al., 2022).
-We collected code generated by Copilot that has been used in projects on GitHub and analyzed the security of the generated code through the lens of a real-world production environment.
+We collected the code generated by Copilot that has been used in projects on GitHub and analyzed the security of the generated code through the lens of a real-world production environment.
 Then, we used static analysis tools to perform security analysis on the collected code snippets and classified the security weaknesses in the code snippets using the Common Weakness Enumeration (CWE).
-Our findings show that: (1) 35.8% of Copilot generated code snippets have security weaknesses, and security weaknesses arise regardless of the programming language used; (2) the security weaknesses are diverse and related to 42 different CWEs, in whichCWE-78: OS Command Injection, CWE-330: Use of Insufficiently Random Values and CWE-703: Improper Check or Handling of Exceptional Conditions are the most frequently occurred; and (3) among the 42 CWEs identified, 11 CWEs belong to the currently recognized 2022 CWE Top-25.
-The contributions of this work: (1) We curated a dataset of code snippets generated by Copilot that has been used in projects on GitHub (a curated data (Fu et al., 2023) is made available online for future research in this area.) and conducted security checks on them, which can to some extent reflect the frequency of security weaknesses encountered by developers when using Copilot to generate code in actual coding; (2) We extensively checked all possible CWEs in the code snippets and analyzed them.
-This can help developers understand the common CWEs caused by using Copilot to generate code in actual coding and how to accept the code suggestions provided by Copilot safely.
+Our findings show that: (1) 29.6% of Copilot-generated code snippets have security weaknesses; (2) the security weaknesses are diverse and related to 38 different CWEs, in whichCWE-330: Use of Insufficiently Random Values, CWE-78: OS Command Injection and CWE-94: Improper Control of Generation of Code (’Code Injection’) are the most frequently occurred; and (3) among the 38 CWEs identified, eight CWEs belong to the currently recognized 2023 CWE Top-25.
+Six CWEs belong to Stubborn Weaknesses in the CWE Top 25.
+The contributions of this work: (1) We curated a dataset of code snippets generated by Copilot that has been used in projects on GitHub (a curated data (Fu et al., 2024) is made available online ) and conducted security checks on them, which can to some extent reflect the frequency of security weaknesses encountered by developers when using Copilot to generate code in actual coding.
+In addition to this, we also categorized the application areas of these code snippets; (2) We extensively checked all possible CWEs in the code snippets and analyzed them.
+This can help developers understand the common CWEs caused by using Copilot to generate code in actual coding and how to safely accept the code suggestions provided by Copilot.
 The rest of this paper is structured as follows: Section 2 presents the related work.
 Section 3 presents the research questions and the research design of this study.
-Section 4 presents our study results, which are further discussed in Section 5.
+Section 4 presents the results of our study, which are further discussed in Section 5.
 The potential threats to validity are clarified in Section 6.
@@
 2. Related Work
+In this section, we present the related work in three aspects, i.e., AI-assisted code generation tools (Section 2.1), security of code generation techniques and LLMs (Section 2.2), and static analysis tools for scanning security weaknesses (Section 2.3).
 2.1. AI-assisted Code Generation Tools
-With the rise of code generation tools integrated with IDEs, many studies have evaluated these code generation systems based on transformer models to better understand their effectiveness in real-world scenarios.
+With the rise of code generation tools integrated with IDEs, many studies have evaluated these systems based on transformer models to better understand their effectiveness in real-world scenarios.
 Previous research mainly focused on whether the code generated by these tools can meet users’ functional requirements.
@@
 Vaithilingam et al.
-(Vaithilingam et al., 2022) studied how programmers use and perceive Copilot, and they found that while Copilot may not necessarily improve task completion time or success rate, it often provides a useful starting point.
-They also noted that participants faced difficulties in understanding, editing, and debugging the code snippets generated by Copilot.
+(Vaithilingam et al., 2022) studied how programmers use and perceive Copilot and found that although Copilot may not necessarily improve task completion time or success rate, it often provides a useful starting point.
+They also noted that the participants had difficulties understanding, editing, and debugging the code snippets generated by Copilot.
 Barke et al.
@@
 They argued that software developers should check the generated code for potentially problematic code that could introduce performance weaknesses.
-These works above conducted relatively extensive evaluations of code generation tools in terms of correctness, effectiveness, and robustness.
-However, there is still room for improvement regarding its security, as detailed in the following section.
+The studies presented above have conducted a relatively extensive evaluation of code-generation tools regarding correctness, effectiveness, and robustness.
+However, its security still has room for improvement, as detailed below.
 2.2. Security of Code Generation Techniques and LLMs
 Code security is an issue that cannot be ignored in the software development process.
-Recent work has primarily focused on evaluating the security of the code generation tools and the security of the LLMs that these tools are based on.
+Recent work has primarily focused on evaluating the security of the code generation tools and the security of the LLMs based on these tools.
 Pearce et al.
@@
 The authors prompted Copilot to generate code for 89 cybersecurity scenarios and evaluated the weaknesses in the generated code.
-They found that 40% of the suggestions in the relevant context contained security-related bugs (i.e., CWE classification from MITRE (The MITRE Corporation, [n.
-d.])).
+They found that 40% of the suggestions in the relevant context contained security-related bugs (i.e., CWE classification from MITRE (The MITRE Corporation, 2023b)).
 Siddiq et al.
@@
 Khoury et al.
-(Khoury et al., 2023) studied the security of the source code generated by the ChatGPT chatbot based on LLMs, and they found that ChatGPT was aware of potential weaknesses but still frequently generated some non-robust code.
-Several researchers also compared the situation where code generation tools produce insecure code with that of human developers.
+(Khoury et al., 2023) studied the security of the source code generated by the ChatGPT chatbot based on LLMs, and found that ChatGPT was aware of potential weaknesses but still frequently generated some non-robust code.
+Elgedawy et al.
+(Elgedawy et al., 2024) compared the capabilities of four code generation models using nine code generation tasks.
+They collected 61 code outputs and studied their security.
+The results revealed that the code generated by different LLMs exhibited disparate levels of security robustness.
+Several researchers also compared the situation in which code generation tools produce insecure code with that of human developers.
 Sandoval et al.
@@
 (Asare et al., 2023) conducted a comparative empirical analysis of these tools and language models from a security perspective and investigated whether Copilot is as bad as humans in generating insecure code.
-They found that while Copilot performs differently across vulnerability types, it is not as bad as human developers when it comes to introducing vulnerabilities in code.
+They found that while Copilot performs differently across vulnerability types, it is not as bad as human developers when introducing code vulnerabilities.
 In addition, researchers have also constructed datasets to test the security of these tools.
@@
 The dataset contains 130 Python code samples.
-Unlike the works above, we studied the security weaknesses exhibited by code generation tools in a real-world production environment (i.e., GitHub).
+Different from prior work, we studied the security weaknesses exhibited by code generation tools in a real-world production environment (i.e., GitHub).
 We collected code snippets from GitHub generated by developers using Copilot in daily production as a source of research data, whereas in the Pearce et al.
-(Pearce et al., 2022) study, the research data came from code generated by the authors using Copilot based on the natural language prompts related to high-risk network security weaknesses.
-In addition to this, Pearce et al.
-configured CodeQL only to examine CWEs targeted by security weaknesses associated with the prompted scenarios.
+(Pearce et al., 2022) study, the research data came from code generated by the authors using Copilot based on natural language prompts related to high-risk network security weaknesses.
+Additionally, Pearce et al.
+configured CodeQL only to examine CWEs targeted by security weaknesses associated with the provided scenarios.
 In contrast, we used various static analysis tools to examine all types of CWEs and analyze them extensively.
@@
 Vulnerabilities detection is critical to improve software security and ensure quality  (Iannone et al., 2022).
-There are two used methods for vulnerability detection in source code: via static and dynamic code analysis.
+Vulnerability can be checked either statically or dynamically.
 Dynamic analysis techniques are more sound and precise but lack coverage (Ernst, 2003).
-On the other hand, static analysis is less precise but offers greater coverage and allows to analyze programs without the need to execute them (Sui et al., 2020).
-Static analysis has been widely used to find security issues in code, given it is cheaper to run and can conduct whole program analyses without the need to execute the program (Dunlap et al., 2023).
+On the other hand, static analysis is less precise but offers greater coverage and allows users to analyze programs without the need to execute them (Sui et al., 2020).
 OWASP (OWASP, [n.
-d.]) provides a list of commonly used static analysis tools.
-This includes tools like CodeQL: a general-purpose automatic scanning tool, FindBugs: a tool for Java programs, ESLint: a tool for JavaScript programs, Bandit: a tool for Python programs, and GoSec: a tool for Go programs.
+d.]) provides a list of commonly used static analysis tools for security analysis.
+This includes tools like CodeQL, a general-purpose automatic scanning tool; FindBugs, a tool for Java programs; ESLint, a tool for JavaScript programs; Bandit, a tool for Python programs; and GoSec, a tool for Go programs.
 Such tools have been widely used in previous security analysis research  (Tómasdóttir et al., 2018; Pearce et al., 2022; Siddiq et al., 2022).
@@
 Pearce et al.
-(Pearce et al., 2022) used CodeQL for security weakness scanning of generated Python and C++ code.
+(Pearce et al., 2022) used CodeQL to scan security weaknesses in the generated Python and C++ code.
 Siddiq et al.
 (Siddiq and Santos, 2022) used Bandit to check Python code generated using a test dataset.
+Lisa et al.
+(Do et al., 2020) reported on users’ goals, motivations, and strategies when using static analysis tools.
 These static analysis tools support different analysis algorithms and techniques.
@@
 CodeQL is an open-source tool that supports multiple languages, including Java, JavaScript, C++, C#, and Python.
-It can find weaknesses in a codebase based on a set of known weaknesses/rules.
+It can find weaknesses in a codebase based on known weaknesses/rules.
 In addition, to obtain more comprehensive scan results, we supplemented the scan of code in different languages with static analysis tools (i.e., Cppcheck and Bandi) tailored to specific languages.
@@
 3.1. Research Goal and Questions
-This empirical study aims to understand the potential security weaknesses in Copilot generated code.
-We first collected code snippets generated by Copilot from GitHub projects as the data source for our research.
+This study aims to understand the potential security weaknesses in Copilot-generated code produced in real-world GitHub projects.
+We first collect code snippets generated by Copilot from GitHub projects as our data source, which is the largest source of open-source projects.
 It should be noted that it is not possible to access all the code generated by Copilot in GitHub projects, as there is no direct way to identify if part of a file was generated by Copilot (i.e., source files do not contain any signatures to indicate if Copilot generates the code).
-However, we can identify many code snippets by searching through the repository description and the comments provided in the code (details provided in Section 3.2.2).
-We then analyzed the identified snippets for security weaknesses.
-We aim to help developers and researchers understand common security weaknesses when using Copilot without focusing on whether the code generation aspects are used correctly.
-We conducted this empirical study following the guidelines proposed in (Easterbrook et al., 2008).
-The RQs, their rationale, and the research process of this study (see Fig.
-1) are detailed in the subsections below.
+However, we can identify many code snippets by searching the repository description and the comments provided in the code (see the details in Section 3.2.2).
+We chose to focus on Python and JavaScript code snippets as they are currently the two most popular programming languages used by developers according to a recent survey in 2022 by GitHub (oct, [n.
+d.]), and also the ones most frequently used with GitHub Copilot (Zhang et al., 2023).
+We first analyzed the functionality and application domains of the collected code snippets to get the demographic information of the Copilot-generated code.
+Next, we performed a security analysis on the identified code.
+We selected static analysis because of its coverage and the ability to analyze programs without executing them.
+Dynamic analysis is more sound and precise (as it can reason about the program behaviour) but also lacks coverage (Sui et al., 2020).
+Using static analysis will allow us to run the analysis on a segment of the program (smaller snippet) without needing the full program to be executed.
+Static security analysis tools have been widely used by developers and companies (Majdinasab et al., 2024; Ernst, 2003; Pearce et al., 2022).
+After obtaining the scan results, we manually checked the results to remove false positives reported by the static analysis tools.
+We finally used CWEs to classify the filtered results for further analysis to answer the RQs.
+We conducted this empirical study by following the guidelines of Easterbrook et al.
+(Easterbrook et al., 2008).
+The RQs, their rationale, and the research process of this study (see Figure 1) are detailed in the subsections below.
 Figure 1. Overview of the research process
 RQ1. How secure is the code generated by Copilot in GitHub projects?
-Rationale: Copilot may produce code suggestions that developers accept but these suggestions may include security weaknesses that could potentially make the program vulnerable.
-The answer to RQ1 helps to understand the frequency of security weaknesses developers encounter when using Copilot in production.
+Rationale: Copilot may produce code suggestions that developers accept, but these suggestions may include security weaknesses that could make the program vulnerable.
+The answer to RQ1 helps understand the frequency of security weaknesses developers encounter when using Copilot in production.
 RQ2. What security weaknesses are present in the code snippets generated by Copilot?
-Rationale: Copilot generated code may contain security weaknesses (Pearce et al., 2022), and developers should conduct a rigorous security review before accepting the code generated by Copilot.
-As clarified in the documentation of GitHub Copilot “the users of Copilot are responsible for ensuring the security and quality of their code (GitHub, [n.
-d.])”.
-The answer to RQ2 can help developers better understand possible security weaknesses in the code generated by Copilot, thereby enabling them to prevent and fix these weaknesses more effectively.
-RQ3. How many security weaknesses belong to the MITRE CWE Top-25?
-Rationale: The MITRE list contains the top 25 most dangerous security weaknesses.
-The answer to RQ3 can help developers understand whether the code generated by Copilot contains widely recognized types of security weaknesses and Copilot’s ability to handle these recent and common weaknesses.
+Rationale: Copilot-generated code may contain a variety of security weaknesses (Pearce et al., 2022), and developers should conduct a rigorous security review before accepting the generated code suggestions.
+Copilot’s documentation notes the following: “users of Copilot are responsible for ensuring the security and quality of their code (GitHub, 2023b)”.
+By identifying common or recurring weaknesses, developers can be more prepared to prevent, mitigate, or fix these security weaknesses.
+The answer to RQ2 can help developers better understand possible security weaknesses in the code generated by Copilot, allowing them to be vigilant about the generated code before it is integrated into their programs.
+RQ3. How many security weaknesses belong to the MITRE CWE Top-25 security weaknesses?
+Rationale: The MITRE list contains the Top 25 most dangerous security weaknesses (The MITRE Corporation, 2023a), offering a benchmark for gauging the severity of weaknesses in the Copilot-generated code (Pearce et al., 2022).
+The answer to RQ3 can help to understand whether the code generated by Copilot contains widely recognized types of security weakness and Copilot’s ability to handle these recent and common weaknesses.
 3.2. Data Collection and Filtering
@@
 GitHub is widely used by developers.
-As the world’s largest code hosting platform, GitHub contains millions of public code repositories and offers access to a large number of code resources, allowing us to cover multiple programming languages and project types in our study.
-We used code snippets generated by Copilot on GitHub as our research object to analyze the relevant security weaknesses of Copilot.
-Our scripts and dataset are provided online in our replication package (Fu et al., 2023).
+As the world’s largest code hosting platform, GitHub contains millions of public code repositories and offers access to a large number of code resources, allowing us to cover multiple programming languages and project types in our study (Cosentino et al., 2017).
+We identify code snippets generated by Copilot from GitHub projects.
 3.2.1. Code Snippets Collection
 Step 1.
-To collect code snippets generated by Copilot, we first conducted a pilot search to formulate our search keywords.
+To identify the code snippets generated by Copilot, we first conducted a pilot search to formulate our search keywords.
 First, we used “GitHub Copilot” and “Copilot” as our search keywords.
-As expected, we found that the term “Copilot” not only refers to the code generation tool launched by GitHub but also to tools in the aviation or telemetry fields.
-Therefore, using the keyword “Copilot” solely may return irrelevant content that may not be related to the use of the tool.
-On the other hand, using “GitHub Copilot” as a search keyword can exclude content unrelated to the code generation tool Copilot and narrow the search scope, which is what we have used to locate the code snippets.
-However, even with this basic search keyword, we still need to carefully filter the search results to ensure they are truly related to GitHub Copilot.
-Although using “GitHub Copilot” as a search keyword increases the relevance of the results to Copilot, these results are not necessarily the code snippets generated by Copilot.
+As expected, we found that the term “Copilot” is a vague term, referring not only to the code generation tool launched by GitHub, but also to tools in the aviation or telemetry fields.
+Therefore, using the keyword solely may return irrelevant content that may not be related to the use of the tool.
+On the other hand, using “GitHub Copilot” can exclude content unrelated to the code generation tool Copilot and narrow the search scope, which is what we have used to locate the code snippets.
+However, even with this basic search keyword, we still need to carefully filter the search results to ensure that they are truly related to GitHub Copilot.
+Although using “GitHub Copilot” increases the relevance of the results to Copilot, these results are not necessarily the code snippets generated by Copilot.
 It should be noted that many code snippets containing the “GitHub Copilot” keyword in the search results display GitHub Copilot as text.
 Developers may use them to describe their experience using Copilot to generate code or showcase information related to Copilot.
-These code snippets are not what we need because they do not directly relate to the code generated by Copilot.
+These code snippets are not what we seek as they do not directly relate to the code generated by Copilot.
 Our target is code generated by Copilot, not code snippets containing the keyword “Copilot”.
-Our observations from the pilot search showed that using keywords such as “by GitHub Copilot”, “use GitHub Copilot”, and “with GitHub Copilot” can improve the accuracy of the search results.
+Our observations from the pilot search showed that using keywords such as “by GitHub Copilot”, “use GitHub Copilot”, and “with GitHub Copilot” can improve the accuracy of search results.
 These keywords enable us to focus more on the code generated using Copilot rather than code snippets that contain other content related to Copilot.
-In addition, since our goal is to use automated analysis tools to perform security scans on the collected code snippets, we further limited the types of code snippets during the search to Python, JavaScript, Java, C++, C#, and Go.
-These are the mainstream languages supported by Copilot, and also the languages supported by CodeQL.
-We collected the Code parts from these search results.
+In addition, since Python and JavaScript are the two most popular programming languages used by developers, and also the most frequently used with GitHub Copilot (see Section 3.1), we further limited the types of code snippets during the search to Python and JavaScript.
+We collected the search results under the Code label.
 Considering that some projects declare using GitHub Copilot generated code in their README files or project description provided in GitHub, we decided to retain the results from the Repository label in the search results.
-Fig.
-2 shows an example of our search process.
+Figure 2 shows an example of our search process.
 Table 1 reports the search terms we used and the number of search results obtained from GitHub.
-In this step, we collected a total of 8,004 results, of which 7,749 were from the Code label, and 255 were from the Repository label.
-The same search result may contain multiple keywords, meaning there are duplicate projects in the collected data.
-After removing duplicate projects, we obtained a total of 4249 search results, of which 4081 were from the Code label, and 168 were from the Repository label.
-Table 2 shows the number of different language types of search results we obtained from the Code label.
+In this step, we collected a total of 8,157 results, of which 7,749 were from the Code label, and 408 were from the Repository label.
+The same search result may contain multiple keywords, which means that there are duplicate projects in the collected data.
+After removing duplicate projects, we obtained a total of 2023 search results, of which 1847 were from the Code label, and 176 were from the Repository label.
 Figure 2. Example of the search process
@@
 # Search Term # Code # Repositories
-1
-“By GitHub Copilot”
-2549
-54
-2
-“Use GitHub Copilot”
-1822
-77
-3
-“With GitHub Copilot”
-3378
-127
-Total 7749 255
-Table 2. Search Results from GitHub
-# Language # Code results
-L1
-Python
-784
-L2
-JavaScript
-863
-L3
-Java
-641
-L4
-C++
-437
-L5
-C#
-386
-L6
-Go
-970
-Total 4081
+1 “By GitHub Copilot” 2549 94
+2 “Use GitHub Copilot” 1822 123
+3 “With GitHub Copilot” 3378 191
+Total 7749 408
 3.2.2. Filtering Code Snippets
 Step 2.
-After obtaining the results from the keyword searches, we further filtered them by not only considering the accuracy of the keywords but also investigating the project’s documentation, code comments, and other metadata in the search results to determine whether they were generated by GitHub Copilot.
-Additionally, since we wanted to obtain code snippets used in real-world projects, we excluded search results used to solve simple algorithmic problems on platforms, such as LeetCode, which generally involve simple code and may not involve security weaknesses.
+After obtaining the data from the keyword searches, we further filtered them by not only considering the accuracy of the keywords but also by investigating the project’s documentation, code comments, and other metadata in the search results to determine whether they were generated by GitHub Copilot.
+Additionally, since we aimed to obtain code snippets used in real-world projects, we excluded search results used to solve simple programming practice problems on platforms, such as bisection lookup and fast sorting tasks from LeetCode.
+We consider these programming exercises to be more closely related to correctness than security.
+We first conducted a pilot data filtration to better filter the search results.
 We begin by explaining the terminology used in data filtering: the search results under the Repository label are the projects that contain code files, and the search results under the Code label are individual code files.
-Those code files contain Copilot generated code snippets.
-In filtering the projects, we followed three rules: (1) for search results under the Repository label, we identified projects that are fully generated by Copilot, as declared in the projects description or the associated README file(s).
-We retained code files for Python, JavaScript, Java, C++, C#, and Go, which are the main languages supported by Copilot.
-(2) For search results under the Code label, we retained code files with comments showing the code generated by Copilot.
-(3) As we mentioned above, we then excluded code files used to solve simple algorithm problems.
-We provide examples for the three rules in Figs.
-3,  4 and  5.
-As shown in the example in Fig.
-3 for the Repository label, we kept all the Python files.
-In the next example in Fig.
-4, we kept the entire file where the Copilot generated code snippet was located.
-In Fig.
-5, the code snippet was removed as it was determined the code just solved a simple algorithmic problem.
-Meanwhile, for the code files retained under the Repository label, we consider the entire code file as code generated by Copilot.
-In other words, we assume that all code in the file is generated by Copilot because it was stated in the README file that it was all generated by Copilot.
+Those code files contain code snippets generated by Copilot.
+In filtering the projects, we followed three criteria including two inclusion criteria and one exclusion criterion.
+Inclusion Criterion 1: for search results under the Repository label, we identified projects that are fully generated by Copilot, as declared in the project description or the associated README file(s).
+We retained code files for Python and JavaScript.
+Inclusion Criterion 2: For search results under the Code label, we retained code files with comments showing the code generated by Copilot.
+Exclusion Criterion 1: We excluded code files used to solve simple programming practice problems.
+We provide examples for the three criteria in Figure 3, Figure 4, and Figure 5.
+As shown in the example in Figure 3 for the Repository label, we kept all the Python files.
+In the next example in Figure 4, we kept the entire file where the Copilot-generated code snippet was located.
+In Figure 5, the code snippet was removed as it was determined that the code just solved a simple algorithmic problem.
+Specifically, the pilot data filtering process consisted of the following steps: (1) The first author randomly selected 40 projects and 400 code files from the search results from the Repository label and the Code label, respectively.
+(2) Two authors independently labelled whether Copilot generated these projects and code files.
+(3) The first author compared the labelling results by the two coders and calculated the level of agreement between them using the Cohen’s Kappa coefficient (Cohen, 1960).
+(4) If there were any results they were unsure of or disagreed with, the two authors discussed until they reached an agreement.
+The Cohen’s Kappa coefficients were 0.79 for the projects (from the Repository label) and 0.85 for the code files (from the Code label), which were both higher than 0.7, indicating a high level of agreement between the two coders and ensuring a good accuracy of the labelling results.
+After completing the pilot data labelling, the first author checked the rest of the search results (1,447 code files and 136 projects).
+After removing the duplicate results, we retained 79 projects under the Repository label and 180 code files under the Code label.
+Step 3.
+To further get the application domains represented in the dataset, we categorized the projects based on the project’s description documents and the specific function of the code.
+Projects were classified into one of the following categories: Games, Web Applications, Utility Tools, AI Applications, Network Communication Applications, and Others.
+Figure 6 shows the classification and distribution of the application domains of the Copilot-generated code in the dataset from the Repository and Code labels respectively.
+For the code files retained under the Repository label, we consider the entire code file as code generated by Copilot.
+In other words, we assume that Copilot generates all the code in the file because it was stated in the README file that it was all generated by Copilot.
 For code files retained under the Code label, we know that the files contain code snippets, perhaps even just a few lines of code, generated by Copilot.
-Instead of identifying the specific Copilot generated code in this step, we combine the warning messages from the security scan and the code comments in the file to determine whether Copilot generates the code snippet with the security problem (this process is explained further in Section 3.3.2).
-After completing the pilot data labeling, the first author checked the rest of the search results, and obtained a total of 465 code snippets.
-After removing duplicate results, we finally obtained 435 different code snippets.
-Among them, 249 are from the Repository label, and 186 are from the Code label.
-Table 3 shows the types and numbers of code snippets obtained.
-Figure 3. Example of rule 1: project is fully written by Copilot
-Figure 4. Example of rule 2: files with comments showing the code generated by Copilot
-Figure 5. Example of rule 3: files used to solve simple algorithm problems
+Instead of identifying the specific Copilot-generated code in this step, we combine the warning messages from the security scan and the code comments in the file to determine whether Copilot generates the code snippet with the security problem (this process is explained further in Section 3.3.2).
+As a result, we obtained 452 code files with different contents, 272 from the Repository label and 180 from the Code label.
+Table 2 gives the type and number of code files.
+A curated dataset, comprising all the data collected during the research process, has been made available (Fu et al., 2024).
+Figure 3. Example of Inclusion Criterion 1: projects fully written by Copilot
+Figure 4. Example of Inclusion Criterion 2: files with comments showing the code generated by Copilot
+Figure 5. Example of Exclusion Criterion 1: files used to solve simple algorithm problems
+(a) Code from the Repository label
+(b) Code from the Code label
+Figure 6. Application domains of the Copilot-generated code in our dataset
 Table 2. Code snippets from GitHub
 # Language # Code Snippets: Repository # Code Snippets: Code Total
-L1 Python 132 119 251
-L2 JavaScript 51 28 79
-L3 Java 25 18 43
-L4 C++ 14 12 26
-L6 Go 19 1 20
-L5 C# 8 8 16
-Total 249 186 435
-3.3. Data Pre-processing and Analysis
-3.3.1. Data Pre-process
-Step 3.
-CodeQL is a scalable static security analysis tool that is widely used in practice, and enables users to analyze code and detect relevant weaknesses using predefined queries and test suites and supports for multiple languages (including Java, JavaScript, C++, C#, Go, and Python (GitHub, 2021a)).
+L1 Python 143 134 277
+L2 JavaScript 129 46 175
+Total 272 180 452
+3.3. Data Processing and Analysis
+3.3.1. Data Processing
+Step 4.
+CodeQL is a scalable static security analysis tool that is widely used in practice and allows users to analyze code and detect relevant weaknesses using predefined queries and test suites and supports for multiple languages (including Java, JavaScript, C++, C#, Go, and Python (GitHub, 2023a)).
 Before using CodeQL to scan the identified code snippets for security weaknesses, we needed to create a CodeQL database for the source code.
-For interpreted languages like Python and JavaScript, the source code can be directly analyzed, while for compiled languages such as Java, the source code will need to be compiled first and then imported into the CodeQL database.
-Therefore, we first compiled the code snippets of all compiled languages (i.e., C#, Java, C++, and Go).
-We removed any code snippets that could not be compiled.
-For successfully compiled files, we generated the CodeQL database required for queries.
-At the same time, for interpreted languages Python and JavaScript files, we stored 20 files in each database to improve efficiency, because if we generate a database for an exceptionally large number of files, this would increase the database compilation and scanning time, which is much longer than partitioning them into small databases.
-In total, we obtained 80 code databases available for CodeQL scanning.
-Table 4 shows the types and numbers of files in each database.
-Table 4. Databases for CodeQL scanning
-# Language # Databases: Repository # Databases: Code
-L1 Python 7 6
-L2 JavaScript 3 2
-L3 Java 8 18
-L4 C++ 5 12
-L5 C# 3 7
-L6 Go 13 1
-Total 39 46
+Source code can be directly analyzed for interpreted languages like Python and JavaScript without being compiled into intermediate code.
+To speed up the scan, we stored 20 files in each database using the command line to generate the database needed for CodeqQL queries.
+It should be noted that generating a database for an exceptionally large number of files would increase the database compilation and scanning time, which is much longer than partitioning them into small databases.
+In total, we obtained 25 code databases available for CodeQL scanning.
 3.3.2. Data Analysis
-Step 4.
+Step 5.
 We used well-known automated static analysis tools listed by OWASP (OWASP, [n.
 d.]) to scan the collected code snippets.
+Static analysis has been widely used to detect security issues in code, for its ability to analyze programs without execution (Dunlap et al., 2023).
+While dynamic analysis offers more precise insights by reasoning about program behavior, it suffers from limited coverage (Sui et al., 2020).
+Employing static analysis enables us to analyze program segments (smaller snippets) without requiring the entire program to be executed.
 Since different static analysis tools may use different algorithms and rules to detect security weaknesses, using multiple tools can increase our chances of discovering security issues in the code.
-To improve the coverage and accuracy of the results, we used two static analysis tools for security checks on each code snippet (i.e., CodeQL plus a dedicated tool for the specific language).
+To improve the coverage and accuracy of the results, we used two static analysis tools for security checks on each code snippet (i.e., CodeQL plus one dedicated tool for the specific language, Bandit for Python and ESLint for JavaScript).
 We first used CodeQL to analyze the code in our dataset.
 The default query suite for the standard CodeQL query package is codeql-suites/<lang>-code-scanning.qls.
-There are several useful query suites in the codeql-suite directory of each package.
-For example, the codeql/cpp-queries package contains the following query suites (GitHub, 2021b):
-  • Cpp-code-scanning.qls, which is the standard code scanning query for C++. It covers various features and syntax of C++ and aims to discover some common weaknesses in the code.
-  • Cpp-security-extended.qls, which includes some more advanced queries than cpp-code-scanning.qls and can detect more security weaknesses.
-  • Cpp-security-and-quality.qls, which combines queries related to security and quality, covering various aspects of C++ development from basic code structure and naming conventions to advanced security and performance weaknesses.
+Each package has several query suites in the codeql-suite directory.
+For example, the codeql/python-queries package contains the following query suites (GitHub, 2023):
+  • Python-code-scanning.qls, the standard scanning query for Python. It covers various features and syntax of Python and aims to discover some common weaknesses in the code.
+  • Python-security-extended.qls, which includes some more advanced queries than Python-code-scanning.qls and can detect more security weaknesses.
+  • Python-security-and-quality.qls, which combines queries related to security and quality, covering various aspects of Python development, from basic code structure and naming conventions to advanced security and performance weaknesses.
 It aims to help developers improve the security and quality of their code.
-In this study, we scanned code snippets using the <language>-security-and-quality.qls test suite related to security weaknesses.
+In this study, we scanned code snippets using the <language>-security-and-quality.qls test suite.
 These test suites check for multiple security properties and cover many CWEs.
-For example, the python-security-and-quality.qls test suite for Python provides 168 security checks, the JavaScript test suite provides 203 security checks, and the C++ test suite provides 163 security checks.
+For example, the python-security-and-quality.qls test suite for Python provides 168 security checks, and the JavaScript test suite provides 203 security checks.
 As the query reports only provide the name and description of the security issues, we manually matched the results in the query reports with the corresponding CWE IDs.
-We then selected other popular static security analysis tools for files in each program languages we analyzed.
-We used the following popular security analysis tools: Bandit for Python, ESLint for JavaScript, Cppcheck for C ++, Findbugs for Java, Roslyn for C# and Gosec for Go.
-In cases where we could not directly obtain the CWE ID related to the security issue from the scan results, we manually mapped the security attributes to the corresponding CWE for later analysis.
-We explain the specific correspondences in detail in Section 4.2.
-Step 5.
+We then selected one additional static security analysis tool for files in each programming language we analyzed: Bandit for Python and ESLint for JavaScript.
+As explained in Section 3.2, we considered the code snippet from the Repository label to be the entire code file, while the code snippet from the Code label exists in the code file, with the exact number of lines unspecified at this stage.
+Step 6.
 We scanned code snippets from the Repository and Code labels, and we filtered the scan results before analyzing them.
-We first removed the scan results that were repeatedly prompted by two of the tools, then removed the results that were unrelated to the security issue, and finally confirmed that the Copilot generated code indeed caused the results related to the security issue.
-As we explained in Section 3.2.2, we considered the code snippet from the Repository label to be the entire code file.
-Therefore, we kept the entire scan results from the Repository label and counted all the security issues they suggested.
-For the code snippet obtained from the Code label, we started by scanning the code file.
-If the static analysis tool found a security issue in the code, we located the code snippet in the file according to the line number of the security issue indicated by the scanning result.
-We determined whether it was generated by Copilot based on the comments before and after the code snippet.
-If Copilot indeed generated the code snippet with a security issue, we kept the scan result for our subsequent statistics.
-We further analyzed the filtered scan results in Step 6, detailed in Section 4 according to the specific RQs.
-We provide our full dataset (including code snippets, full scan results, and filtered results) in our replication package (Fu et al., 2023).
+In this step, we adopted part of the strategies used by developers when they perform different tasks with static analysis tools: warning prioritization and determining whether a warning is a false positive or a true report (Do et al., 2020).
+We first performed an initial filtering of the results based on the priority of the warnings.
+Specifically we removed the repeated scan results that were reported by the two tools, then removed the scan results that were not security issues, such as recommendations, to get the initial scan results.
+Subsequently, to confirm that the security issues were actually caused by the Copilot-generated code, the first author manually checked the scan results.
+For the initial results obtained under the Repository label, we checked them one by one.
+Specifically, we determined whether the corresponding location of the code snippet indeed had a security issue based on the line number information provided by the scan results.
+For the initial results obtained under the Code label, we confirmed whether the security issues truly existed and verified if the security issues were caused by Copilot-generated code.
+Specifically, after scanning the code file, we pinpointed the code snippet within the file based on the line number of the security issue indicated in the scan results.
+We assessed whether it was generated by Copilot by checking the surrounding comments and determined whether a security issue existed in that particular context.
+We further analyzed the filtered scan results in Step 7, detailed in Section 4, according to the specific RQs.
+We provide the complete dataset (including code snippets, full scan results, and filtered results) in our replication package (Fu et al., 2024).
 4. Results
-We present the results of three RQs formulated in Section 3.1 below.
-For each RQ, we first explain how we analyzed the collected code snippets to answer the RQ.
-We then provide a detailed presentation of the final results for each RQ.
+We present the results of three RQs formulated in Section 3.1 below. For each RQ, we first explain how we analyzed the collected code snippets to answer the RQ. We then answer each of our three RQs.
 4.1. RQ1: How secure is the code generated by Copilot?
 Approach.
-To answer this RQ, we collected 435 code snippets generated by Copilot from GitHub projects.
-These snippets cover six common programming languages.
-We used two static analysis tools (CodeQL + another language-dedicated tool) to scan and analyze the code snippets and then combine the results obtained from the two tools.
-The aim is to achieve a better coverage of security issues.
+To answer this RQ, we collected 452 Python and JavaScript code snippets generated by Copilot from GitHub projects.
+We used two static analysis tools (CodeQL + another language-dedicated tool, Bandit for Python and ESLint for JavaScript) to scan and analyze the code snippets and then combine the results obtained from the two tools.
+The aim is to achieve better coverage of security issues.
 Therefore, as long as one of the tools detected the presence of a security issue, the code snippet was considered vulnerable.
-In the analysis results obtained from the CodeQL tool, three types of warnings were used to describe the detected weaknesses: Recommendation, which provides suggestions for improving code quality; Warning, which alerts to potential weaknesses that could cause code to run abnormally or unsafely; and Error, which is the highest level of warning and alert to inform that the error could cause code to fail to compile or run incorrectly.
-Since our research primarily focused on security weaknesses, we only counted code snippets that had warnings and errors, and we ignored the other code quality recommendations.
-For the scanning results from the Code label, we also needed to identify whether the security issues obtained from the scan were from Copilot generated code snippets based on the comment that appears before the method.
-We provide an example of the scan results filtration in Fig.
-6.
-In Step 1, we first went to the corresponding file to locate the specific code snippet based on the start and end lines of the scan results that suggested a security issue.
-In Step 2, we located the code at Line 103 and found no comment indicating that Copilot generated it.
-In Step 3, we also found that the code snippet generated by Copilot in the file was located at Lines 226 to 239, and we then determined that the code snippet generated by Copilot did not cause this security issue and discarded this scan result from further analysis.
-Finally, we aggregated the filtered results obtained using multiple analysis tools to calculate the number of code snippets with security issues detected.
+We identified three types of warnings from CodeQL analysis:
+  • Recommendation, which provides suggestions for improving code quality;
+  • Warning, which alerts to potential weaknesses that could cause code to run abnormally or unsafely;
+  • Error, which is the highest level of warning and alert to inform that the error could cause code to fail to compile or run incorrectly.
+Since our research primarily focused on security weaknesses, we only counted code snippets that had warnings and errors, and we ignored recommendations on code quality.
+After obtaining the initial scan results related to security issues, we manually checked the scan results.
+For the scan results from the Repository label, we marked them as 1 (security issue exists) and 0 (no security issue exists).
+For the scan results from Code label, we marked them as 1 (security issue exists) and 0 (no security issue exists or the code snippet with a security issue was not generated by Copilot).
+Figure 7 shows the outcomes of the manual verification on the initial security scan results under the Repository and Code labels.
+When manually checking the scan results from the Code label, we also needed to identify whether the security issues obtained from the scan were from Copilot-generated code snippets based on the comment (prompt message) that appears before the method.
+We provide a working example of the filtration of the scan results in Figure 8.
+In Step 1, we first went to the corresponding file to locate the specific code snippet based on the start line numbers of the scan results.
+In Step 2, we located the code at Line 149 in aggregation.py.
+We found that this code does have a security issue and determined that it was generated by Copilot based on the prompt messages above.
+Consequently, we marked the corresponding security scan result as “1”.
+We also located the code at Line 145 in _lytools.py.
+We found that this code had a security issue, but there were no prompt messages that would indicate that it was generated by Copilot.
+Consequently, we marked the corresponding security scan result as “0” and discarded this scan result from further analysis.
+Finally, we kept the results marked as 1 and aggregated the filtered results obtained using multiple analysis tools to calculate the number of code snippets with security issues detected.
+(a) Results from the Repository label
+(b) Results from the Code label
+Figure 7. Manual verification of the initial security scan results for the Repository and Code label
 Figure 8. Example of filtering scan results from the Code label that are generated by Copilot
@@
 Table 3 shows the numbers of code snippets for different types and the numbers and percentages of code snippets with security weaknesses.
-From the statistical results, we found that out of the 435 code snippets generated by Copilot, 35.8% of them have security weaknesses, regardless of the programming language.
-There is a higher proportion of security weaknesses in Python and JavaScript code, which are the most popular Copilot languages developers use (Stackscale, 2021).
-Out of the 251 Python code snippets, 39.4% have security weaknesses.
-Among the 79 JavaScript code snippets we collected, 27.8% have security weaknesses.
-Among all programming languages, C++ code snippets have the highest proportion of security weaknesses, reaching 46.1%.
-Go also has a relatively high proportion of security weaknesses, at 45.0%.
-In comparison, the proportion of files with security issues is lower for C# and Java code, at 25% and 23.2%, respectively.
+From the statistical results, we found that out of the 452 code snippets generated by Copilot, 25.9% of them have security weaknesses.
+Among the Python and JavaScript code snippets, Python exhibits a higher proportion of security issues compared to JavaScript.
+Out of the 277 Python code snippets, 91 (32.8%) have security weaknesses.
+Among the 175 JavaScript code snippets we collected, 43 (24.6%) have security weaknesses.
+Note that one related code snippet may contain multiple instances of a specific CWE.
 Table 3. The number and percentage of code snippets with security weaknesses generated by Copilot
@@
 security weaknesses
-Python 251 99 39.4%
-JavaScript 79 22 27.8%
-Java 43 10 23.2%
-C++ 26 12 46.1%
-Go 20 9 45.0%
-C# 16 4 25%
-Total 435 156 35.8%
+Python 277 91 32.8%
+JavaScript 175 43 24.6%
+Total 452 134 29.6%
 4.2. RQ2: What security weaknesses are present in the code snippets generated by Copilot?
 Approach.
-To answer RQ2, we processed the results of the scans conducted for RQ1, eliminating duplicate security issues detected at the same code snippet location.
-In total, we identified 600 security weaknesses across 435 code snippets.
-Table 6 shows the number of security weaknesses found in code files of different programming languages.
+To answer RQ2, we processed the scan results collected by RQ1 that were manually checked to contain vulnerabilities.
+In total, we found 544 security issues in 452 code snippets, and Table 4 shows the number of security weaknesses found in code files of different programming languages.
 Table 4. The number of security weaknesses in code snippets generated by Copilot
@@
 weaknesses
-Python 99 352
-JavaScript 22 93
-Java 10 56
-C++ 12 70
-Go 9 18
-C# 4 11
-Total 156 600
-For each code snippet, we used CWEs to classify the security issues identified by the static analysis.
-Each CWE has a unique ID and a set of related descriptions, including its potential impact and how to detect and fix the CWE (The MITRE Corporation, [n.
-d.]).
-Some static analysis tools we used, such as Bandit and Gosec, provide a CWE ID corresponding to the detected security issues in their scan results.
-For other scan tools that do not directly give a CWE ID, such as Codeql, ESLint, and FindBugs, we manually associated the provided security issue information with a CWE ID, which is detailed below.
-The QL queries used in CodeQL often point to a specific CWE, and the scanning process typically displays the CWE-ID associated with the QL query.
-Although only the name of the QL query is shown in the scan results, we can manually correlate the name field with the CWE ID.
-For example, the result shows “Hard-coded credential in API call”, and we know that Hardcodedcredential.ql query belongs to CWE-798, this security issue can be mapped to the same CWE (798).
-In addition, while FindBugs and ESLint also identify security issues, their scan results only provide descriptions of the security issues.
-We manually associated the descriptions of the security issues in the results with relevant CWE descriptions to determine the specific CWE category to which these security issues belong.
+Python 91 277
+JavaScript 43 175
+Total 134 544
+For each code snippet, we used CWEs to classify the security issues identified by the static analysis tools.
+Each CWE has a unique ID and a set of related descriptions, including its potential impact and how to detect and fix the CWE (The MITRE Corporation, 2023b).
+To more accurately identify the type of security issue with the corresponding CWE, we did not directly accept the provided information of the static analysis tools (such as the CWE number reported by Bandit).
+Instead, we manually identified the type of security issue at the specific location reported in the warning message of the static analysis tool and matched it with the corresponding CWE number.
 Initially, two authors independently matched each description of the security issue with a CWE ID.
@@
 This process continued until all the descriptions of the security issues in the results were matched with CWE IDs.
-Table 7 shows the list of manually matched CWE IDs and warning messages from CodeQL, ESLint, and FindBugs results.
-In the final stage, we performed a statistical analysis of CWE weaknesses in 156 code snippets that contained security weaknesses.
+In the final stage, we performed a statistical analysis of CWE weaknesses in 134 code snippets that contained security weaknesses.
 Results.
-Table 8 shows the distribution of CWEs in the code snippets, including the number of code snippets that contain a certain CWE (Related Snippets) and the total number of occurrences (Frequency) of the CWE in the code snippets (we put those CWEs whose Frequency = 1 in “Others”).
-Note that one related code snippets may contain multiple instances of specific CWE.
-In total, we found 600 CWEs in 435 code snippets.
-These security weaknesses were related to 42 types of CWE, indicating that developers face a variety of security weaknesses when using Copilot.
-CWE-78: OS Command Injection is the most frequently occurred CWE, as it was detected in 15 code snippets (representing 14% of the security weaknesses), followed by CWE-330: Use of Insufficiently Random Values, CWE-703: Improper Check or Handling of Exceptional Conditions , CWE-400: Uncontrolled Resource Consumption and CWE-502: Deserialization of Untrusted Data.
-Some CWEs appeared less frequently, such as CWE-95: Eval Injection, and CWE-22: Improper Limitation of a Pathname to a Restricted Directory.
-Additionally, many CWEs occur with a probability of less than 1%, for example, CWE-176: Improper Handling of Unicode Encoding, CWE-312: Cleartext Storage of Sensitive Information, and CWE-326: Inadequate Encryption Strength.
-This indicates that the types of security issues are closely related to the specific scenarios in which developers use Copilot and make the security issues become apparent, emphasizing the importance of maintaining vigilance and caution when programming.
-Table 7. The warning messages of the scan results by CodeQL, ESLint, and FindBugs manually matched with corresponding CWE IDs
-Tool Warning Message CWE-ID
-CodeQL
-Reflected server-side cross-site scripting
-CWE-79
-CodeQL
-Flask app is run in debug mode
-CWE-215
-CodeQL
-Clear-text logging of sensitive information
-CWE-532
-CodeQL
-Clear-text storage of sensitive information
-CWE-312
-CodeQL
-Information exposure through an exception
-CWE-209
-CodeQL
-Request without certificate validation
-CWE-295
-CodeQL
-Assignment to constant
-CWE-682
-CodeQL
-Log injection
-CWE-117
-CodeQL
-Identical operands
-CWE-570
-CodeQL
-Incomplete string escaping or encoding
-CWE-176
-CodeQL
-DOM text reinterpreted as HTML
-CWE-79
-CodeQL
-Arbitrary file write during archive extraction (“Zip Slip”)
-CWE-22
-CodeQL
-Hard-coded credential in API call
-CWE-798
-CodeQL
-Hard-coded credentials
-CWE-798
-CodeQL
-Uncontrolled data used in path expression
-CWE-22
-CodeQL
-Resource not released in destructor
-CWE-416
-CodeQL
-Missing Dispose call on local IDisposable
-CWE-690
-CodeQL
-Use of the return value of a procedure
-CWE-252
-CodeQL
-Dereferenced variable may be null
-CWE-476
-ESLint
-Generic Object Injection Sink
-CWE-502
-ESLint
-Function Call Object Injection Sink
-CWE-20
-ESLint
-Unsafe Regular Expression
-CWE-20
-ESLint
-Variable Assigned to Object Injection Sink
-CWE-95
-FindBugs
-M S Dm: Hardcoded constant database password
-CWE-798
-FindBugs
-H I Dm: Found reliance on default encoding
-CWE-116
-FindBugs
-M D ICAST: Integral division result
-CWE-682
-FindBugs
-H C IL: There is an apparent infinite loop
-CWE-835
-FindBugs
-M B RV:Exceptional return value of java.io.File.mkdirs()ignored
-CWE-252
-FindBugs
-M D NP: Possible null pointer dereference
-CWE-476
-FindBugs
-M B ODR: DatabaseOperate.readDatabase() may fail to close Connection
-CWE-404
+Table LABEL:Distribution_of_CWEs_in_code_snippets shows the distribution of CWEs in the code snippets and the total number of occurrences (Frequency) of the CWE in the code snippets (we put those CWEs whose Frequency = 1 in “Others”).
+In total, we found 544 CWEs in 452 code snippets.
+These security weaknesses were related to 38 CWEs, indicating that developers face a variety of security weaknesses when using Copilot.
+CWE-330: Use of Insufficiently Random Values is the most frequently occurring CWE, as it represents 23.3% of the security weaknesses, followed by CWE-94: Improper Control of Generation of Code (‘Code Injection’) , CWE-78: OS Command Injection and CWE-95: Improper Neutralization of Directives in Dynamically Evaluated Code (‘Eval Injection’).
+Some CWEs appear less frequently, for example CWE-117: Improper Output Neutralization for Logs only occurred twice.
+Furthermore, many CWEs occur with a frequency of less than 1%, for example, CWE-22: Improper Limitation of a Pathname to a Restricted Directoryand CWE-770: Allocation of Resources Without Limits or Throttling.
+This indicates that the types of security issue are closely related to the specific scenarios in which developers use Copilot, emphasizing the importance of maintaining vigilance and caution when programming.
+Besides, Table 6 presents the top 5 CWEs that appear in Python and JavaScript code snippets.
 Table 5. Distribution of CWEs in code snippets
-CWE-ID # Related Code Snippets Frequency of Specific CWE Percentage
-CWE-78
-15
-84
-14.0%
-CWE-330
-34
-81
-13.5%
-CWE-703
-20
-78
-13.0%
-CWE-398
-11
-60
-10.0%
-CWE-502
-15
-56
-9.3%
-CWE-400
-22
-50
-8.3%
-CWE-20
-10
-19
-3.1%
-CWE-252
-2
-13
-2.1%
-CWE-259
-6
-13
-2.1%
-CWE-404
-3
-13
-2.1%
-CWE-451
-3
-13
-2.1%
-CWE-682
-2
-13
-2.1%
-CWE-116
-3
-11
-1.8%
-CWE-690
-3
-9
-1.5%
-CWE-798
-5
-9
-1.5%
-CWE-561
-4
-8
-1.3%
-CWE-95
-4
-8
-1.3%
-CWE-22
-4
-6
-1.0%
-CWE-327
-5
-5
-¡1%
-CWE-532
-3
-5
-¡1%
-CWE-563
-4
-4
-¡1%
-CWE-605
-4
-4
-¡1%
-CWE-89
-3
-4
-¡1%
-CWE-295
-1
-3
-¡1%
-CWE-476
-1
-3
-¡1%
-CWE-775
-2
-3
-¡1%
-CWE-117
-1
-2
-¡1%
-CWE-209
-1
-2
-¡1%
-CWE-215
-2
-2
-¡1%
-CWE-416
-1
-2
-¡1%
-CWE-570
-2
-2
-¡1%
-CWE-664
-1
-2
-¡1%
-CWE-676
-2
-2
-¡1%
-CWE-79
-1
-2
-¡1%
-CWE-94
-2
-2
-¡1%
-Others
-=1
-¡1%
-42 Types Total: 600
+CWE-ID Frequency of Specific CWE Percentage
+CWE-330 127 23.3%
+CWE-94 115 21.1%
+CWE-78 81 14.9%
+CWE-95 30 5.5%
+CWE-772 23 4.2%
+CWE-89 15 2.8%
+CWE-457 15 2.8%
+CWE-259 14 2.6%
+CWE-670 13 2.4%
+CWE-396 11 2.0%
+CWE-617 8 1.5%
+CWE-628 7 1.3%
+CWE-563 7 1.3%
+CWE-561 7 1.3%
+CWE-798 6 1.1%
+CWE-703 6 1.1%
+CWE-502 6 1.1%
+CWE-20 6 1.1%
+CWE-312 5 0.9%
+CWE-605 4 0.7%
+CWE-295 4 0.7%
+CWE-252 4 0.7%
+CWE-367 3 0.5%
+CWE-327 3 0.5%
+CWE-185 3 0.5%
+CWE-79 2 0.4%
+CWE-682 2 0.4%
+CWE-665 2 0.4%
+CWE-400 2 0.4%
+CWE-215 2 0.4%
+CWE-200 2 0.4%
+CWE-117 2 0.4%
+CWE-116 2 0.4%
+Others =1 ¡1%
+38 Types Total: 544
+Table 6. Top 5 CWEs in Python and JavaScript
+Rank CWE Type
+Python JavaScript
+1 CWE-330 (Use of Insufficiently Random Values Weakness) CWE-94 (Improper Control of Generation of Code)
+2 CWE-78 (Improper Neutralization of Special Elements used in an OS Command) CWE-95 (Improper Neutralization of Directives in Dynamically Evaluated Code)
+3 CWE-772 (Missing Release of Resource after Effective Lifetime) CWE-563 (Assignment to Variable without Use)
+4 CWE-89 (Improper Neutralization of Special Elements used in an SQL Command) CWE-20 (Improper Input Validation)
+5 CWE-259 (Use of Hard-coded Password) CWE-185 (Incorrect Regular Expression)
 4.3. RQ3: How many security weaknesses belong to the CWE Top-25?
 Approach.
-The code in our collected dataset was generated between June 2021 and June 2023.
-To compare whether the security issues in Copilot generated code are widespread in this period, we chose MITRE 2022 CWE Top-25 list (The MITRE Corporation, 2022) as our baseline.
-Then, we compared the CWEs obtained in RQ2 with the CWE Top-25.
+The code snippets in our collected dataset were mainly generated in 2022 and 2023.
+To compare whether the security issues in Copilot-generated code are widespread in this period, we chosed MITRE 2023 CWE Top-25 list (The MITRE Corporation, 2023a) as our baseline.
+In addition, MITRE maintains another list of “Stubborn Weaknesses in the CWE Top 25” (The MITRE Corporation, 2023d), which includes those CWEs that have consistently appeared in the Top 25 most dangerous software weaknesses of the past five years (2019-2023).
+We compared the CWEs obtained in RQ2 with the 2023 CWE Top 25 and “Stubborn Weaknesses” in the CWE Top 25.
 Results.
 The distribution of CWEs found compared to the MITER list is shown in Table 7.
-The results show that the CWE weaknesses present in the code generated by Copilot belong to eleven CWE types included in the MITER CWE Top-25 list.
+The results show that the CWE weaknesses present in the code generated by Copilot belong to eight CWEs.
+We highlight those CWEs that are “Stubborn Weaknesses” (SW) in bold.
 This means these are present issues and are currently among the most common and serious security weaknesses in practice.
-It is worth noting that the 237 security issues present in the code snippet correspond to these 11 CWEs, while another 31 CWEs cover the remaining 363 security issues.
+It is worth noting that the 218 security issues present in the code snippet correspond to these eight CWEs, while another 30 CWEs cover the remaining 326 security issues.
 This indicates that the CWE Top-25 weaknesses are also prevalent in the code generated by Copilot.
-Therefore, developers using Copilot must pay close attention to these weaknesses and take appropriate measures to prevent them before they are integrated into their codebase.
-At the same time, we can see that CWE-78: OS Command Injection is the most frequently occurring weakness from the Top-25 security weaknesses, ranking sixth in the Top-25 list and first in our RQ2 results.
-Although CWE-400: Uncontrolled Resource Consumption is ranked towards the end of the Top-25 list, it is one of the weaknesses with a high occurrence frequency.
-Some CWEs with a higher ranking in the Top-25 list do not appear frequently in Copilot generated code, such as CWE-79: Cross-site Scripting and CWE-89: SQL Injection.
-Table 9. The CWEs that belong to the 2022 CWE Top-25 list
-CWE-ID Description # Related Snippets Frequency
-CWE-78
-OS Command Injection
-15
-84
-CWE-502
-Deserialization of Untrusted Data
-15
-56
-CWE-400
-Uncontrolled Resource Consumption
-22
-50
-CWE-20
-Improper Input Validation
-10
-19
-CWE-798
-Use of Hard-coded Credentials
-5
-9
-CWE-22
-Improper Limitation of a Pathname to a Restricted Directory
-4
-6
-CWE-89
-SQL Injection
-3
-4
-CWE-476
-NULL Pointer Dereference
-1
-3
-CWE-94
-Code Injection
-2
-2
-CWE-416
-Use After Free
-1
-2
-CWE-79
-Cross-site Scripting
-1
-2
-Total 237
+At the same time, we can see that CWE-78: OS Command Injection occurs more frequently in Copilot-generated code.
+CWE-94: Improper Control of Generation of Code is in the list of the CWE Top-25 (The MITRE Corporation, 2023a), and meanwhile it is also one of the weaknesses with a high occurrence frequency in our results.
+Some CWEs with a higher ranking in the Top-25 list do not frequently appear in Copilot-generated code, such as CWE-79: Cross-site Scripting and CWE-89: SQL Injection.
+Table 7. The CWEs that belong to the 2023 CWE Top-25 list
+CWE-ID Description Is SW # Related Snippets Frequency
+CWE-94 Improper Control of Generation of Code No 34 115
+CWE-78 Improper Neutralization of Special Elements used in an OS Command Yes 18 81
+CWE-798 Use of Hard-coded Credentials Yes 3 6
+CWE-502 Deserialization of Untrusted Data Yes 2 6
+CWE-20 Improper Input Validation Yes 3 6
+CWE-79 Improper Neutralization of Input During Web Page Generation Yes 2 2
+CWE-276 Incorrect Default Permissions No 1 1
+CWE-22 Improper Limitation of a Pathname to a Restricted Directory Yes 1 1
+Total 218
 5. Discussion
@@
 RQ1: How secure is the code generated by Copilot?
-Among the 435 code snippets generated by Copilot, we found that 35.8% of these code snippets contain security weaknesses.
-Those weaknesses appear in all six top-used programming languages supported by Copilot.
-Furthermore, when it comes to the occurrence of security issues in code snippets of different programming languages, it is important to analyze them in conjunction with the popularity of the languages (mend.io, 2023).
-In code snippets written in languages like Python and JavaScript, which are frequently used with Copilot, there may be a slightly higher number of security issues.
-However, overall the proportion of security issues across these six languages ranges from 25% to 45%, showing no significant difference.
-Besides, we also found that CWE-502: Deserialization of Untrusted Data and CWE-400: Uncontrolled Resource Consumption problems mainly appeared in code snippets written in Python and JavaScript.
-This could be attributed to certain features that made their code more flexible, such as dynamic typing and dynamic interpretation.
-Therefore, developers should pay special attention to the security of their JavaScript and Python-generated code, taking appropriate measures to validate input data and manage resources effectively to minimize security risks.
-The results of RQ1 suggest that in practical production, although Copilot can help developers write code faster and increase productivity, additional security assessments and fixes are also required to ensure that the generated code does not introduce potential security risks.
+Among the 452 code snippets generated by Copilot, we found that 29.6% of these code snippets contain security weaknesses.
+Although in our results, the proportion of security issues in Python code snippets is higher than in JavaScript, there is no significant difference.
+Using Copilot to write code in Python or JavaScript can generally lead to security issues.
+This could be attributed to features that made their code more flexible, such as dynamic typing and interpretation.
+From a Copilot point of view, the generated code does not need to reason about the whole program because it is dynamically typed, consequently, it does not need to read the flow of the whole program to generate working suggestions.
+In summary, developers should pay special attention to the security of Copilot-generated code, taking appropriate measures to validate input data and manage resources effectively to minimize security risks.
+The results of RQ1 suggest that in practical production, Copilot can help developers write code faster and increase productivity (GitHub, 2024; Perry et al., 2022).
+Still, additional security assessments and fixes are also required to ensure that the generated code does not introduce potential security risks.
 RQ2: What security weaknesses are present in the code snippets generated by GitHub Copilot?
-After conducting a security evaluation of 425 code snippets generated by Copilot, a total of 600 security weaknesses were identified, involving 42 CWE types, which is around 10% of the CWEs (439 CWE types) in software development (The MITRE Corporation, 2023).
-This may be due to the reason that Copilot generates code in different programming languages and application scenarios, and a wide variety of application scenarios may lead to various types of security issues.
+After conducting a security evaluation of 452 code snippets generated by Copilot, a total of 544 security weaknesses were identified, involving 38 CWEs, which is around 10% of the CWEs (439 CWEs) (The MITRE Corporation, 2023c).
+This may be because Copilot generates code in different programming languages and application scenarios.
 In addition, since the Copilot base model (Codex) is trained on publicly available data that potentially contain various types of security weaknesses, this can lead to the presence of multiple CWEs in the generated code by Copilot.
-This set of 42 CWE types covers many types of security issues, and Table 10 shows the types of security issues that these 42 CWEs are relevant to.
-Table 10. The CWEs and Types of Security issues
+This set of 38 CWEs covers many security issues, and Table 8 shows the types of security issues to which these CWEs are relevant.
+Table 8. CWEs and security issue types in Copilot-generated code
 Type of Security Issue Relevant CWEs
-Web security issue
-CWE-79, CWE-94, CWE-690, CWE-732
-Access control issue
-CWE-252, CWE-259, CWE-327, CWE-338
-Input validation and representation issue
-CWE-20, CWE-89, CWE-116
-Command injection issue
-CWE-78, CWE-563, CWE-835
-SQL injection issue
-CWE-89, CWE-95
-File handling issue
-CWE-22, CWE-570
-Insecure storage
-CWE-502, CWE-775
-Improper error handling
-CWE-398, CWE-400
-Encryption issue
-CWE-327
-Memory management issue
-CWE-416
-Buffer errors
-CWE-476
-Insecure random number
-CWE-330
-Incorrect type conversion
-CWE-703
+Data Neutralization Issues CWE-94, CWE-78, CWE-117, CWE-89
+Resource Management Errors CWE-772, CWE-502
+Error Conditions, Return Values CWE-396, CWE-617
+Bad Coding Practices CWE-628, CWe-563
+Credentials Management Errors CWE-798
+Information Management Errors CWE-312, CWE-215
+Authentication Errors CWE-295
+Concurrency Issues CWE-367
+File Handling Issues CWE-22
 The diversity of security weaknesses indicates that developers using Copilot face various security risks.
 These risks are diverse, covering different development environments and application scenarios.
-At the same time, it also reflects the inevitability of security weaknesses in Copilot generated code.
-Developers need to have corresponding security awareness and skills and take appropriate security measures to avoid these risks in a timely and targeted manner.
-In addition, we can see that developers often encounter CWE-78: OS Command Injection, CWE-330: Use of Insufficiently Random Values, andCWE-703: Improper Check or Handling of Exceptional Conditions, CWE-502: Deserialization of Untrusted Data, and CWE-400:Uncontrolled Resource Consumption, which appear in multiple code snippets and have a high frequency of occurrence.
-This can remind developers to take timely and targeted security measures to mitigate these risks.
+In addition, we find that CWEs in Python are mainly related to data processing and system calls.
+In contrast, CWEs in JavaScript are often associated with dynamic code generation problems and security issues in web development.
+This may be because the two languages are designed differently.
+JavaScript is more commonly used for web development, while Python is more widely used in areas such as data processing and scientific computing.
+Overall, while Python and JavaScript differ in some common types of security weaknesses, they require developers to be aware of and take timely and targeted security measures to mitigate these risks.
 For example, developers should perform adequate validation of user inputs.
-In addition to this, it is also necessary to restrict the program’s permissions so that they only access essential resources.
-The results of RQ2 reveal the security weaknesses that developers may encounter in an actual production environment and their frequency of occurrence, which can help developers be aware of security aspects of code generated by Copilot and take appropriate measures to address the security weaknesses in an informed manner.
+It is also necessary to restrict the program’s permissions so that they only access essential resources.
+The results of RQ2 reveal the security weaknesses that developers may encounter in an actual production environment and their frequency of occurrence, which can help developers to be aware of the security aspects of the code generated by Copilot and to take appropriate measures to address the security weaknesses in an informed manner.
 RQ3: How many security weaknesses belong to the CWE Top-25?
-As shown in Table 9, eleven of the CWEs in Copilot generated code can be found in the 2022 CWE Top-25 list, covering more than 237 security issues (39.5% of 600 identfied CWEs) in our dataset.
-This indicates that the commonly acknowledged top 25 weaknesses in software development, which are considered the most prevalent and dangerous, are also prevalent in the code generated by Copilot.
-Therefore, developers need to pay special attention to these frequently occurring weaknesses and take corresponding measures to avoid and fix them.
-We also observed that some vulnerabilities from the CWE top-25 list were not detected in our scans, indicating that Copilot may sanitize and prevent specific weaknesses from being suggested to developers.
-GitHub is gradually enhancing the security of Copilot and its underlying model (Codex) (GitHub, 2023).
-We also identified 31 security weaknesses in the code that do not belong to the CWE Top-25 list.
+As shown in Table 7, eight of the CWEs in the Copilot-generated code can be found in the 2023 CWE Top-25 list, covering more than 218 security issues (40.1% of 544 identified CWEs) in our dataset.
+This indicates that the commonly acknowledged CWE Top-25 weaknesses in software development, which are considered the most prevalent and dangerous, are also prevalent in the code generated by Copilot.
+Therefore, developers using Copilot must pay close attention to these weaknesses and take appropriate measures to prevent them before they are integrated into their code.
+Meanwhile, we found that some higher-ranked CWEs in the CWE Top-25 list, such as CWE-89: SQL Injection, did not frequently appear in our scan results.
+This could be because the generated code does not establish connections with external resources (e.g., databases).
+Besides, we also observed that some vulnerabilities from the CWE Top-25 list were not detected in our scans, indicating that Copilot may sanitize and prevent specific weaknesses from being suggested to developers.
+GitHub is gradually improving the security of Copilot and its underlying model (Codex) (GitHub, 2023).
+However, it has also been found that Copilot’s security layer can potentially handle some CWEs better than others, leaving applications vulnerable to some critical vulnerabilities (Majdinasab et al., 2024).
+We also identified 30 security weaknesses in the code that do not belong to the CWE Top-25 list.
 Although these less common security weaknesses may not be as widespread as CWE Top-25, attackers can still exploit them.
@@
 5.2. Implications
-Code Snippets with Security Weaknesses: In practical production, practitioners often use Copilot to generate code in six languages: Python, JavaScript, Java, C++, Go, and C#.
-These languages all inevitably produce security weaknesses.
-We conjecture that practitioners using Copilot will likely encounter security weaknesses, regardless of the programming language used, and security checks are mandatory.
-When using Copilot, practitioners should conduct their own assessment of the generated code with the support of security analysis tools.
-They should exercise extreme caution when attempting to rely entirely on Copilot’s behavior, especially for the most commonly used languages with Copilot: Python and JavaScript.
-Types of Security Weaknesses in Copilot Generated Code: Practitioners using Copilot may encounter a variety of security weaknesses.
-Our results show that these weaknesses are related to over 40 CWEs.
-This finding indicates that there are diverse security scenarios in production, and practitioners must have the corresponding security awareness and skills and adopt multiple security prevention measures to address security risks so that they do not simply accept vulnerable code suggestions.
-At the same time, our study reveals the frequency of related CWEs.
-When using Copilot to generate code, practitioners should pay particular attention to specific weaknesses, such as CWE-78: OS Command Injection, CWE-330: Use of Insufficiently Random Values, CWE-502: Deserialization of Untrusted Data, CWE-703: Improper Check or Handling of Exceptional Conditions, and CWE-400: Uncontrolled Resource Consumption.
-Our findings can assist practitioners in proactively preventing and addressing security issues in a targeted manner.
-The CWEs in Copilot Generated Code from the Top-25 CWE List: Common security weaknesses in software development are also prevalent in code generated by Copilot.
-As a good practice, developers can use the CWE Top-25 list as a guide to understand which security weaknesses are most common and dangerous in the generated code and take appropriate measures to improve the code security.
-Additionally, the CWE Top-25 provides a standardized approach for security assessment, and developers can also use it to conduct security audits of the code generated by Copilot.
-Developers should also follow best practices and use code analysis tools (static, dynamic, or hybrid) to check the suggested code by Copilot (or any code generation tools) before integrating any code suggestions.
-Such tools can safeguard the code and help in discovering weaknesses early.
+Importance of continuous security analysis of Copilot-generated code.
+We conjecture practitioners using Copilot will likely encounter security weaknesses, regardless of the programming language used.
+Our study results reflect the inevitability of security weaknesses in Copilot-generated code, and there are a variety of security weaknesses in Copilot-generated code.
+Practitioners must be aware of the diverse security scenarios in production and adopt multiple security prevention measures to address security risks before accepting vulnerable code suggestions.
+Maintaining a rigorous process of security checks concurrently with code generation can identify potential vulnerability risks and rectify the security issues in time.
+Developers should follow the best practices for using code generation tools and always check the code suggestions generated by Copilot (or any code generation tools).
+For example, developers can establish a gated check-in build process for checking and preventing security issues when committing code generated by Copilot.
+Initially, we can turn to automated tools to continuously scan the Copilot-generated code for known security issues, such as CWEs.
+Recognizing that these automated tools may not always detect all security issues, especially newly emerging ones or those that are context-specific, it is necessary to conduct a subsequent manual assessment, including manual security code review for Copilot-generated code.
+By embracing this combined strategy for continuous security analysis, we can ensure a robust security shield for the code committing process with Copilot-generated code.
+Prevention of security issues in Copilot-generated code.
+According to our study results, we provide the following suggestions on how to prevent potential security issues in Copilot-generated code: (1) Targeted security countermeasures: Based on the frequency of related CWEs in Copilot-generated code, practitioners can proactively prevent and address security issues in a targeted manner.
+When using Copilot to generate Python or JavaScript code, developers should focus on dedicated security weaknesses (CWEs) in Copilot-generated code of different programming languages (see Table 6).
+Furthermore, a recent study shows that security weaknesses appear in certain code suggestions, but not all (Majdinasab et al., 2024); consequently, developers should carefully select potentially more secure suggestions, with the assistance of tools, that do not expose the code to vulnerabilities.
+(2) Standardized security assessment: Common security weaknesses in software development are also prevalent in the code generated by Copilot.
+As a good practice, developers can use the CWE Top-25 list as a guide to understand which security weaknesses are most common and dangerous in the generated code and follow the mitigation measures for related CWEs provided by the MITRE (The MITRE Corporation, 2023b).
+Additionally, the CWE Top-25 provides a standardized approach for security assessment, and developers can also use it as a guide to perform security audits of the code generated by Copilot.
+(3) Enhancing prompt engineering to generate security code: The instruction tuning schemes in code generation not only impacts the utility of the code but also its security.
+By combining the security fine-tuning with standard instruction tuning, joint optimization of security and utility can be promoted (He et al., 2024).
+We should incorporate security considerations from the initial stages of code generation.
+For specific security scenarios that are prone to CWEs, we can improve the security of code by enhancing prompt engineering, e.g., with the prompt patterns proposed in (White et al., 2023).
 6. Threats to Validity
@@
 Note that we did not consider internal validity threats since we did not investigate any relationships between variables and results.
-Construct Validity is the degree to which a measurement can explain the theoretical structure and characteristics of the measurement, reflecting the extent to which the studied operational measures truly represent the researcher’s ideas and the content investigated based on the research questions.
-This study has three threats to the construct validity: (1) Using the keyword-based search – We used a keyword-based search to collect relevant code snippets from GitHub.
-The results obtained through the keyword-based search may not cover all code snippets generated by Copilot on GitHub.
+Construct Validity: This study has three threats to construct validity: (1) Using keyword-based search – We used a keyword-based search to collect relevant code snippets from GitHub.
+The results obtained through the keyword-based search may not cover all the code snippets generated by Copilot on GitHub.
 We tried to mitigate this threat by constantly and iteratively refining the keywords and using synonyms.
@@
 Since this process was manually done, it may have been influenced by personal bias.
-At the same time, we assumed that all code files contained in projects declared to be written by Copilot in markdown files were generated by Copilot, and we included them in the research data.
-This has an impact on the construct validity, as we could not have excluded the possibility of human-written code files in these projects.
-(3) Manual association of CWEs – We manually associated some warning messages prompted by static analysis tools with CWEs.
-Some warning messages could be associated with multiple CWEs, while we only focused on assigning one most suitable CWE for each warning message.
-Personal bias may occur with this step, impacting the final association, and we mitigated the bias by two authors conducting the association with the assessment by a security expert.
-External Validity refers to the extent to which research results can be generalized and the degree to which people outside the investigated cases are interested in the research results.
-It indicates whether the research results are representative and can be validated in similar contexts.
-Our dataset consists of Copilot generated code snippets collected from open-source projects on GitHub.
+In this regard, two authors conducted the pilot experiment independently to minimize the impact on the construct validity.
+(3) Manual association of CWEs – We manually associated the warning messages reported by the static analysis tools with a particular CWE, which may introduce personal subjective bias, threatening the construct validity.
+We employed two measures to mitigate this threat.
+First, since the list of CWEs is a tree structure with interconnections between them, we first matched the warning messages to a higher-level CWE, and further checked whether we can match the warning messages to a lower-level CWE with more specific definition.
+Second, to mitigate the personal bias, two authors independently assigned each security issue description a CWE ID.
+In case of disagreement, the two authors discussed it with the assessment by a third author (a security expert).
+External Validity: Our dataset consists of Copilot-generated code snippets collected from open-source projects on GitHub.
 During the filtering process, we excluded code that utilized Copilot to solve algorithmic problems, aiming to ensure that the collected data genuinely reflected real-world production environments.
-Due to the reason that the data from GitHub is not diversified enough, there are many code snippets from Game projects.
+Since the data from GitHub are not diversified enough, we had a higher number of code snippets originating from the Game projects.
+This could result in a lack of comprehensiveness in the security scenarios involved.
 The peculiarity of the data source may make the dataset incomplete, thereby threatening the external validity of the results.
-However, GitHub is one of the largest code hosting platforms in the world, with hundreds of millions of public code repositories, and is popular among developers and the technology community.
-The code snippets obtained from GitHub are diverse, which mitigates this threat.
 Furthermore, we acknowledge the need to collect more diverse code snippets from different platforms to increase the generalizability of the results.
 We will consider adopting more diversified ways or platforms to collect code.
-Additionally, due to the limitations of static analysis tools themselves, these tools could not scan all CWEs, and there is a degree of false-positives in the scanning results (as the case with static analysis tools (Sui et al., 2020; Kang et al., 2022)).
-Although we attempted to use two static analysis tools to increase the coverage of weaknesses, these tools may have limited abilities in analyzing some CWEs with specific error rates, which may impact the completeness and correctness of the results.
-Reliability refers to the extent to which a specific research method can produce consistent results.
-We used multiple automated static analysis tools to analyze the Copilot generated code snippets to improve security weaknesses detection.
+Additionally, due to the limitations of static analysis tools themselves, these tools could not scan all CWEs, and there is a degree of false positives in the scan results (as the case with static analysis, in general, (Sui et al., 2020; Kang et al., 2022)).
+Although we used two widely used static analysis tools to increase the comprehensiveness of the scans and manually checked the results of the tool scans, the results may suffer from incompleteness.
+Reliability: We used multiple automated static analysis tools to analyze the Copilot-generated code snippets to improve security weaknesses detection.
 Developers have widely used these automated tools.
 The querying mechanism of these tools ensures that the scan results remain consistent when used multiple times.
-In addition, we performed two rounds of scanning with two tools for security checks on each code snippet, intending to complement the results of one tool with the other one.
+In addition, we performed two rounds of scanning with two tools for security checks on each code snippet, intending to complement the results of one tool with the other.
 By implementing these measures, we believe that our research results are reliable and these threats to reliability are mitigated.
 7. Conclusions
-Automatic code generation and recommendation has been an active research area due to the advancement of AI and specifically LLMs.
-AI code generation tools such as Copilot can greatly improve the development efficiency of programmers, but they can also introduce vulnerabilities and security risks.
-In this paper, we present the results of an empirical study to analyze security weaknesses in Copilot generated code found in public GitHub projects.
-We identified 435 code snippets generated by Copilot from GitHub projects and analyzed those snippets for security weaknesses using static analysis tools.
+Automatic code generation and recommendation have been an active research area due to the advancement of AI and, specifically, LLMs.
+AI code generation tools, such as Copilot, can significantly improve developers’ development efficiency, but can also introduce vulnerabilities and security risks.
+In this paper, we present the results of an empirical study to analyze security weaknesses in Copilot-generated code found in public GitHub projects.
+We identified 452 code snippets generated by Copilot from GitHub projects and analyzed those snippets for security weaknesses using static analysis tools.
 This study aims to help developers understand the security risks of weaknesses introduced in the code generated by Copilot (and potentially similar code generation tools).
-Our results show: (1) 35.8% of the 435 Copilot generated code snippets contain security weaknesses, spreading across six programming languages.
-(2) The detected security weaknesses are diverse in nature and are associated with 42 different CWEs.
-The CWEs that occurred most frequently are CWE-78: OS Command Injection, CWE-330: Use of Insufficiently Random Values, and CWE-703: Improper Check or Handling of Exceptional Conditions (3) Among these CWEs, 11 appear in the MITRE CWE Top-25 list, demonstrating their commonality, current and severity.
-These are: CWE-78: OS Command Injection, CWE-502: Deserialization of Untrusted Data, CWE-400: Uncontrolled Resource Consumption, CWE-89: SQL Injection, CWE-20: Improper Input Validation, CWE-22: Improper Limitation of a Pathname to a Restricted Directory, CWE-94: Code Injection, CWE-476: NULL Pointer Dereference, CWE-798: Use of Hard-coded Credentials, CWE-79: Cross-site Scripting, and CWE-416: Use After Free.
-In the future, we plan to: (1) collect additional code snippets from other open source repositories and industrial projects, and code snippets generated by newer releases of Copilot; (2) analyze and summarize the application scenarios of these code snippets, studying how practitioners use Copilot and fix the issues in development; and (3) compare the results with other emerging Generative AI code generation tools such as CodeWhisperer, aiXcoder, and Code Llama.
+Our results show that (1) 29.8% of the 452 Copilot-generated code snippets contain security weaknesses.
+Developers have a high risk of raising security issues when using Copilot, regardless of the programming language, so security checks are necessary.
+(2) The detected security weaknesses are diverse in nature and are associated with 38 different CWEs.
+Developers face a variety of development scenarios and application environments in production and need the appropriate security awareness and skills.
+(3) Among these CWEs, eight appear in the MITRE CWE Top-25 list, and six belong to the Stubborn Weaknesses, demonstrating their severity.
+In the future, we plan to: (1) collect additional code snippets from other open source repositories and industrial projects and code snippets generated by newer releases of Copilot; (2) analyze and summarize the application scenarios of these code snippets, studying how practitioners use Copilot and fix the issues in development; and (3) compare the results with other emerging Generative AI code generation tools such as CodeWhisperer, aiXcoder, and Code Llama.
+Data availability
+We have shared the link to our dataset in the reference (Fu et al., 2024).
+Acknowledgments
+This work is funded by the National Natural Science Foundation of China (NSFC) under Grant No. 62172311 and the Special Fund of Hubei Luojia Laboratory.

### v3

@@
-Security Weaknesses of Copilot Generated Code in GitHub
+Security Weaknesses of Copilot-Generated Code in GitHub Projects: An Empirical Study
 Abstract.
-Modern code generation tools, utilizing AI models like Large Language Models (LLMs), have gained popularity for producing functional code.
+Modern code generation tools utilizing AI models like Large Language Models (LLMs) have gained increased popularity due to their ability to produce functional code.
 However, their usage presents security challenges, often resulting in insecure code merging into the code base.
-Evaluating the quality of generated code, especially its security, is crucial.
-While prior research explored various aspects of code generation, the focus on security has been limited, mostly examining code produced in controlled environments rather than real-world scenarios.
-To address this gap, we conducted an empirical study, analyzing code snippets generated by GitHub Copilot from GitHub projects.
-Our analysis identified 452 snippets generated by Copilot, revealing a high likelihood of security issues, with 32.8% of Python and 24.5% of JavaScript snippets affected.
-These issues span 38 different Common Weakness Enumeration (CWE) categories, including significant ones like CWE-330: Use of Insufficiently Random Values, CWE-78: OS Command Injection, and CWE-94: Improper Control of Generation of Code.
-Notably, eight CWEs are among the 2023 CWE Top-25, highlighting their severity.
-Our findings confirm that developers should be careful when adding code generated by Copilot and should also run appropriate security checks as they accept the suggested code.
-It also shows that practitioners should cultivate corresponding security awareness and skills.
+Thus, evaluating the quality of generated code, especially its security, is crucial.
+While prior research explored various aspects of code generation, the focus on security has been limited, mostly examining code produced in controlled environments rather than open source development scenarios.
+To address this gap, we conducted an empirical study, analyzing code snippets generated by GitHub Copilot and two other AI code generation tools (i.e., CodeWhisperer and Codeium) from GitHub projects.
+Our analysis identified 733 snippets, revealing a high likelihood of security weaknesses, with 29.5% of Python and 24.2% of JavaScript snippets affected.
+These issues span 43 Common Weakness Enumeration (CWE) categories, including significant ones like CWE-330: Use of Insufficiently Random Values, CWE-94: Improper Control of Generation of Code, and CWE-79: Cross-site Scripting.
+Notably, eight of those CWEs are among the 2023 CWE Top-25, highlighting their severity.
+We further examined using Copilot Chat to fix security issues in Copilot-generated code by providing Copilot Chat with warning messages from the static analysis tools, and up to 55.5% of the security issues can be fixed.
+We finally provide the suggestions for mitigating security issues in generated code.
 Code Generation, Security Weakness, CWE, GitHub Copilot, GitHub Project
@@
 This approach can save development time and accelerate the software development process.
-Automated code generation tools have always been a topic of active research (Siddiq and Santos, 2022; Li et al., 2022b).
+Automated code generation tools have always been a topic of active research (Siddiq and Santos, 2022; Li et al., 2023).
 Some of the earliest work can be traced back to the 1960s, when Waldinger and Lee proposed a program synthesizer called PROW, which automatically generated LISP programs based on specifications provided by users in the form of a predicate calculus (Waldinger and Lee, 1969).
 As computer languages continued to evolve, more and more programming languages began to support meta-programming, making automated code generation technology more efficient and flexible.
-In recent years, the rapid development of artificial intelligence technologies, particularly machine learning and deep learning models, has accelerated the development of code generation technologies.
+In recent years, the rapid development of artificial intelligence technologies, particularly deep learning models, has accelerated the development of code generation technologies.
 Recent advancements in code generation came with the emergence of Large Language Models (LLMs).
 LLMs are deep learning models trained on a large code/text corpus with powerful language understanding capabilities that can be used for tasks such as natural language generation, text classification, and question-answer systems (Chen et al., 2021).
-Compared to previous deep learning methods, the latest developments in LLMs, such as Generative Pre-trained Transformer (GPT) models, have opened up new opportunities to address the limitations of existing automated code generation technology (Mozannar et al., 2022).
-Currently, LLM-based code generation tools have also been widely applied, such as Codex by OpenAI (OpenAI, [n.
-d.]), AlphaCode by DeepMind (Li et al., 2022a), and CodeWhisperer by Amazon (Becker et al., 2022).
-These models are trained on billions of public open-source lines of code, which includes public code with unsafe coding patterns (He and Vechev, 2023).
-Therefore, code generation tools based on such models can pose code quality issues (Liu et al., 2024), and the code generated by these tools may also suffer from security weaknesses (She et al., 2023).
+Compared to previous deep learning methods, the latest developments in LLMs, such as Generative Pre-trained Transformer (GPT) models, have opened up new opportunities to address the limitations of existing automated code generation technology (Mozannar et al., 2024).
+Currently, LLM-based code generation tools have also been widely applied, such as Codex by OpenAI (OpenAI, 2021), AlphaCode by DeepMind (Li et al., 2022), and CodeWhisperer by Amazon (Becker et al., 2023).
+These models are trained on billions of public open-source lines of code, which include public code with unsafe coding patterns (He and Vechev, 2023).
+Therefore, code generation tools based on such models can pose code quality issues (Liu et al., 2024), and the AI generated code may also suffer from security weaknesses (She et al., 2023).
 For example, GitHub Copilot may produce some insecure code, as its underlying Codex model is pre-trained on untrusted data from GitHub (Brown et al., 2020), which is known to contain buggy programs (Rokon et al., 2020).
-According to the developer security company Snyk, GitHub Copilot may also replicate existing security issues in code to suggest insecure code when the user’s existing codebase contains security issues (Paul Krill, 2024).
+According to the developer security company Snyk, GitHub Copilot may also replicate existing security weaknesses in code to suggest insecure code when the user’s existing codebase contains security weaknesses (Krill, 2024).
 In addition, the code with vulnerabilities generated by these code-generation tools may continue to be used to train the model, thus further generating code with vulnerabilities, leading to a vicious cycle.
-Previous research has studied code generation tools, with more focus on the correctness of the results (Lertbanjongngam et al., 2022; Wong et al., 2022; Dakhel et al., 2022; Pudari and Ernst, 2023), and relatively less attention has been paid to security aspects (Pearce et al., 2022; Siddiq et al., 2022; Perry et al., 2022).
-To the best of our knowledge, potential security weaknesses in practical scenarios have not been fully considered and addressed in previous work, and GitHub Copilot clarifies that “the users of Copilot are responsible for ensuring the security and quality of their code” (GitHub, 2023b).
-Code generation algorithms of Copilot are incentivized to suggest code to be accepted rather than other code qualities, e.g., easy to read and understand, which has an adverse impact for the code quality generated by Copilot (Harding and Kloster, 2024).
-GitHub also provides tools such as CodeQL to help developers scan for security issues in their code.
-To this end, we conducted an empirical study on the security weaknesses of the generated code by GitHub Copilot, which is available on GitHub.
-We chose Copilot as our research subject because it is a commercial instance of AI-assisted programming and has gained much attention and popularity among developers since its launch in 2021.
+Previous research has studied code generation tools, with more focus on the correctness of the results (Lertbanjongngam et al., 2022; Wong et al., 2022; Dakhel et al., 2023; Pudari and Ernst, 2023), and relatively less attention has been paid to security aspects (Pearce et al., 2022; Siddiq et al., 2022; Perry et al., 2023).
+To the best of our knowledge, potential security weaknesses in practical scenarios have not been fully considered and addressed in previous work, and GitHub Copilot clarifies that “the users of Copilot are responsible for ensuring the security and quality of their code” (GitHub, 2023).
+Code generation algorithms of Copilot are incentivized to suggest code to be accepted rather than other code qualities, e.g., easy to read and understand, which has an adverse impact on the code quality generated by Copilot (Harding and Kloster, 2024).
+GitHub also provides tools such as CodeQL to help developers scan for security weaknesses in their code.
+To this end, we conducted an empirical study on the security weaknesses of the generated code by GitHub Copilot and other AI code generation tools (CodeWhisperer and Codeium), which is available on GitHub.
+We chose Copilot as our main research subject because it is a commercial instance of AI programming and has gained much attention and popularity among developers since its launch in 2021.
 The security weaknesses of code generated by Copilot have also gained attention in the research and practice community.
-Furthermore, thousands of developers in the GitHub community have shared their experiences of using Copilot in real-world systems (Dakhel et al., 2022).
-We collected the code generated by Copilot that has been used in projects on GitHub and analyzed the security of the generated code through the lens of a real-world production environment.
+Furthermore, thousands of developers in the GitHub community have shared their experiences of using Copilot in open source projects (Dakhel et al., 2023).
+We also include the code generated by other two popular AI code generation tools, CodeWhisperer and Codeium, for the purpose of increasing the generalizability of this research.
+We then collected the code generated by Copilot and other generation tools that has been used in projects on GitHub and analyzed the security of the generated code through the lens of open source development environment.
 Then, we used static analysis tools to perform security analysis on the collected code snippets and classified the security weaknesses in the code snippets using the Common Weakness Enumeration (CWE).
-Our findings show that: (1) 29.6% of Copilot-generated code snippets have security weaknesses; (2) the security weaknesses are diverse and related to 38 different CWEs, in whichCWE-330: Use of Insufficiently Random Values, CWE-78: OS Command Injection and CWE-94: Improper Control of Generation of Code (’Code Injection’) are the most frequently occurred; and (3) among the 38 CWEs identified, eight CWEs belong to the currently recognized 2023 CWE Top-25.
+Our findings show that: (1) around 30% of code snippets have security weaknesses; (2) the security weaknesses are diverse and related to 43 different CWEs, in whichCWE-330: Use of Insufficiently Random Values, CWE-94: Improper Control of Generation of Code (’Code Injection’) and CWE-79: Cross-site Scripting and are the most frequently occurred; and (3) among the 38 CWEs identified, eight CWEs belong to the currently recognized 2023 CWE Top-25.
 Six CWEs belong to Stubborn Weaknesses in the CWE Top 25.
-The contributions of this work: (1) We curated a dataset of code snippets generated by Copilot that has been used in projects on GitHub (a curated data (Fu et al., 2024) is made available online ) and conducted security checks on them, which can to some extent reflect the frequency of security weaknesses encountered by developers when using Copilot to generate code in actual coding.
-In addition to this, we also categorized the application areas of these code snippets; (2) We extensively checked all possible CWEs in the code snippets and analyzed them.
+(4) It is possible to use Copilot Chat to fix security issues in Copilot-generated code, but the success of Copilot Chat’s fixes varies between CWEs.
+(5) Providing Copilot Chat with a warning message from the static analysis tool resulted in a better fix.
+Using the fix command can fix 19.3% of security issues, while using the enhanced prompt raises the rate to 55.5%.
+Contributions: (1) We curated a dataset of code snippets generated by Copilot that has been used in projects on GitHub (a curated data (Fu et al., 2024) is made available online ) and conducted security checks on them, which can to some extent reflect the frequency of security weaknesses encountered by developers when using Copilot to generate code in actual coding.
+In addition to this, we also categorized the application areas of these code snippets.
+(2) We extensively checked all possible CWEs in the code snippets and analyzed them.
 This can help developers understand the common CWEs caused by using Copilot to generate code in actual coding and how to safely accept the code suggestions provided by Copilot.
-The rest of this paper is structured as follows: Section 2 presents the related work.
+(3) We utilized Copilot Chat to fix code snippets generated by Copilot, demonstrating Copilot’s ability to fix security issues to a certain extent.
+The rest of this paper is structured as follows: Section 2 presents the background and related work.
 Section 3 presents the research questions and the research design of this study.
-Section 4 presents the results of our study, which are further discussed in Section 5.
-The potential threats to validity are clarified in Section 6.
-Section 7 concludes this work with future work directions.
-2. Related Work
-In this section, we present the related work in three aspects, i.e., AI-assisted code generation tools (Section 2.1), security of code generation techniques and LLMs (Section 2.2), and static analysis tools for scanning security weaknesses (Section 2.3).
-2.1. AI-assisted Code Generation Tools
-With the rise of code generation tools integrated with IDEs, many studies have evaluated these systems based on transformer models to better understand their effectiveness in real-world scenarios.
+Section 4 presents the results of our study, which are further discussed in Section LABEL:Discussion.
+The potential threats to validity are clarified in Section LABEL:Threats_to_Validity.
+Section LABEL:Conclusions_and_Future_Work concludes this work with future work directions.
+2. Background and Related Work
+In this section, we present the related work in four aspects, i.e., AI code generation tools (Section 2.1), security of code generation techniques and LLMs (Section 2.2), how AI code generation tools work (Section 2.3), and static analysis tools for analyzing security weaknesses (Section 2.4).
+2.1. AI Code Generation Tools
+With the rise of code generation tools integrated with IDEs, many studies have evaluated these systems based on transformer models to better understand their effectiveness in open source development scenarios.
 Previous research mainly focused on whether the code generated by these tools can meet users’ functional requirements.
@@
 Burak et al.
-(Yetiştiren et al., 2023) evaluated the code quality of AI-assisted code generation tools (GitHub Copilot, Amazon CodeWhisperer, and ChatGPT).
+(Yetiştiren et al., 2023) evaluated the code quality of AI code generation tools.
 They compared the improvements between the latest and older versions of Copilot and CodeWhisperer and found that the quality of generated code had improved.
-In recent years, researchers have also started to focus on the experience of developers when using AI-assisted code generation tools and how the tools can improve productivity by observing their behavior.
+In recent years, researchers have also started to focus on the experience of developers when using AI code generation tools and how the tools can improve productivity by observing their behavior.
 Vaithilingam et al.
-(Vaithilingam et al., 2022) studied how programmers use and perceive Copilot and found that although Copilot may not necessarily improve task completion time or success rate, it often provides a useful starting point.
+(Vaithilingam et al., 2022) studied how developer use and perceive Copilot and found that although Copilot may not necessarily improve task completion time or success rate, it often provides a useful starting point.
 They also noted that the participants had difficulties understanding, editing, and debugging the code snippets generated by Copilot.
 Barke et al.
-(Barke et al., 2022) presented the first theoretical analysis of how programmers interact with Copilot based on the observations of 20 participants.
+(Barke et al., 2023) presented the first theoretical analysis of how developer interact with Copilot based on the observations of 20 participants.
 Sila et al.
@@
 The authors prompted Copilot to generate code for 89 cybersecurity scenarios and evaluated the weaknesses in the generated code.
-They found that 40% of the suggestions in the relevant context contained security-related bugs (i.e., CWE classification from MITRE (The MITRE Corporation, 2023b)).
+They found that 40% of the suggestions in the relevant context contained security-related bugs (i.e., CWE classification from MITRE (MITRE, 2024)).
 Siddiq et al.
@@
 Khoury et al.
-(Khoury et al., 2023) studied the security of the source code generated by the ChatGPT chatbot based on LLMs, and found that ChatGPT was aware of potential weaknesses but still frequently generated some non-robust code.
+(Khoury et al., 2023) studied the security of the source code generated by the ChatGPT chatbot based on LLMs and found that ChatGPT was aware of potential weaknesses but still frequently generated some non-robust code.
 Elgedawy et al.
@@
 Sandoval et al.
-(Sandoval et al., 2022) conducted a security-driven user study, and their results showed that the rate at which AI-assisted user programming produced critical security errors was no more than 10% of the control group, indicating that the use of LLMs does not introduce new security risks.
+(Sandoval et al., 2022) conducted a security-driven user study, and their results showed that the rate at which AI user programming produced critical security errors was no more than 10% of the control group, indicating that the use of LLMs does not introduce new security risks.
 Asare et al.
@@
 The dataset contains 130 Python code samples.
-Different from prior work, we studied the security weaknesses exhibited by code generation tools in a real-world production environment (i.e., GitHub).
-We collected code snippets from GitHub generated by developers using Copilot in daily production as a source of research data, whereas in the Pearce et al.
-(Pearce et al., 2022) study, the research data came from code generated by the authors using Copilot based on natural language prompts related to high-risk network security weaknesses.
+Natella et al.
+(Natella et al., 2024) provide a dataset and conduct an experimental evaluation of the security of three popular LLMs (GitHub Copilot, Amazon CodeWhisperer, and CodeBERT).
+Different from prior work, we studied the security weaknesses exhibited by code generation tools in open source development environment (i.e., GitHub).
+We collected code snippets from GitHub generated by developers using Copilot in daily production as a source of research data, whereas in the study by Pearce et al.
+(Pearce et al., 2022), the research data came from code generated by the authors using Copilot based on natural language prompts related to high-risk network security weaknesses.
 Additionally, Pearce et al.
@@
 Our research results may help developers understand what common CWEs are prone to result from using Copilot to generate code in real coding.
+2.3. How AI Code Generation Tools Work: Copilot as an Example
+Existing automated code generation tools, such as Copilot, CodeWhisperer, and Codeium, are often integrated into IDEs.
+These tools provide developer with real-time code suggestions based on given prompts.
+Those suggestions can range from code snippets to complete functions.
+We use GitHub Copilot as an example to demonstrate how these tools are used.
+When developers write code in their IDE, Copilot continuously scans the program and performs the following actions (Copilot, 2024): (1) complete existing code based on function or variable names; (2) generate code suggestions from inline natural language comments; (3) automatically fill in repetitive code based on the context of existing code; and (4) generate test cases.
+Figure 1 shows an example of Copilot usage in action.
+In Figure 1(a), Copilot automatically suggests the entire function body (shown in gray text) when the function name is provided.
+A user can then accept the suggestion if desired.
+In Figure 1(b), a function description is given in the comments (Lines 1-2), and Copilot provides code suggestions starting from Line 3.
+After accepting a suggestion, Copilot will continue to suggest the next line of code.
+A user can also view alternative suggestions - Copilot will provide 10 different suggestions at a time, as shown in Figure 2.
+(a) Method 1: Copilot completes a function using the function name
+(b) Method 2: Copilot implements the function in the comment
+Figure 1. Two methods of using Copilot in action
+Figure 2. More suggestions provided by Copilot
 2.4. Security Static Analysis
-Vulnerabilities detection is critical to improve software security and ensure quality  (Iannone et al., 2022).
-Vulnerability can be checked either statically or dynamically.
-Dynamic analysis techniques are more sound and precise but lack coverage (Ernst, 2003).
-On the other hand, static analysis is less precise but offers greater coverage and allows users to analyze programs without the need to execute them (Sui et al., 2020).
-OWASP (OWASP, [n.
-d.]) provides a list of commonly used static analysis tools for security analysis.
-This includes tools like CodeQL, a general-purpose automatic scanning tool; FindBugs, a tool for Java programs; ESLint, a tool for JavaScript programs; Bandit, a tool for Python programs; and GoSec, a tool for Go programs.
-Such tools have been widely used in previous security analysis research  (Tómasdóttir et al., 2018; Pearce et al., 2022; Siddiq et al., 2022).
+Detecting software vulnerabilities is critical to improve security and ensure quality  (Iannone et al., 2022).
+Vulnerabilities can be detected through static, dynamic, or hybrid analysis.
+Static analysis offers greater coverage and allows users to analyze programs without the need to execute them, but it can be unsound and imprecise  (Sui et al., 2020).
+On the other hand, dynamic analysis is more sound and precise (as it captures actual program behaviour at runtime) but can be incomplete and lacks coverage (Ernst, 2003).
+Due to their ease of use and low cost compared to dynamic analysis, static analysis tools are widely used for security analysis (Nunes et al., 2018; Huth and Nielson, 2019).
+There are several security static analysis tools available for different programming languages and application types.
+OWASP and Snyk (OWASP, 2024; Snyk, 2024) provide a list of commonly used static analysis tools for security analysis.
+This includes tools like CodeQL, a general-purpose automatic analysis tool; FindBugs/SpotBugs for Java; ESLint for JavaScript programs; Bandit for Python; and GoSec for Go programs.
+Those tools have been widely used in previous security analysis research  (Siddiq et al., 2022; Tómasdóttir et al., 2018; Pearce et al., 2022).
 Kaur et al.
-(Kaur and Nayyar, 2020) compared static analysis tools for vulnerability detection in scanning C/C++ and Java source code.
+(Kaur and Nayyar, 2020) compared static analysis tools for vulnerability detection in analyzing C/C++ and Java source code.
 Tomasdottir et al.
@@
 It can find weaknesses in a codebase based on known weaknesses/rules.
-In addition, to obtain more comprehensive scan results, we supplemented the scan of code in different languages with static analysis tools (i.e., Cppcheck and Bandi) tailored to specific languages.
+In addition, to obtain more comprehensive analysis results, we supplemented the scan of code in different languages with static analysis tools (i.e., Cppcheck and Bandi) tailored to specific languages.
 3. Research Design
 In this section, we describe our research design in detail.
-In Section 3.1, we first define our Research Questions (RQs), followed by the process of collecting and filtering the code snippets generated by Copilot in Section 3.2.
+In Section 3.1, we first define our Research Questions (RQs), followed by the process of collecting and filtering the code snippets generated by code generation tools such as Copilot in Section 3.2.
 We then explain the security analysis performed on the identified snippets and the process of filtering the raw results generated by static analysis tools in Section 3.3.
+Finally, in Section 3.4, we introduce the process of using Copilot Chat to fix the security weaknesses identified by the static analysis tools.
 3.1. Research Goal and Questions
-This study aims to understand the potential security weaknesses in Copilot-generated code produced in real-world GitHub projects.
-We first collect code snippets generated by Copilot from GitHub projects as our data source, which is the largest source of open-source projects.
-It should be noted that it is not possible to access all the code generated by Copilot in GitHub projects, as there is no direct way to identify if part of a file was generated by Copilot (i.e., source files do not contain any signatures to indicate if Copilot generates the code).
+This study aims to understand the potential security weaknesses in AI-generated code produced in open source GitHub projects.
+We first collect code snippets generated by code generation tools such as Copilot from GitHub projects, which is our data source.
+It should be noted that it is not possible to access all the AI-generated code in GitHub projects, as there is no direct way to identify if part of a file was generated by code generation tools such as Copilot (i.e., source files do not contain any signatures to indicate if Copilot generates the code).
 However, we can identify many code snippets by searching the repository description and the comments provided in the code (see the details in Section 3.2.2).
-We chose to focus on Python and JavaScript code snippets as they are currently the two most popular programming languages used by developers according to a recent survey in 2022 by GitHub (oct, [n.
-d.]), and also the ones most frequently used with GitHub Copilot (Zhang et al., 2023).
-We first analyzed the functionality and application domains of the collected code snippets to get the demographic information of the Copilot-generated code.
-Next, we performed a security analysis on the identified code.
-We selected static analysis because of its coverage and the ability to analyze programs without executing them.
-Dynamic analysis is more sound and precise (as it can reason about the program behaviour) but also lacks coverage (Sui et al., 2020).
-Using static analysis will allow us to run the analysis on a segment of the program (smaller snippet) without needing the full program to be executed.
-Static security analysis tools have been widely used by developers and companies (Majdinasab et al., 2024; Ernst, 2003; Pearce et al., 2022).
-After obtaining the scan results, we manually checked the results to remove false positives reported by the static analysis tools.
-We finally used CWEs to classify the filtered results for further analysis to answer the RQs.
-We conducted this empirical study by following the guidelines of Easterbrook et al.
+We first used keyword-based search to collect projects and files containing code generated by Copilot and other AI code generation tools (CodeWhisperer and Codeium) from GitHub and filtered them manually.
+We subsequently analyzed the functionality and application domains of the collected code snippets to get a comprehensive understanding of the dataset.
+Next, we performed a static security analysis on the identified code.
+After running the analysis and obtaining the analysis results, we manually checked the results to remove false positives reported by the static analysis tools.
+We then used CWEs to classify the filtered results for further analysis to answer our Research Questions (RQs).
+In addition, we used Copilot Chat (GitHub, 2024a) to fix code snippets containing security weaknesses.
+Copilot Chat is an interactive tool launched by OpenAI based on the GPT-4 model, designed to enable natural language-driven coding as part of Copilot.
+Developers can use Copilot Chat to perform tasks such as code analysis and fixing bugs.
+We used the warnings from static analysis tools as prompts and fed them to Copilot Chat in order to get the fixes.
+We then performed a security analysis on the fixed code to evaluate the effectiveness of those fixes.
+To conduct our empirical study, we followed the guidelines of Easterbrook et al.
 (Easterbrook et al., 2008).
-The RQs, their rationale, and the research process of this study (see Figure 1) are detailed in the subsections below.
+The RQs, their rationale, and the research process of this study (see Figure 3) are detailed in the following subsections.
 Figure 3. Overview of the research process
@@
 The answer to RQ1 helps understand the frequency of security weaknesses developers encounter when using Copilot in production.
-RQ2. What security weaknesses are present in the code snippets generated by Copilot?
-Rationale: Copilot-generated code may contain a variety of security weaknesses (Pearce et al., 2022), and developers should conduct a rigorous security review before accepting the generated code suggestions.
-Copilot’s documentation notes the following: “users of Copilot are responsible for ensuring the security and quality of their code (GitHub, 2023b)”.
+We also include the code generated by other two popular AI code generation tools, CodeWhisperer and Codeium, in order to increase the generalizability of this research.
+RQ2. What security weaknesses are present in the code generated by Copilot across different application domains?
+Rationale: Copilot-generated code may contain security weaknesses (Pearce et al., 2022; Majdinasab et al., 2024), and developers should conduct a rigorous security review before accepting the generated code suggestions.
+Copilot’s documentation notes the following: “users of Copilot are responsible for ensuring the security and quality of their code” (GitHub, 2023).
 By identifying common or recurring weaknesses, developers can be more prepared to prevent, mitigate, or fix these security weaknesses.
-The answer to RQ2 can help developers better understand possible security weaknesses in the code generated by Copilot, allowing them to be vigilant about the generated code before it is integrated into their programs.
-RQ3. How many security weaknesses belong to the MITRE CWE Top-25 security weaknesses?
-Rationale: The MITRE list contains the Top 25 most dangerous security weaknesses (The MITRE Corporation, 2023a), offering a benchmark for gauging the severity of weaknesses in the Copilot-generated code (Pearce et al., 2022).
-The answer to RQ3 can help to understand whether the code generated by Copilot contains widely recognized types of security weakness and Copilot’s ability to handle these recent and common weaknesses.
+The answer to RQ2 can help developers better understand possible security weaknesses in the code generated by Copilot and other AI code generation tools, as well as the characteristics of code that contains certain types of security weaknesses, allowing them to be vigilant about the generated code before it is integrated into their code base.
+RQ3. Can Copilot Chat help fix the security weaknesses found in the code generated by Copilot?
+Rationale: The purpose of Copilot Chat is to assist developers by answering questions about their code or providing fixes for issues identified in the code.
+Copilot Chat is integrated with Copilot to form a powerful AI assistant that can help developers build at their speed of thought in the natural language of their choice (GitHub, 2023).
+It is claimed that Copilot Chat can assist in fixing security weaknesses.
+In this RQ, we utilize Copilot Chat to examine whether it can help fix the security weaknesses identified in the Copilot-generated code because most of the security weaknesses in RQ1 and RQ2 come from Copilot.
+Our approach is to utilize Copilot Chat in the loop by collect generated code →analyze the code for security weaknesses →check for fixes in Copilot Chat →re-analyze the code to check if the fix is correct.
+Answering RQ3 can help to understand the capability of Copilot Chat in fixing Copilot-generated code that contains security issues.
 3.2. Data Collection and Filtering
 We chose GitHub as the primary data source for answering our RQs.
-GitHub is widely used by developers.
 As the world’s largest code hosting platform, GitHub contains millions of public code repositories and offers access to a large number of code resources, allowing us to cover multiple programming languages and project types in our study (Cosentino et al., 2017).
-We identify code snippets generated by Copilot from GitHub projects.
+At the same time, thousands of developers in the GitHub community have shared their experiences of using Copilot in open-source projects.
+We identified code snippets generated by AI code generation tools from GitHub projects.
+We chose Copilot as our main research subject as it is a well-known commercial instance of an AI pair programming tool that has gained increased popularity among developers since its launch in 2021.
+Copilot is one of the most popular general-purpose AI code generation tools and only second to ChatGPT in terms of overall AI assistance tools developers use in practice (Overflow, 2024).
+The security drawbacks of code generated by Copilot have been investigated in previous studies (Pearce et al., 2022; McDaniel, 2021; Dow, 2021).
+To increase the generalizability of the research, we also attempted to collect data from two popular AI code generation tools, CodeWhisperer and Codeium - both are general-purpose, AI-driven code generators.
+We did not collect data from specialized AI code generation tools, such as AlphaCode as it focuses on solving complex coding tasks rather than everyday code generation.
 3.2.1. Code Snippets Collection
 Step 1.
-To identify the code snippets generated by Copilot, we first conducted a pilot search to formulate our search keywords.
-First, we used “GitHub Copilot” and “Copilot” as our search keywords.
-As expected, we found that the term “Copilot” is a vague term, referring not only to the code generation tool launched by GitHub, but also to tools in the aviation or telemetry fields.
-Therefore, using the keyword solely may return irrelevant content that may not be related to the use of the tool.
-On the other hand, using “GitHub Copilot” can exclude content unrelated to the code generation tool Copilot and narrow the search scope, which is what we have used to locate the code snippets.
-However, even with this basic search keyword, we still need to carefully filter the search results to ensure that they are truly related to GitHub Copilot.
-Although using “GitHub Copilot” increases the relevance of the results to Copilot, these results are not necessarily the code snippets generated by Copilot.
-It should be noted that many code snippets containing the “GitHub Copilot” keyword in the search results display GitHub Copilot as text.
-Developers may use them to describe their experience using Copilot to generate code or showcase information related to Copilot.
+We utilized GitHub REST API to automatically collect projects and files containing AI-generated code.
+We started with a pilot search on GitHub to formulate our search keywords.
+First, we used the names of the code generation tools, such as “GitHub Copilot” and “CodeWhisperer” as the basic search terms.
+It should be noted that many code snippets in the search results contained the “GitHub Copilot” keyword, but the keyword was not used to declare that the code snippets were generated by Copilot.
+Developers may use them to describe their experience of using Copilot to generate code or showcase information related to Copilot.
 These code snippets are not what we seek as they do not directly relate to the code generated by Copilot.
-Our target is code generated by Copilot, not code snippets containing the keyword “Copilot”.
-Our observations from the pilot search showed that using keywords such as “by GitHub Copilot”, “use GitHub Copilot”, and “with GitHub Copilot” can improve the accuracy of search results.
-These keywords enable us to focus more on the code generated using Copilot rather than code snippets that contain other content related to Copilot.
-In addition, since Python and JavaScript are the two most popular programming languages used by developers, and also the most frequently used with GitHub Copilot (see Section 3.1), we further limited the types of code snippets during the search to Python and JavaScript.
-We collected the search results under the Code label.
-Considering that some projects declare using GitHub Copilot generated code in their README files or project description provided in GitHub, we decided to retain the results from the Repository label in the search results.
-Figure 2 shows an example of our search process.
-Table 1 reports the search terms we used and the number of search results obtained from GitHub.
-In this step, we collected a total of 8,157 results, of which 7,749 were from the Code label, and 408 were from the Repository label.
-The same search result may contain multiple keywords, which means that there are duplicate projects in the collected data.
-After removing duplicate projects, we obtained a total of 2023 search results, of which 1847 were from the Code label, and 176 were from the Repository label.
-Figure 2. Example of the search process
-Table 1. Search results based on different terms used
-# Search Term # Code # Repositories
-1 “By GitHub Copilot” 2549 94
-2 “Use GitHub Copilot” 1822 123
-3 “With GitHub Copilot” 3378 191
-Total 7749 408
+By manually analyzing the search results, we found that restricting the search terms provided more accurate results.
+We decided to use keyword combinations such as {by, use, with} + {GitHub Copilot, CodeWhisperer, Codeium} as search keywords.
+This approach enabled us to focus more on the code generated using AI code generation tools (e.g., Copilot) rather than code snippets that contain other content related to the generation tools.
+In addition, we further limited the types of code snippets during the search to Python and JavaScript.
+Both are dynamically typed languages where a full analysis can be conducted without the need for compilation.
+On the other hand, statically typed languages such as Java and C++ require the code to be compiled first to generate the database required for analysis.
+Our data collection method makes it difficult to properly compile individual code snippets, given that some code is unavailable when conducting the analysis.
+We selected Python and JavaScript as they are currently the two most popular programming languages used by developers (GitHub, 2024b), and the most frequently used with GitHub Copilot (Zhang et al., 2023).
+The popularity of the two languages ensures that we can collect a substantial amount of generated code from GitHub projects, providing a significantly larger sample size and statistically meaningful results compared to other languages.
+We collected the search results under the filter Code with file names, file paths, and forks and stars of the repositories where the files were located.
+Considering that some projects declared using GitHub Copilot generated code in their README files or project descriptions provided in GitHub, we decided to retain the results from the filter Repo with project name, project path, project description, and forks and stars of the projects.
+Table 1 shows a breakdown of the search terms we used and the number of search results per language.
+The same search result may contain multiple keywords, meaning there are duplicate projects in the collected data.
+After removing duplications, we ended up with 3,589 search results, of which 3,141 were from the Code label and 445 from the Repository label.
+Table 1. Search results based on different terms
+Search Term Code (Py) Code (Js) Repo (Py) Repo (Js) Total
+by GitHub Copilot 1320 1256 48 43 2667
+use GitHub Copilot 1344 1384 84 93 2905
+with GitHub Copilot 1532 1876 95 177 3680
+by CodeWhisperer 39 44 0 1 84
+use CodeWhisperer 41 46 4 2 93
+with CodeWhisperer 68 39 4 0 111
+by Codeium 85 73 0 1 159
+use Codeium 31 26 0 1 58
+with Codeium 99 69 0 2 170
+Total 4559 4813 235 320 9927
 3.2.2. Filtering Code Snippets
 Step 2.
-After obtaining the data from the keyword searches, we further filtered them by not only considering the accuracy of the keywords but also by investigating the project’s documentation, code comments, and other metadata in the search results to determine whether they were generated by GitHub Copilot.
-Additionally, since we aimed to obtain code snippets used in real-world projects, we excluded search results used to solve simple programming practice problems on platforms, such as bisection lookup and fast sorting tasks from LeetCode.
+After obtaining the data from the keyword searches, we further filtered them by not only considering the accuracy of the keywords but also by investigating the project’s documentation, code comments, and other metadata in the search results to determine whether they were generated by GitHub Copilot and other code generation tools.
+Additionally, since we aimed to obtain code snippets used in open-source GitHub projects, we manually excluded search results related to solving simple programming practice problems on platforms that provide foundational exercises for enhancing algorithmic thinking and coding skills, such as bisection lookup and fast sorting tasks from LeetCode.
 We consider these programming exercises to be more closely related to correctness than security.
-We first conducted a pilot data filtration to better filter the search results.
 We begin by explaining the terminology used in data filtering: the search results under the Repository label are the projects that contain code files, and the search results under the Code label are individual code files.
@@
 Exclusion Criterion 1: We excluded code files used to solve simple programming practice problems.
+Simple programming practice problems mainly refer to those standard coding exercises on platforms such as LeetCode, which are often named BOJ-number.
 We provide examples for the three criteria in Figure 4, Figure 5, and Figure 6.
@@
 In Figure 6, the code snippet was removed as it was determined that the code just solved a simple algorithmic problem.
-Specifically, the pilot data filtering process consisted of the following steps: (1) The first author randomly selected 40 projects and 400 code files from the search results from the Repository label and the Code label, respectively.
-(2) Two authors independently labelled whether Copilot generated these projects and code files.
-(3) The first author compared the labelling results by the two coders and calculated the level of agreement between them using the Cohen’s Kappa coefficient (Cohen, 1960).
-(4) If there were any results they were unsure of or disagreed with, the two authors discussed until they reached an agreement.
-The Cohen’s Kappa coefficients were 0.79 for the projects (from the Repository label) and 0.85 for the code files (from the Code label), which were both higher than 0.7, indicating a high level of agreement between the two coders and ensuring a good accuracy of the labelling results.
-After completing the pilot data labelling, the first author checked the rest of the search results (1,447 code files and 136 projects).
-After removing the duplicate results, we retained 79 projects under the Repository label and 180 code files under the Code label.
-Step 3.
-To further get the application domains represented in the dataset, we categorized the projects based on the project’s description documents and the specific function of the code.
-Projects were classified into one of the following categories: Games, Web Applications, Utility Tools, AI Applications, Network Communication Applications, and Others.
-Figure 6 shows the classification and distribution of the application domains of the Copilot-generated code in the dataset from the Repository and Code labels respectively.
-For the code files retained under the Repository label, we consider the entire code file as code generated by Copilot.
-In other words, we assume that Copilot generates all the code in the file because it was stated in the README file that it was all generated by Copilot.
-For code files retained under the Code label, we know that the files contain code snippets, perhaps even just a few lines of code, generated by Copilot.
-Instead of identifying the specific Copilot-generated code in this step, we combine the warning messages from the security scan and the code comments in the file to determine whether Copilot generates the code snippet with the security problem (this process is explained further in Section 3.3.2).
-As a result, we obtained 452 code files with different contents, 272 from the Repository label and 180 from the Code label.
-Table 2 gives the type and number of code files.
-A curated dataset, comprising all the data collected during the research process, has been made available (Fu et al., 2024).
 Figure 4. Example of Inclusion Criterion 1: projects fully written by Copilot
@@
 Figure 6. Example of Exclusion Criterion 1: files used to solve simple algorithm problems
-(a) Code from the Repository label
-(b) Code from the Code label
-Figure 6. Application domains of the Copilot-generated code in our dataset
+We manually filtered all the data, labeling whether they were generated by Copilot and other code generation tools based on the three criteria mentioned above.
+Two authors independently completed this manual filtration process for all the data within a period of two weeks.
+The first author compared the labeling results by the two coders and calculated the level of agreement between them using Cohen’s Kappa coefficient (Cohen, 1960).
+Cohen’s Kappa coefficient was 0.84, which was higher than 0.8, indicating a high level of agreement between the two coders and ensuring a good accuracy of the labeling results.
+Besides, for those results where the two authors disagreed on the labeling, they resolved the disagreements through a negotiated agreement approach (Campbell et al., 2013) during a meeting.
+After manual filtration, we retained 116 projects under the Repository label and 335 code files under the Code label.
+For the code files retained under the Repository label, we consider the entire code file as code generated by Copilot and other code generation tools.
+In other words, we assume that Copilot and other code generation tools generate all the code in the file because it was stated in the README file that it was all generated by Copilot and other code generation tools.
+For code files retained under the Code label, we know that the files contain code snippets, perhaps even just a few lines of code, generated by Copilot and other code generation tools.
+Instead of identifying the specific AI-generated code in this step, we combine the warning messages from the security analysis and the code comments in the file to determine whether Copilot and other code generation tools generate the code snippet with the security problem (this process is explained further in Section 3.3.2).
+As a result, our dataset consisted of 733 distinct code files, with 672 from Copilot, 38 from CodeWhisperer, and 23 from Codeium.
+Table 2 gives the type and number of code files.
+A curated dataset, comprising all the data collected during the research process, has been made available (Fu et al., 2024).
 Table 2. Code snippets from GitHub
 # Language # Code Snippets: Repository # Code Snippets: Code Total
-L1 Python 143 134 277
-L2 JavaScript 129 46 175
-Total 272 180 452
-3.3. Data Processing and Analysis
-3.3.1. Data Processing
+L1 Python 190 229 419
+L2 JavaScript 208 106 314
+Total 398 335 733
+Step 3.
+To understand the features of our dataset, we first conducted a statistical analysis of the stars and forks of the projects where AI-generated code was present.
+The distribution of stars and forks are shown in Figure 7 and Figure 8, respectively.
+We found that the projects containing AI-generated code often tend to be small and have low popularity in the GitHub community, possibly often developed by individuals or small teams.
+This finding is consistent with the recent study by Yu et al.
+about the characteristics of LLM-generated code on GitHub (Yu et al., 2024).
+We speculated that the low popularity may be correlated with the age of the project.
+Most of these projects are relatively new as Copilot itself is also a recent development.
+Despite the low popularity and user engagement of most projects, there are a few exceptions in the dataset, such as the project OnmyojiAutoScript, which received 1,672 stars and 505 forks.
+In addition, we also counted the lines of code (LoC) of the collected code files.
+Figure 9 shows the distribution of LoC in different languages (Python and JavaScript respectively) and then the LoC of all code files (Python + JavaScript) in the dataset.
+We found that the median LoC of code is 74, while the average reaches 181.
+Although the files obtained from GitHub are relatively small in size, they are still significantly larger than the size of code generated using designed scenario prompts.
+For example, when Majdinasab et al.
+(Majdinasab et al., 2024) replicated the experiment by Pearce et al.
+(Pearce et al., 2022), the median LoC of code generated by their designed prompts through Copilot was 42, with an average of 40.
+While using the SecurityEval dataset (Siddiq and Santos, 2022) as prompts to test Copilot, the LoC of generated code is concentrated between 5 and 26.
+This indicates that our collected dataset has an advantage in terms of the size of the generated code.
+Figure 7. Distribution of stars of the projects where AI-generated code was present
+Figure 8. Distribution of forks of the projects where AI-generated code was present
+Figure 9. Distribution of LoC of the files where AI-generated code was present
+We also analyzed the application domains of the code snippets in our dataset.
+To understand the application domains, we classified the code in the dataset based on the project description and the specific functions of the code files.
+Two authors independently classified all code snippets within a period of ten days.
+If their categorization results diverged, the two authors discussed until they reached a consensus.
+In total, we categorized the generated code into six application domain categories: Game, Web Application, Utility Tool, AI Application, Network Communication, and Others.
+Table 3 shows the descriptions of application domains and the categorization of code snippets in different programming languages.
+Table 3. Application domains of the files where AI-generated code was present
+Category Description Python JavaScript
+Utility Tool Involves scripts, command-line tools, and other programs that simplify daily tasks. 39.0% 30.1%
+Web Application Code related to front-end, back-end development, and server-side operations. 15.6% 49.8%
+AI Application Code for machine learning, deep learning, natural language processing, or other intelligent applications. 15.6% 2.8%
+Game Contains code related to game development, using engines such as Unity, Unreal, and Godot. 15.9% 8.5%
+Network Communication Involves code for implementing network protocols, client/server communication. 8.2% 5.7%
+Others Code that does not fit into the above categories. 5.4% 2.8%
+3.3. Data Analysis and Results Filtering
+3.3.1. Data Analysis
 Step 4.
-CodeQL is a scalable static security analysis tool that is widely used in practice and allows users to analyze code and detect relevant weaknesses using predefined queries and test suites and supports for multiple languages (including Java, JavaScript, C++, C#, Go, and Python (GitHub, 2023a)).
+We used well-known automated static analysis tools suggested by OWASP (OWASP, 2024) to scan the collected code snippets.
+The reason why we chose a static rather than dynamic approach is that the static analysis has higher coverage and is able to analyze programs without executing them (Dunlap et al., 2023).
+While dynamic analysis offers more precise insights by reasoning about program behavior, it suffers from limited coverage and can be expensive (Sui et al., 2020; Ernst, 2003; Tahir and MacDonell, 2012).
+Static security analysis tools have been widely used by developers and companies (Distefano et al., 2019; Sadowski et al., 2018).
+Employing static analysis enables us to analyze program segments (smaller snippets) without requiring the entire program to be executed.
+Since different static analysis tools may use different algorithms and rules to detect security weaknesses, using multiple tools can increase our chances of discovering security weaknesses in the code.
+To improve the coverage and accuracy of the results, we used two static analysis tools for security checks on each code snippet (i.e., CodeQL plus one dedicated tool for the specific language, Bandit for Python, and ESLint for JavaScript).
+CodeQL is a scalable static security analysis tool that is widely used in practice and allows users to analyze code and detect relevant weaknesses using predefined queries and test suites and supports for multiple languages (including Java, JavaScript, C++, C#, Go, and Python (GitHub, 2023)).
 Before using CodeQL to scan the identified code snippets for security weaknesses, we needed to create a CodeQL database for the source code.
 Source code can be directly analyzed for interpreted languages like Python and JavaScript without being compiled into intermediate code.
-To speed up the scan, we stored 20 files in each database using the command line to generate the database needed for CodeqQL queries.
-It should be noted that generating a database for an exceptionally large number of files would increase the database compilation and scanning time, which is much longer than partitioning them into small databases.
-In total, we obtained 25 code databases available for CodeQL scanning.
-3.3.2. Data Analysis
-Step 5.
-We used well-known automated static analysis tools listed by OWASP (OWASP, [n.
-d.]) to scan the collected code snippets.
-Static analysis has been widely used to detect security issues in code, for its ability to analyze programs without execution (Dunlap et al., 2023).
-While dynamic analysis offers more precise insights by reasoning about program behavior, it suffers from limited coverage (Sui et al., 2020).
-Employing static analysis enables us to analyze program segments (smaller snippets) without requiring the entire program to be executed.
-Since different static analysis tools may use different algorithms and rules to detect security weaknesses, using multiple tools can increase our chances of discovering security issues in the code.
-To improve the coverage and accuracy of the results, we used two static analysis tools for security checks on each code snippet (i.e., CodeQL plus one dedicated tool for the specific language, Bandit for Python and ESLint for JavaScript).
-We first used CodeQL to analyze the code in our dataset.
-The default query suite for the standard CodeQL query package is codeql-suites/<lang>-code-scanning.qls.
-Each package has several query suites in the codeql-suite directory.
+Thus, we directly use the command line to generate the database needed for CodeQL queries.
+Then, we used CodeQL to analyze the generated database.
+We can find the available query suites for different programming languages in the codeql repository.
 For example, the codeql/python-queries package contains the following query suites (GitHub, 2023):
@@
 For example, the python-security-and-quality.qls test suite for Python provides 168 security checks, and the JavaScript test suite provides 203 security checks.
-As the query reports only provide the name and description of the security issues, we manually matched the results in the query reports with the corresponding CWE IDs.
-We then selected one additional static security analysis tool for files in each programming language we analyzed: Bandit for Python and ESLint for JavaScript.
-As explained in Section 3.2, we considered the code snippet from the Repository label to be the entire code file, while the code snippet from the Code label exists in the code file, with the exact number of lines unspecified at this stage.
-Step 6.
-We scanned code snippets from the Repository and Code labels, and we filtered the scan results before analyzing them.
+As the query reports only provide the name and description of the security weaknesses, we manually matched the results in the query reports with the corresponding CWE IDs based on the official documentation provided (CodeQL, 2024).
+For Python files, we used Bandit (PyCQA, 2024) – a Python’s static security analysis tool.
+When running Bandit, we enabled all security rules by default.
+Bandit had a total of 73 security check rules, and each Bandit rule was typically associated with a corresponding CWE number, helping developers better understand the categories of detected vulnerabilities and align with industry standards.
+For JavaScript files, we used ESLint (ESLint Contributors, 2024) – a widely used JavaScript static analysis tool that can detect common coding issues and security vulnerabilities.
+Since we needed to configure the security check rules ourselves, we used the eslint-plugin-security plugin for security checks, which is an ESLint plugin containing 14 security check rules designed to detect common security vulnerabilities in JavaScript.
+In addition, we also configured the plugins eslint-plugin-no-secrets and eslint-plugin-xss with more precise rules to identify security weaknesses more accurately.
+For example, some of the rules in eslint-plugin-security also deal with XSS protection (such as avoiding unsafe eval() usage), but eslint-plugin-xss provides more precise rules against front-end XSS attacks.
+Both tools accept raw code snippets as input, and there is no need to pre-process the data before using them.
+As explained in Section 3.2, we considered the code snippet from the Repository label to be the entire code file, while the code snippet from the Code label exists in the code file, with the exact location of the code snippets unspecified at this stage.
+3.3.2. Results Filtering
+Step 5.
+We scanned code snippets from the Repository and Code labels, and we filtered the analysis results before analyzing them.
 In this step, we adopted part of the strategies used by developers when they perform different tasks with static analysis tools: warning prioritization and determining whether a warning is a false positive or a true report (Do et al., 2020).
 We first performed an initial filtering of the results based on the priority of the warnings.
-Specifically we removed the repeated scan results that were reported by the two tools, then removed the scan results that were not security issues, such as recommendations, to get the initial scan results.
-Subsequently, to confirm that the security issues were actually caused by the Copilot-generated code, the first author manually checked the scan results.
-For the initial results obtained under the Repository label, we checked them one by one.
-Specifically, we determined whether the corresponding location of the code snippet indeed had a security issue based on the line number information provided by the scan results.
-For the initial results obtained under the Code label, we confirmed whether the security issues truly existed and verified if the security issues were caused by Copilot-generated code.
-Specifically, after scanning the code file, we pinpointed the code snippet within the file based on the line number of the security issue indicated in the scan results.
-We assessed whether it was generated by Copilot by checking the surrounding comments and determined whether a security issue existed in that particular context.
-We further analyzed the filtered scan results in Step 7, detailed in Section 4, according to the specific RQs.
-We provide the complete dataset (including code snippets, full scan results, and filtered results) in our replication package (Fu et al., 2024).
-4. Results
-We present the results of three RQs formulated in Section 3.1 below. For each RQ, we first explain how we analyzed the collected code snippets to answer the RQ. We then answer each of our three RQs.
-4.1. RQ1: How secure is the code generated by Copilot?
-Approach.
-To answer this RQ, we collected 452 Python and JavaScript code snippets generated by Copilot from GitHub projects.
-We used two static analysis tools (CodeQL + another language-dedicated tool, Bandit for Python and ESLint for JavaScript) to scan and analyze the code snippets and then combine the results obtained from the two tools.
-The aim is to achieve better coverage of security issues.
-Therefore, as long as one of the tools detected the presence of a security issue, the code snippet was considered vulnerable.
+Specifically, we manually removed the repeated analysis results that were reported by the two tools (14 duplicate results by Bandit and 2 duplicate results by ESLint), keeping only one for further analysis.
+The first author eliminated analysis results that were not security weaknesses.
 We identified three types of warnings from CodeQL analysis:
@@
 Since our research primarily focused on security weaknesses, we only counted code snippets that had warnings and errors, and we ignored recommendations on code quality.
-After obtaining the initial scan results related to security issues, we manually checked the scan results.
-For the scan results from the Repository label, we marked them as 1 (security issue exists) and 0 (no security issue exists).
-For the scan results from Code label, we marked them as 1 (security issue exists) and 0 (no security issue exists or the code snippet with a security issue was not generated by Copilot).
-Figure 7 shows the outcomes of the manual verification on the initial security scan results under the Repository and Code labels.
-When manually checking the scan results from the Code label, we also needed to identify whether the security issues obtained from the scan were from Copilot-generated code snippets based on the comment (prompt message) that appears before the method.
-We provide a working example of the filtration of the scan results in Figure 8.
-In Step 1, we first went to the corresponding file to locate the specific code snippet based on the start line numbers of the scan results.
+After obtaining preliminary analysis results related to security weaknesses, we manually checked the analysis results to confirm that the security weaknesses were actually caused by the automatically generated code.
+For the initial results obtained under the Repository label, we checked them one by one.
+Specifically, we determined whether the corresponding location of the code snippet indeed had a security weakness based on the line number information provided by the analysis results.
+For the initial results obtained under the Code label, we confirmed whether the security weaknesses truly existed and verified if the security weaknesses were caused by Copilot-generated code.
+Specifically, after scanning the code file, we pinpointed the code snippet within the file based on the line number of the security weakness indicated in the analysis results.
+We assessed whether it was generated by Copilot or other generation tools by checking the surrounding comments and determined whether a security weakness existed in that particular context.
+We further analyzed the filtered analysis results in Step 6, detailed in Section 4, according to the specific RQs.
+3.4. Fixing Code with Security Weaknesses
+Step 7.
+Based on RQ2 results, we randomly selected a subset of code snippets containing security weaknesses as the repair dataset since our purpose is to explore the feasibility of Copilot Chat in fixing the security weaknesses.
+These code snippets were all generated by Copilot, as most of the code obtained from GitHub was produced by Copilot.
+Since we aimed to fix as many security weaknesses as possible, we did not use clean data (i.e., code snippets without known security weaknesses).
+We randomly selected 50 Python and 40 JavaScript code snippets from the Repo label, which contained 163 and 132 security weaknesses, respectively.
+We utilized Copilot Chat to remediate the security weaknesses identified by the static analysis tools.
+We conducted several experiments on the selected dataset with 295 security weaknesses using three different prompts.
+Copilot Chat provides developers with basic slash commands, avoiding the need to write complex prompts for common scenarios.
+The slash commands include: /tests: generates unit tests for the selected code.
+/fix: provides fix suggestions for issues in the selected code.
+/explain: explains the selected code.
+/optimize: analyzes and improves the runtime of the selected code.
+We first used the slash command /fix to provide fix suggestions for the selected code.
+Then, we performed the fixes using our designed basic prompts and enhanced prompts with additional context.
+The prompt designs are as follows:
+Basic Prompt Is there a security weakness with this code snippet? If so, please fix it and return the fixed code.
+Enhanced Prompt This code snippet is known to have some security weaknesses.
+The warning messages are as follows: <insert warning messages here>.
+Please fix these security weaknesses and return the fixed code.
+Figure 10. An example of using Copilot Chat to fix security weaknesses in Visual Studio Code
+Figure 10 shows an example of how to use Copilot Chat to fix the security weaknesses in the code.
+We first select code snippets with security weaknesses and provide the three prompts (/fix, basic prompt, and enhanced prompt) in the IDE chat window.
+Then, Copilot Chat uses the selected code as a reference (input) to generate responses to our prompts, which we can accept or reject.
+Considering the token limitation issue, we always selected the function block containing the security weaknesses as the reference for Copilot Chat.
+After accepting the fix suggestions from Copilot Chat, we used the static analysis tools to recheck the modified code snippets.
+We recorded the positions of the modified code, indicating where the accepted code is located.
+Then we examined whether security weaknesses still exist in the corresponding positions of the modified code according to the rechecking results of the static analysis tools.
+We provided the complete dataset (including code snippets, full analysis results, and filtered results) in our replication package (Fu et al., 2024).
+4. Results
+In this section, We present the results of the three RQs formulated in Section 3.1.
+For each RQ, we first explain how we analyzed the collected code snippets to answer the RQ, and we then answer each of our three RQs.
+4.1. RQ1: How secure is the code generated by Copilot in GitHub projects?
+Approach.
+To answer this RQ, we collected 733 Python and JavaScript code snippets generated by Copilot and two other tools from GitHub projects, with 672 from Copilot, 38 from CodeWhisperer, and 23 from Codeium.
+We used two static analysis tools (CodeQL + another language-dedicated tool, Bandit for Python, and ESLint for JavaScript) to scan.
+The aim is to achieve better coverage of security weaknesses.
+Therefore, as long as one of the tools detected the presence of a security weakness, the code snippet was considered vulnerable.
+Subsequently, we filtered the analysis results as described in Section 3.3.2.
+For the analysis results from the Repository label, we marked them as 1 (security weakness exists) and 0 (no security weakness exists).
+For the analysis results from Code label, we marked them as 1 (security weakness exists) and 0 (no security weakness exists or the code snippet with a security weakness was not generated by code generation tools).
+Figure 11. Example of filtering analysis results from the Code label that are generated by Copilot
+When manually checking the analysis results from the Code label, we also needed to identify whether the security weaknesses obtained from the scan were from automatically generated code snippets based on the comment (prompt message) that appears before the method.
+We provide a working example of the filtration of the analysis results in Figure 11.
+In Step 1, we first went to the corresponding file to locate the specific code snippet based on the start line numbers of the analysis results.
 In Step 2, we located the code at Line 149 in aggregation.py.
-We found that this code does have a security issue and determined that it was generated by Copilot based on the prompt messages above.
-Consequently, we marked the corresponding security scan result as “1”.
+We found that this code does have a security weakness and determined that it was generated by Copilot based on the prompt messages above.
+Consequently, we marked the corresponding security analysis result as “1”.
 We also located the code at Line 145 in _lytools.py.
-We found that this code had a security issue, but there were no prompt messages that would indicate that it was generated by Copilot.
-Consequently, we marked the corresponding security scan result as “0” and discarded this scan result from further analysis.
-Finally, we kept the results marked as 1 and aggregated the filtered results obtained using multiple analysis tools to calculate the number of code snippets with security issues detected.
-(a) Results from the Repository label
-(b) Results from the Code label
-Figure 7. Manual verification of the initial security scan results for the Repository and Code label
-Figure 8. Example of filtering scan results from the Code label that are generated by Copilot
+We found that this code had a security weakness, but there were no prompt messages that would indicate that it was generated by Copilot.
+Consequently, we marked the corresponding security analysis result as “0” and discarded this analysis result from further analysis.
+Two authors independently filtered the analysis results within a period of two weeks, and if there were any results that they were unsure of or disagreed with, the two authors discussed the results until they reached an agreement.
+Cohen’s Kappa coefficients (Cohen, 1960) were 0.85 for the analysis results from the Repository label, and 0.82 for the analysis results from the Code label, which were both higher than 0.8, ensuring a good accuracy of the filtering results.
+Finally, we kept the results marked as 1 and aggregated the filtered results obtained using multiple analysis tools to calculate the number of code snippets with security weaknesses detected.
+All labeling results can be found in our online replication package (Fu et al., 2024).
 Results.
-Table 3 shows the numbers of code snippets for different types and the numbers and percentages of code snippets with security weaknesses.
-From the statistical results, we found that out of the 452 code snippets generated by Copilot, 25.9% of them have security weaknesses.
-Among the Python and JavaScript code snippets, Python exhibits a higher proportion of security issues compared to JavaScript.
-Out of the 277 Python code snippets, 91 (32.8%) have security weaknesses.
-Among the 175 JavaScript code snippets we collected, 43 (24.6%) have security weaknesses.
-Note that one related code snippet may contain multiple instances of a specific CWE.
-Table 3. The number and percentage of code snippets with security weaknesses generated by Copilot
-Language # Snippets
+Table 4 shows the number of code snippets containing security weaknesses and the number of security weaknesses identified in the code snippets.
+From the statistical results, we found that out of the 733 generated code snippets, 27.3% of them contained security weaknesses.
+Around 30% of Python code snippets (124 of the 419) and around 25% of JavaScript code snippets (76 of the 314) contained security weaknesses.
+At the same time, we can see that more than half of the code snippets containing security weaknesses (102 of 200) have more than one security issue.
+In Figure 12, we present the distribution of application domains for the 200 code snippets (27.3% of all) identified with security weaknesses.
+We also found 628 security weaknesses in the 200 code snippets.
+Note that one code snippet may contain multiple security weaknesses attributed to multiple CWEs.
+Figure 13 shows the distribution of security weaknesses per code snippet.
+Our analysis reveals an average incidence of 3 security weaknesses per code snippet, irrespective of the programming language.
+The maximum count of security weaknesses within a single code snippet, which consisted of 1,595 LoC, reached 22.
+Table 4. Statistics of the code snippets with security weaknesses
+Language # Snippets # Security weaknesses
 # Snippets containing
 security weaknesses
-Python 277 91 32.8%
-JavaScript 175 43 24.6%
-Total 452 134 29.6%
-4.2. RQ2: What security weaknesses are present in the code snippets generated by Copilot?
+# containing more than
+one security weakness
+Python 419 387 124 (29.5%) 65 (15.5%)
+JavaScript 314 241 76 (24.2%) 37 (11.8%)
+Total 733 628 200 (27.3%) 102 (13.9%)
+(a) application domains of Python code
+(b) application domains of JavaScript code
+Figure 12. Application domains of the code snippets with security weaknesses
+Figure 13. Distribution of the number of security weaknesses per code snippet
+4.2. RQ2: What security weaknesses are present in the code generated by Copilot across different application domains?
 Approach.
-To answer RQ2, we processed the scan results collected by RQ1 that were manually checked to contain vulnerabilities.
-In total, we found 544 security issues in 452 code snippets, and Table 4 shows the number of security weaknesses found in code files of different programming languages.
-Table 4. The number of security weaknesses in code snippets generated by Copilot
-Language
-# Snippets containing
-security weaknesses
-# Total security
-weaknesses
-Python 91 277
-JavaScript 43 175
-Total 134 544
-For each code snippet, we used CWEs to classify the security issues identified by the static analysis tools.
-Each CWE has a unique ID and a set of related descriptions, including its potential impact and how to detect and fix the CWE (The MITRE Corporation, 2023b).
-To more accurately identify the type of security issue with the corresponding CWE, we did not directly accept the provided information of the static analysis tools (such as the CWE number reported by Bandit).
-Instead, we manually identified the type of security issue at the specific location reported in the warning message of the static analysis tool and matched it with the corresponding CWE number.
-Initially, two authors independently matched each description of the security issue with a CWE ID.
+To answer RQ2, we processed the analysis results collected by RQ1, which were manually checked to see if they contained security weaknesses.
+For each code snippet, we used CWEs to classify the security weaknesses identified by the static analysis tools.
+Each CWE has a unique ID and a set of related descriptions, including its potential impact and how to detect and fix the CWE (MITRE, 2024).
+To accurately identify the type of security weakness with the corresponding CWE, we did not directly accept the information provided by the static analysis tools (for example, only Bandit provided the CWE IDs of the security issues identified, and we did not follow the CWE IDs of 55 out of 235 security issues reported by Bandit).
+Instead, we manually identified the type of security weakness at the specific location reported in the warning message of the static analysis tool.
+We utilized the warning information provided by static analysis tools, such as the test_name and message fields indicated by CodeQL, in conjunction with relevant descriptions from the CWE, to perform manual correspondences.
+Initially, two authors independently matched each description of the security issue with a CWE ID within a period of ten days.
+Cohen’s Kappa coefficient (Cohen, 1960) was 0.82, which was higher than 0.8 and indicated a high level of agreement between the two authors.
 In case of disagreement, a discussion was initiated between the two authors, and one other author (a security expert) was then involved to provide his assessment.
-This process continued until all the descriptions of the security issues in the results were matched with CWE IDs.
-In the final stage, we performed a statistical analysis of CWE weaknesses in 134 code snippets that contained security weaknesses.
+We provided correspondences between the warning prompt messages and CWEs in our online replication package (Fu et al., 2024).
+In the final stage, we performed a statistical analysis of CWE weaknesses in 200 code snippets that contained security weaknesses.
+Additionally, we analyzed the distribution of different CWEs across different application domains of the code snippets.
+We also discussed whether the security weaknesses present in the code generated by Copilot are widely prevalent in the code produced during software development.
+The code snippets in our collected dataset were mainly generated in 2022 and 2023.
+To compare whether the security weaknesses in in our dataset are widespread in this period, we chose MITRE 2023 CWE Top-25 list (MITRE, 2023) as our baseline.
+We compared the CWEs obtained in RQ2 with the 2023 CWE Top 25.
+Table 5. Distribution of CWEs in code snippets
+CWE-ID Name Frequency Percentage
+CWE-330 Use of Insufficiently Random Values Weakness 114 18.15%
+CWE-94 Improper Control of Generation of Code (’Code Injection’) 62 9.87%
+CWE-79 Improper Neutralization of Input During Web Page Generation (’Cross-site Scripting’) 60 9.55%
+CWE-78 Improper Neutralization of Special Elements used in an OS Command (’OS Command Injection’) 39 6.21%
+CWE-427 Uncontrolled Search Path Element 35 5.57%
+CWE-457 Use of Uninitialized Variable 30 4.78%
+CWE-22 Improper Limitation of a Pathname to a Restricted Directory (’Path Traversal’) 29 4.62%
+CWE-772 Missing Release of Resource after Effective Lifetime 29 4.62%
+CWE-89 Improper Neutralization of Special Elements used in an SQL Command (’SQL Injection’) 27 4.30%
+CWE-259 Use of Hard-coded Password 16 2.55%
+CWE-685 Function Call with Incorrect Number of Arguments 14 2.23%
+CWE-284 Improper Access Control 13 2.07%
+CWE-312 Cleartext Storage of Sensitive Information 12 1.91%
+CWE-390 Detection of Error Condition Without Action 11 1.75%
+CWE-456 Missing Initialization of Variable 11 1.75%
+CWE-607 Use of Wrong Operator in String Comparison 11 1.75%
+CWE-770 Allocation of Resources Without Limits or Throttling 11 1.75%
+CWE-20 Improper Input Validation 8 1.27%
+CWE-665 Improper Initialization 8 1.27%
+CWE-117 Improper Output Neutralization for Logs 7 1.11%
+CWE-400 Uncontrolled Resource Consumption 7 1.11%
+CWE-561 Dead Code 7 1.11%
+CWE-732 Incorrect Permission Assignment for Critical Resource 6 0.96%
+CWE-798 Use of Hard-coded Credentials 6 0.96%
+CWE-215 Information Exposure Through Debug Information 5 0.80%
+CWE-290 Authentication Bypass by Spoofing 5 0.80%
+CWE-295 Improper Certificate Validation 5 0.80%
+CWE-209 Information Exposure Through an Error Message 4 0.64%
+CWE-252 Unchecked Return Value 4 0.64%
+CWE-571 Expression is Always True 4 0.64%
+CWE-605 Multiple Binds to the Same Port 4 0.64%
+CWE-200 Information Exposure 3 0.48%
+CWE-327 Use of a Broken or Risky Cryptographic Algorithm 3 0.48%
+CWE-367 Time-of-check Time-of-use Race Condition 3 0.48%
+CWE-563 Assignment of a Fixed Address to a Pointer 3 0.48%
+CWE-116 Improper Encoding or Escaping of Output 2 0.32%
+CWE-480 Use of Incorrect Operator 2 0.32%
+CWE-502 Deserialization of Untrusted Data 2 0.32%
+CWE-601 URL Redirection to Untrusted Site (’Open Redirect’) 2 0.32%
+CWE-208 Observable Timing Discrepancy 1 0.16%
+CWE-570 Expression is Always False 1 0.16%
+CWE-628 Function Not Implemented Correctly 1 0.16%
+CWE-682 Incorrect Calculation 1 0.16%
+43 Types Total: 628
 Results.
-Table LABEL:Distribution_of_CWEs_in_code_snippets shows the distribution of CWEs in the code snippets and the total number of occurrences (Frequency) of the CWE in the code snippets (we put those CWEs whose Frequency = 1 in “Others”).
-In total, we found 544 CWEs in 452 code snippets.
-These security weaknesses were related to 38 CWEs, indicating that developers face a variety of security weaknesses when using Copilot.
-CWE-330: Use of Insufficiently Random Values is the most frequently occurring CWE, as it represents 23.3% of the security weaknesses, followed by CWE-94: Improper Control of Generation of Code (‘Code Injection’) , CWE-78: OS Command Injection and CWE-95: Improper Neutralization of Directives in Dynamically Evaluated Code (‘Eval Injection’).
-Some CWEs appear less frequently, for example CWE-117: Improper Output Neutralization for Logs only occurred twice.
-Furthermore, many CWEs occur with a frequency of less than 1%, for example, CWE-22: Improper Limitation of a Pathname to a Restricted Directoryand CWE-770: Allocation of Resources Without Limits or Throttling.
-This indicates that the types of security issue are closely related to the specific scenarios in which developers use Copilot, emphasizing the importance of maintaining vigilance and caution when programming.
-Besides, Table 6 presents the top 5 CWEs that appear in Python and JavaScript code snippets.
-Table 5. Distribution of CWEs in code snippets
-CWE-ID Frequency of Specific CWE Percentage
-CWE-330 127 23.3%
-CWE-94 115 21.1%
-CWE-78 81 14.9%
-CWE-95 30 5.5%
-CWE-772 23 4.2%
-CWE-89 15 2.8%
-CWE-457 15 2.8%
-CWE-259 14 2.6%
-CWE-670 13 2.4%
-CWE-396 11 2.0%
-CWE-617 8 1.5%
-CWE-628 7 1.3%
-CWE-563 7 1.3%
-CWE-561 7 1.3%
-CWE-798 6 1.1%
-CWE-703 6 1.1%
-CWE-502 6 1.1%
-CWE-20 6 1.1%
-CWE-312 5 0.9%
-CWE-605 4 0.7%
-CWE-295 4 0.7%
-CWE-252 4 0.7%
-CWE-367 3 0.5%
-CWE-327 3 0.5%
-CWE-185 3 0.5%
-CWE-79 2 0.4%
-CWE-682 2 0.4%
-CWE-665 2 0.4%
-CWE-400 2 0.4%
-CWE-215 2 0.4%
-CWE-200 2 0.4%
-CWE-117 2 0.4%
-CWE-116 2 0.4%
-Others =1 ¡1%
-38 Types Total: 544
-Table 6. Top 5 CWEs in Python and JavaScript
+Table LABEL:Distribution_of_CWEs_in_code_snippets shows the distribution of CWEs in the code snippets and the total number of occurrences (frequency) of the CWEs in the code snippets.
+In total, we found 628 CWEs in 200 code snippets with security weaknesses.
+These security weaknesses were related to 43 CWE types, indicating that developers face a variety of security weaknesses when using Copilot and other code generation tools.
+CWE-330: Use of Insufficiently Random Values is the most frequently occurring CWE, as it represents 18.15% of the security weaknesses, followed by CWE-94: Code Injection , CWE-79: Cross-site Scripting, and CWE-78: OS Command Injection.
+Some CWEs appear less frequently.
+For example, CWE-117: Improper Output Neutralization for Logs only occurred twice.
+Furthermore, many CWEs occur with a frequency of less than 1%, for example, CWE-117: Improper Output Neutralization for Logsand CWE-290: Authentication Bypass by Spoofing.
+This indicates that the types of security weaknesses are closely related to the specific scenarios in which developers use code generation tools, emphasizing the importance of maintaining vigilance and caution when programming.
+In addition, we highlighted those CWEs that are the “Top-25 CWEs” in bold.
+It is worth noting that 233 security weaknesses in the code snippet correspond to these eight CWEs, which belong to the Top-25 CWE list, indicating that the CWE Top-25 weaknesses are also prevalent in the code generated by Copilot and other tools.
+At the same time, we can see that CWE-79: Cross-site Scripting and CWE-78: OS Command Injection are among the most frequently occurring security weaknesses in our results and rank high on the Top-25 CWE list.
+Table LABEL:tab:cwe_types presents the top 5 CWEs that appear in Python and JavaScript code snippets.
+We found that the detected CWEs vary across programming languages.
+CWEs in Python are mainly related to data processing and system calls, such as CWE-78: OS Command Injection and CWE-427: Uncontrolled Search Path Element.
+In contrast, CWEs in JavaScript are primarily associated with dynamic code generation problems and security issues in Web development, such as CWE-94: Code Injection and CWE-79: Cross-site Scripting.
+Furthermore, Figure LABEL:fig:cwe-py and Figure LABEL:fig:cwe-js show the distribution of CWEs across different application domains.
+We can observe that the number and type of instances of CWE vary from one application domain to another.
+In addition to this, the distribution of CWEs within the same application domain varies by programming language.
+For example, in Python, the application domain with the most CWEs is Utility Tool, while in JavaScript, Web Applications have the most CWEs.
+Meanwhile, we can see that in the Web Application category, CWE-79: Cross-site Scripting is the most frequent in JavaScript code snippets, while CWE-89: SQL Injection is the most present in Python code snippets.
 Rank CWE Type
@@
 1 CWE-330 (Use of Insufficiently Random Values Weakness) CWE-94 (Improper Control of Generation of Code)
-2 CWE-78 (Improper Neutralization of Special Elements used in an OS Command) CWE-95 (Improper Neutralization of Directives in Dynamically Evaluated Code)
-3 CWE-772 (Missing Release of Resource after Effective Lifetime) CWE-563 (Assignment to Variable without Use)
-4 CWE-89 (Improper Neutralization of Special Elements used in an SQL Command) CWE-20 (Improper Input Validation)
-5 CWE-259 (Use of Hard-coded Password) CWE-185 (Incorrect Regular Expression)
-4.3. RQ3: How many security weaknesses belong to the CWE Top-25?
+2 CWE-78 (Improper Neutralization of Special Elements used in an OS Command) CWE-79 (Improper Neutralization of Directives in Dynamically Evaluated Code)
+3 CWE-427(Uncontrolled Search Path Element) CWE-22 (Improper Limitation of a Pathname to a Restricted Directory (’Path Traversal’))
+4 CWE-457 (Use of Uninitialized Variable) CWE-685 (Function Call with Incorrect Number of Arguments)
+5 CWE-772 (Missing Release of Resource after Effective Lifetime) CWE-770 (Allocation of Resources Without Limits or Throttling)
+4.3. RQ3: Can Copilot Chat fix the security weaknesses of code generated by Copilot?
 Approach.
-The code snippets in our collected dataset were mainly generated in 2022 and 2023.
-To compare whether the security issues in Copilot-generated code are widespread in this period, we chosed MITRE 2023 CWE Top-25 list (The MITRE Corporation, 2023a) as our baseline.
-In addition, MITRE maintains another list of “Stubborn Weaknesses in the CWE Top 25” (The MITRE Corporation, 2023d), which includes those CWEs that have consistently appeared in the Top 25 most dangerous software weaknesses of the past five years (2019-2023).
-We compared the CWEs obtained in RQ2 with the 2023 CWE Top 25 and “Stubborn Weaknesses” in the CWE Top 25.
+To answer RQ3, we conducted experiments using Copilot Chat on the randomly selected code snippets containing security weaknesses and asked Copilot Chat to provide fix suggestions based on the three prompts (see Section 3.4).
+We then fixed the code according to the suggestion provided by Copilot Chat.
+Specifically, we used Copilot Chat to repair snippets with security weaknesses.
+The repaired code snippets are then checked for security weaknesses using the same static analysis tools.
+After getting the analysis results, The first author used the same filtering steps described in Section 3.3.2 to filter the results.
+We first removed the results that were repeatedly reported by two of the tools.
+Then, we removed the analysis results that were not security weaknesses.
+Among the three types of warnings from CodeQL analysis, we only counted code snippets that had warnings and errors, and ignored recommendations on code quality.
+Since the code snippets we selected all come from the Repo label, it is easier to filter the results related to security weaknesses, as we did not need to locate the specific line numbers of the code generated by Copilot in each file.
+Based on the results of the two successive security analysis, we determine whether the security weakness was successfully removed and the code is fixed.
 Results.
-The distribution of CWEs found compared to the MITER list is shown in Table 7.
-The results show that the CWE weaknesses present in the code generated by Copilot belong to eight CWEs.
-We highlight those CWEs that are “Stubborn Weaknesses” (SW) in bold.
-This means these are present issues and are currently among the most common and serious security weaknesses in practice.
-It is worth noting that the 218 security issues present in the code snippet correspond to these eight CWEs, while another 30 CWEs cover the remaining 326 security issues.
-This indicates that the CWE Top-25 weaknesses are also prevalent in the code generated by Copilot.
-At the same time, we can see that CWE-78: OS Command Injection occurs more frequently in Copilot-generated code.
-CWE-94: Improper Control of Generation of Code is in the list of the CWE Top-25 (The MITRE Corporation, 2023a), and meanwhile it is also one of the weaknesses with a high occurrence frequency in our results.
-Some CWEs with a higher ranking in the Top-25 list do not frequently appear in Copilot-generated code, such as CWE-79: Cross-site Scripting and CWE-89: SQL Injection.
-Table 7. The CWEs that belong to the 2023 CWE Top-25 list
-CWE-ID Description Is SW # Related Snippets Frequency
-CWE-94 Improper Control of Generation of Code No 34 115
-CWE-78 Improper Neutralization of Special Elements used in an OS Command Yes 18 81
-CWE-798 Use of Hard-coded Credentials Yes 3 6
-CWE-502 Deserialization of Untrusted Data Yes 2 6
-CWE-20 Improper Input Validation Yes 3 6
-CWE-79 Improper Neutralization of Input During Web Page Generation Yes 2 2
-CWE-276 Incorrect Default Permissions No 1 1
-CWE-22 Improper Limitation of a Pathname to a Restricted Directory Yes 1 1
-Total 218
+Table LABEL:table:fix shows the proportion of CWEs that were fixed under the three prompts.
+We can see that using the slash command fix can fix 19.3% of security weaknesses, while our designed basic and enhanced prompts can fix 31.8% and 55.5% of security weaknesses, respectively.
+Copilot Chat demonstrates promising capability in fixing security weaknesses regardless of the prompt used.
+At the same time, providing enhanced information in the prompts helps improve the proportion of fixes.
+Figure LABEL:fig:cwe-repair shows the number of security weaknesses fixed in code snippets of different languages when using various prompts.
+With basic prompts, security weaknesses were fixed in 29.4% (48 of 163) of Python code and 34.8% (46 of 132) in JavaScript.
+When using the enhanced prompt, the percentage of security weaknesses fixed in both languages rose to 58.3% (95 of 163) and 51.5% (68 of 132), respectively.
+We can see that Copilot Chat’s ability to fix security weaknesses in Python and JavaScript snippets does not differ significantly.
+Prompt Total CWE Before Fix Total CWE After Fix Fix Rate (%)
+/fix 295 238 19.3%
+basic 295 201 31.8%
+enhanced 295 132 55.5%
+Furthermore, Figure LABEL:fig:cwe-fix shows a heatmap with the number of different CWEs that were fixed when using different prompts.
+We can visually observe that Copilot Chat’s effectiveness in fixing different CWEs varies.
+For certain CWEs, Copilot Chat can perform well in repairing those snippets and provide a complete fix (i.e., remove the CWE from the code) such as CWE-259: Use of Hard-coded Password and CWE-685: Function Call with Incorrect Number of Arguments, which can be 100% fixed with the enhanced prompt.
+However, for some security weaknesses, the repair performance of Copilot Chat is considerably poor as it fails to aviod certain CWEs.
+For example, for CWE-94: Code Injection, Copilot Chat’s fixes were successful in less than 20% of cases, and Copilot Chat was unable to fix CWE-78: OS Command Injection.
+In addition, we also observe that as the information provided in the prompts increases, the percentages of certain CWEs being fixed improve, such as CWE-79: Cross-site Scripting and CWE-330: Use of Insufficiently Random Values Weakness.
+The percentage of successful fixes for CWE-79: Cross-site Scripting increases from 0% to 70%, while the percentage of successful fixes for CWE-330: Use of Insufficiently Random Values Weakness increases from less than 10% to 98%.
 5. Discussion
-In this section, we explain the study results in Section 5.1 and then discuss their implications in Section 5.2.
 5.1. Interpretation of Results
-RQ1: How secure is the code generated by Copilot?
-Among the 452 code snippets generated by Copilot, we found that 29.6% of these code snippets contain security weaknesses.
-Although in our results, the proportion of security issues in Python code snippets is higher than in JavaScript, there is no significant difference.
-Using Copilot to write code in Python or JavaScript can generally lead to security issues.
+RQ1: How secure is the code generated by Copilot in GitHub projects?
+Around quarter of the code snippets in our dataset contained security weaknesses.
+We found that the code with security weaknesses in Python primarily involves Utility Tools and Game applications, whereas, in JavaScript, Web Application code constitutes the majority of security weaknesses followed by Utility Tools.
+This disparity may arise from the two languages’ differing purposes and design philosophies (Reichenbach, 2022; Halfnine, 2024).
+JavaScript is mainly used for front-end development and Web applications (PubNub, 2024), leading to its security weaknesses being concentrated in scenarios related to the web and user interaction.
+In contrast, Python is more broadly applied in fields such as data processing and scientific computing (Nagpal and Gabrani, 2019; Simplilearn, 2024), resulting in security weaknesses occurring across a wider range of application domains.
+Although our results show that the proportion of security weaknesses in Python code snippets is higher than in JavaScript, there is no significant difference between the two languages.
+Using AI code generation tools to write code in Python or JavaScript can generally lead to security weaknesses.
 This could be attributed to features that made their code more flexible, such as dynamic typing and interpretation.
-From a Copilot point of view, the generated code does not need to reason about the whole program because it is dynamically typed, consequently, it does not need to read the flow of the whole program to generate working suggestions.
-In summary, developers should pay special attention to the security of Copilot-generated code, taking appropriate measures to validate input data and manage resources effectively to minimize security risks.
-The results of RQ1 suggest that in practical production, Copilot can help developers write code faster and increase productivity (GitHub, 2024; Perry et al., 2022).
-Still, additional security assessments and fixes are also required to ensure that the generated code does not introduce potential security risks.
-RQ2: What security weaknesses are present in the code snippets generated by GitHub Copilot?
-After conducting a security evaluation of 452 code snippets generated by Copilot, a total of 544 security weaknesses were identified, involving 38 CWEs, which is around 10% of the CWEs (439 CWEs) (The MITRE Corporation, 2023c).
-This may be because Copilot generates code in different programming languages and application scenarios.
-In addition, since the Copilot base model (Codex) is trained on publicly available data that potentially contain various types of security weaknesses, this can lead to the presence of multiple CWEs in the generated code by Copilot.
-This set of 38 CWEs covers many security issues, and Table 8 shows the types of security issues to which these CWEs are relevant.
-Table 8. CWEs and security issue types in Copilot-generated code
-Type of Security Issue Relevant CWEs
-Data Neutralization Issues CWE-94, CWE-78, CWE-117, CWE-89
-Resource Management Errors CWE-772, CWE-502
-Error Conditions, Return Values CWE-396, CWE-617
-Bad Coding Practices CWE-628, CWe-563
-Credentials Management Errors CWE-798
-Information Management Errors CWE-312, CWE-215
-Authentication Errors CWE-295
-Concurrency Issues CWE-367
-File Handling Issues CWE-22
+From a tool point of view, the generated code does not need to reason about the whole program because it is dynamically typed, and consequently, it does not need to read the flow of the whole program to generate working suggestions.
+In summary, developers should pay careful attention to the potential security weaknesses present in automatically generated code, taking appropriate measures to validate input data and manage resources effectively to minimize security risks.
+RQ1 results suggest that using Copilot is prone to security issues in production and additional security assessments are required to ensure the generated code does not introduce potential security risks.
+RQ2: What security weaknesses are present in the code generated by Copilot across different application domains?
+After conducting a security analysis of 733 code snippets generated by Copilot and other tools, a total of 628 security weaknesses were identified, involving 43 CWEs, which is around 10% of the CWEs (439 CWEs) (The MITRE Corporation, 2023).
+This may be because Copilot and the other two tools generate code in different programming languages and application scenarios.
+In addition, since the base models of AI code generation tools (such as Codex) are trained on publicly available data that potentially contain various types of security weaknesses, this can lead to the presence of multiple CWEs in the generated code.
+This set of 43 CWEs covers many security weaknesses.
 The diversity of security weaknesses indicates that developers using Copilot face various security risks.
-These risks are diverse, covering different development environments and application scenarios.
-In addition, we find that CWEs in Python are mainly related to data processing and system calls.
-In contrast, CWEs in JavaScript are often associated with dynamic code generation problems and security issues in web development.
-This may be because the two languages are designed differently.
-JavaScript is more commonly used for web development, while Python is more widely used in areas such as data processing and scientific computing.
-Overall, while Python and JavaScript differ in some common types of security weaknesses, they require developers to be aware of and take timely and targeted security measures to mitigate these risks.
-For example, developers should perform adequate validation of user inputs.
-It is also necessary to restrict the program’s permissions so that they only access essential resources.
-The results of RQ2 reveal the security weaknesses that developers may encounter in an actual production environment and their frequency of occurrence, which can help developers to be aware of the security aspects of the code generated by Copilot and to take appropriate measures to address the security weaknesses in an informed manner.
-RQ3: How many security weaknesses belong to the CWE Top-25?
-As shown in Table 7, eight of the CWEs in the Copilot-generated code can be found in the 2023 CWE Top-25 list, covering more than 218 security issues (40.1% of 544 identified CWEs) in our dataset.
-This indicates that the commonly acknowledged CWE Top-25 weaknesses in software development, which are considered the most prevalent and dangerous, are also prevalent in the code generated by Copilot.
-Therefore, developers using Copilot must pay close attention to these weaknesses and take appropriate measures to prevent them before they are integrated into their code.
-Meanwhile, we found that some higher-ranked CWEs in the CWE Top-25 list, such as CWE-89: SQL Injection, did not frequently appear in our scan results.
-This could be because the generated code does not establish connections with external resources (e.g., databases).
-Besides, we also observed that some vulnerabilities from the CWE Top-25 list were not detected in our scans, indicating that Copilot may sanitize and prevent specific weaknesses from being suggested to developers.
-GitHub is gradually improving the security of Copilot and its underlying model (Codex) (GitHub, 2023).
-However, it has also been found that Copilot’s security layer can potentially handle some CWEs better than others, leaving applications vulnerable to some critical vulnerabilities (Majdinasab et al., 2024).
-We also identified 30 security weaknesses in the code that do not belong to the CWE Top-25 list.
-Although these less common security weaknesses may not be as widespread as CWE Top-25, attackers can still exploit them.
+These risks are diverse, covering different development environments and scenarios.
+At the same time, eight of the CWEs identified in our dataset can be found in the 2023 CWE Top-25 list, covering more than 200 security weaknesses (37.1% of 628 identified).
+This indicates that the commonly acknowledged Top-25 CWEs, which are considered the most prevalent and dangerous weaknesses, are also prevalent in the AI-generated code.
+Therefore, developers using Copilot must pay close attention to these weaknesses and take appropriate measures to prevent them before they are integrated into their code base.
+Meanwhile, we observed that some vulnerabilities from the CWE Top-25 list were not detected in our analysis, indicating that Copilot may sanitize and prevent specific weaknesses from being suggested to developers.
+We also identified 30 security weaknesses in the generated code that do not belong to the CWE Top-25 list.
+Although these are less common security weaknesses and may not be as widespread as CWE Top-25, attackers can still exploit them.
 For example, we only detected one instance of CWE-732: Incorrect Permission Assignment for Critical Resource in our dataset.
@@
 However, it can lead to significant security risks when it does occur.
-Developers should also be aware of these less common security weaknesses to fully protect their code from attacks.
+Besides, we observed that the types of CWEs in generated code are closely related to the application domains.
+Code snippets used in Game development are prone to issues such as CWE-330: Use of Insufficiently Random Values Weakness and CWE-284: Improper Access Control.
+This may be due to the complex logic and player inputs typically involved in game applications, which can lead to security problems related to memory management and input validation.
+In the Utility Tool category, the proportion of security weaknesses is relatively high, which may be attributed to the diversity of user inputs.
+Furthermore, we found that Python code snippets often exhibit security weaknesses such as CWE-330: Use of Insufficiently Random Values Weakness, CWE-427: Uncontrolled Search Path Element, and CWE-78: OS Command Injection, whereas JavaScript code snippets primarily encounter issues like CWE-22: Path Traversal and CWE-94: Code Injection.
+This may be related to the differences in the application characteristics of the two languages: JavaScript, particularly through environments like Node.js, is use for file operations and dynamic code execution (Chhetri, 2016), while Python tends to focus more on system-level operations and automation scripts.
+In addition, the measures we need to take are different: Utility Tools in JavaScript need to emphasize path validation and the security of dynamic execution to prevent attacks such as path traversal and code injection, while those in Python should focus on ensuring safe library loading paths and validating user input for command execution to prevent system-level vulnerabilities caused by malicious input.
+In Web development, Python is commonly used for frequent interactions between the back-end and the database (Simplilearn, 2024), which makes it prone to CWE-89: SQL Injection.
+CWE-89: SQL Injection typically occurs when user input is directly embedded into SQL query statements without parameterized queries or input validation.
+On the other hand, JavaScript, as a front-end language that interacts directly with users, often faces security weaknesses such as CWE-79: Cross-site Scripting, CWE-94: Code Injection, and CWE-22: Path Traversal.
+This may be attributed to the frequent interactions between users and data in Web development.
+Additionally, the complexity of dependencies and the handling of sensitive data further increase the likelihood of security weaknesses.
+Regarding the Network Communication category, the proportion of security weaknesses is relatively lower than other categories.
+This may be attributed to the fact that code in the Network Communication domain typically demands a high level of security, benefiting from mature protocols, stringent industry standards.
+Additionally, the protocols involved in network communication inherently incorporate certain security mechanisms to safeguard the data transmission process, further mitigating common attack risks.
+Overall, while Python and JavaScript differ in some common types of security weaknesses, they require developers to be aware of and take timely and targeted security measures to mitigate these risks.
+RQ2 findings note the security weaknesses that developers may encounter in an actual production environment and their occurrence frequency.
+These findings can help developers be aware of the security aspects of the code generated by AI code generation tools and take appropriate measures to address the security weaknesses in an informed manner.
+RQ3: Can Copilot Chat help fix the security weaknesses found in the code generated by Copilot?
+As shown in Table LABEL:table:fix, we can see that regardless of the prompts used, Copilot Chat can fix a certain proportion of security weaknesses in the code we analyzed.
+Even the slash command /fix can resolve nearly 20% of security problems, indicating that Copilot Chat has some capability to fix certain weaknesses.
+Copilot Chat is an enhanced feature that extends the capabilities of the original Copilot (GitHub, 2023).
+In fact, the security of Copilot and its underlying model (Codex) is continually improving (GitHub, 2023).
+However, it has also been found that Copilot’s security layer can potentially handle some CWEs better than others, leaving applications vulnerable to some critical vulnerabilities (Majdinasab et al., 2024).
+It is believed that similar programming-based chatbots, such as Amazon Q (which works with CodeWhisperer), also offer similar repair capabilities.
+Furthermore, we can observe that using prompts with richer information helps improve the repair effectiveness, which guides us to leverage static analysis tools in combination with Copilot Chat to enhance the security of automatically generated code.
+When using the fix command, security weaknesses in JavaScript code are resolved more frequently than those in Python, possibly because JavaScript code is easier to trace in context, allowing Copilot Chat to more readily identify and fix these issues.
+In contrast, Python often involves data structures, libraries, and system calls, which may make security weakness resolution less intuitive compared to JavaScript.
+However, when enhanced information is provided, the repair rates for security weaknesses in Python and JavaScript code segments do not differ significantly.
+From the heatmap in Figure LABEL:fig:cwe-fix, we can see that the effectiveness of Copilot Chat in fixing different CWEs varies.
+Copilot Chat is particularly effective at fixing CWE-259: Use of Hard-coded Password and CWE-685: Function Call with Incorrect Number of Arguments, which pertain to access control issues and environment configuration.
+These issues are relatively straightforward and have standard repair methods, making them easier for automated tools to identify and fix.
+On the other hand, other weaknesses such as CWE-78: command injection and CWE-94: code injection involve executing commands or code provided by users, which typically require better complex contextual understanding and detailed input validation.
+Due to the complexity of these issues, Copilot Chat may encounter difficulties in automatically providing a fix.
+Additionally, we found that using more detailed instructions significantly improves repair outcomes for certain CWEs, such as CWE-330: Use of Insufficiently Random Values Weakness and CWE-79: Cross-site Scripting.
+This may be because our designed prompts provide richer information and context (the warning messages from static analysis tools), reducing ambiguity.
+This allows Copilot to effectively identify issues and generate correct fixes.
 5.2. Implications
-Importance of continuous security analysis of Copilot-generated code.
-We conjecture practitioners using Copilot will likely encounter security weaknesses, regardless of the programming language used.
-Our study results reflect the inevitability of security weaknesses in Copilot-generated code, and there are a variety of security weaknesses in Copilot-generated code.
+5.2.1. Importance of continuous security analysis of automatically-generated code
+We conjecture that practitioners using Copilot and other AI code generation tools are likely to encounter security weaknesses in their generated code, regardless of the programming language used.
+Our study results reflect the inevitability of security weaknesses in automatically generated code.
 Practitioners must be aware of the diverse security scenarios in production and adopt multiple security prevention measures to address security risks before accepting vulnerable code suggestions.
-Maintaining a rigorous process of security checks concurrently with code generation can identify potential vulnerability risks and rectify the security issues in time.
+Some CWEs, such as CWE-78: OS Command Injection and CWE-94: Code Injection, involve direct control of code execution by user inputs and may lead to severe security risks.
+Developers should be particularly vigilant about these high-risk operations and manually review these parts of the code.
+Input validation is the foundation of code security.
+Many security weaknesses (such as CWE-22: Path Traversal) arise from inadequate input validation.
+When using AI code generation tools, developers should avoid directly using user input as the basis for code execution or file path operations and manually add input validation or adopt a more secure coding approach.
+Moreover, Copilot performs relatively well in handling permission control (such as CWE-259: Use of Hard-coded Password) and environment configuration issues.
+Nevertheless, developers should be familiar with best practices related to permissions and verify that the configurations generated by AI code generation tools meet the project’s security requirements.
+Maintaining a rigorous process of security analysis concurrently with code generation can help identify potential vulnerability issues and rectify the security weaknesses in time.
 Developers should follow the best practices for using code generation tools and always check the code suggestions generated by Copilot (or any code generation tools).
-For example, developers can establish a gated check-in build process for checking and preventing security issues when committing code generated by Copilot.
-Initially, we can turn to automated tools to continuously scan the Copilot-generated code for known security issues, such as CWEs.
-Recognizing that these automated tools may not always detect all security issues, especially newly emerging ones or those that are context-specific, it is necessary to conduct a subsequent manual assessment, including manual security code review for Copilot-generated code.
-By embracing this combined strategy for continuous security analysis, we can ensure a robust security shield for the code committing process with Copilot-generated code.
-Prevention of security issues in Copilot-generated code.
-According to our study results, we provide the following suggestions on how to prevent potential security issues in Copilot-generated code: (1) Targeted security countermeasures: Based on the frequency of related CWEs in Copilot-generated code, practitioners can proactively prevent and address security issues in a targeted manner.
-When using Copilot to generate Python or JavaScript code, developers should focus on dedicated security weaknesses (CWEs) in Copilot-generated code of different programming languages (see Table 6).
+For example, developers can establish a gated check-in build process to check and prevent security weaknesses when committing code generated by AI code generation tools.
+Initially, we can turn to automated tools to continuously scan the Copilot-generated code for known security weaknesses, such as CWEs.
+Then, Copilot Chat or other LLMs can be used to fix identified code with security weaknesses, followed by a code re-analysis with security analysis tools.
+At the same time, the severity of security weaknesses and the limitations of static analysis tools remind us that it is essential to conduct a subsequent manual assessment, including manual security code review for auto-generated code.
+By embracing this combined strategy for continuous security analysis, we can ensure a robust security shield for the code-committing process with AI-generated code.
+5.2.2. Prevention of security weaknesses in AI-generated code
+We offer the following suggestions on how to prevent potential security weaknesses in generated code:
+Targeted security countermeasures: Based on the frequency of related CWEs in our dataset, practitioners can proactively prevent and address security weaknesses in a targeted manner.
+Developers should focus on dedicated CWEs in AI-generated code of different programming languages when using Copilot to generate e.g., Python or JavaScript code.
 Furthermore, a recent study shows that security weaknesses appear in certain code suggestions, but not all (Majdinasab et al., 2024); consequently, developers should carefully select potentially more secure suggestions, with the assistance of tools, that do not expose the code to vulnerabilities.
-(2) Standardized security assessment: Common security weaknesses in software development are also prevalent in the code generated by Copilot.
-As a good practice, developers can use the CWE Top-25 list as a guide to understand which security weaknesses are most common and dangerous in the generated code and follow the mitigation measures for related CWEs provided by the MITRE (The MITRE Corporation, 2023b).
-Additionally, the CWE Top-25 provides a standardized approach for security assessment, and developers can also use it as a guide to perform security audits of the code generated by Copilot.
-(3) Enhancing prompt engineering to generate security code: The instruction tuning schemes in code generation not only impacts the utility of the code but also its security.
+Security weaknesses can vary across programming languages and application domains, and developers should pay special attention to these differences when using these automated code generation tools.
+For example, the security weaknesses path validation (such as CWE-22: Path Traversal) and dynamic execution (such as CWE-94: Code Injection) are more likely to occur in JavaScript code.
+At the same time, system calls and library loading in Python code are more likely to raise security concerns.
+Standardized security assessment: Common security weaknesses in software development are also prevalent in the code generated by AI code generation tools.
+As a good practice, developers can use the CWE Top-25 list as a guide to understand which security weaknesses are most common and dangerous in the generated code.
+Additionally, the CWE Top-25 provides a standardized approach for security assessment, and developers can also use it as a guide to perform security audits of the AI-generated code.
+Using a standardized remediation strategy for common security weaknesses (such as privilege control and environment configuration) can effectively reduce the likelihood of generating insecure code.
+Developers can refer to the fixes provided in the static analysis tool documentation (CodeQL, 2024) or the mitigation measures for related CWEs provided by the MITRE (MITRE, 2024).
+Enhancing prompt engineering to generate more secure code: The instruction tuning schemes in code generation not only impact the utility of the code but also its security.
 By combining the security fine-tuning with standard instruction tuning, joint optimization of security and utility can be promoted (He et al., 2024).
-We should incorporate security considerations from the initial stages of code generation.
-For specific security scenarios that are prone to CWEs, we can improve the security of code by enhancing prompt engineering, e.g., with the prompt patterns proposed in (White et al., 2023).
+It is recommended to incorporate security considerations from the initial stages of code generation.
+For specific security scenarios prone to CWEs, we can improve code security by enhancing prompt engineering, e.g., with the prompt patterns proposed in (White et al., 2023).
+Distinguishing the complexity of security weaknesses: Some security issues, such as access control and environment configuration, have standard solutions and are well suited for quick fixes using tools like Copilot Chat.
+However, for complex logic controls (such as logic processing related to multi-layered inputs), Copilot Chat’s repair capabilities may be limited, requiring developers’ review and design.
+When using Copilot Chat, developers should assess the complexity of issues to determine which code can be automatically fixed and which requires manual intervention.
+Static analysis tools combined with LLM-based chatbots for fixing security weaknesses: When using Copilot Chat to address security weaknesses, developers should provide more specific context and clearer instructions.
+More detailed information can significantly enhance Copilot Chat’s ability to fix security problems.
+Meanwhile, integrating static analysis tools allows for a more comprehensive identification of potential risks in the code.
+Developers can leverage static analysis tools to identify specific security risks and provide detailed instructions to Copilot Chat, enabling it to effectively fix the issues in Copilot-generated code.
 6. Threats to Validity
@@
 Construct Validity: This study has three threats to construct validity: (1) Using keyword-based search – We used a keyword-based search to collect relevant code snippets from GitHub.
-The results obtained through the keyword-based search may not cover all the code snippets generated by Copilot on GitHub.
+The results obtained through the keyword-based search may not cover all the code snippets generated by Copilot and other tools on GitHub.
 We tried to mitigate this threat by constantly and iteratively refining the keywords and using synonyms.
-(2) Manual data filtering – We manually screened the results obtained from the keyword-based search by analyzing the comments, tags, and other metadata of the code snippets to determine whether they were generated by Copilot.
+(2) Manual data filtering – We manually screened the results obtained from the keyword-based search by analyzing the comments, tags, and other metadata of the code snippets to determine whether they were generated by Copilot or other tools.
 Since this process was manually done, it may have been influenced by personal bias.
-In this regard, two authors conducted the pilot experiment independently to minimize the impact on the construct validity.
+In this regard, two authors conducted the experiment independently to minimize the impact on the construct validity.
 (3) Manual association of CWEs – We manually associated the warning messages reported by the static analysis tools with a particular CWE, which may introduce personal subjective bias, threatening the construct validity.
 We employed two measures to mitigate this threat.
-First, since the list of CWEs is a tree structure with interconnections between them, we first matched the warning messages to a higher-level CWE, and further checked whether we can match the warning messages to a lower-level CWE with more specific definition.
-Second, to mitigate the personal bias, two authors independently assigned each security issue description a CWE ID.
+First, since the list of CWEs is a tree structure with interconnections between them, we first matched the warning messages to a higher-level CWE and further checked whether we can match the warning messages to a lower-level CWE with a more specific definition.
+We refer to the description of the test suite and CWE coverage in the static analysis tool documentation.
+Second, to mitigate the personal bias, two authors independently assigned each security weakness description a CWE ID within a period of ten days.
 In case of disagreement, the two authors discussed it with the assessment by a third author (a security expert).
-External Validity: Our dataset consists of Copilot-generated code snippets collected from open-source projects on GitHub.
-During the filtering process, we excluded code that utilized Copilot to solve algorithmic problems, aiming to ensure that the collected data genuinely reflected real-world production environments.
+External Validity: Our dataset consists of the code snippets generated by Copilot and other code generation tools collected from open-source projects on GitHub.
+During the filtering process, we excluded code that utilized AI code generation tools to solve simple programming practice problems, aiming to ensure that the collected data genuinely reflected open source development on GitHub.
 Since the data from GitHub are not diversified enough, we had a higher number of code snippets originating from the Game projects.
@@
 Furthermore, we acknowledge the need to collect more diverse code snippets from different platforms to increase the generalizability of the results.
-We will consider adopting more diversified ways or platforms to collect code.
-Additionally, due to the limitations of static analysis tools themselves, these tools could not scan all CWEs, and there is a degree of false positives in the scan results (as the case with static analysis, in general, (Sui et al., 2020; Kang et al., 2022)).
+We will consider adopting more diversified ways or platforms to collect code.Additionally, due to the limitations of static analysis tools themselves, these tools could not scan all CWEs, and there is a degree of false positives in the analysis results (as the case with static analysis, in general, (Sui et al., 2020; Kang et al., 2022)).
 Although we used two widely used static analysis tools to increase the comprehensiveness of the scans and manually checked the results of the tool scans, the results may suffer from incompleteness.
@@
 Developers have widely used these automated tools.
-The querying mechanism of these tools ensures that the scan results remain consistent when used multiple times.
-In addition, we performed two rounds of scanning with two tools for security checks on each code snippet, intending to complement the results of one tool with the other.
+The querying mechanism of these tools ensures that the analysis results remain consistent when used multiple times.
+In addition, we performed two rounds of analyzing with two tools for security checks on each code snippet, intending to complement the results of one tool with the other.
 By implementing these measures, we believe that our research results are reliable and these threats to reliability are mitigated.
@@
 Automatic code generation and recommendation have been an active research area due to the advancement of AI and, specifically, LLMs.
-AI code generation tools, such as Copilot, can significantly improve developers’ development efficiency, but can also introduce vulnerabilities and security risks.
+AI code generation tools, such as Copilot, can significantly improve developers’ development efficiency but can also introduce vulnerabilities and security risks.
+Existing studies on the security of AI-generated code mainly focused on the security issues in the generated code using crafted scenarios, and potential security weaknesses of AI-generated code in practical scenarios in open source development environment (e.g., GitHub) have not been fully considered.
 In this paper, we present the results of an empirical study to analyze security weaknesses in Copilot-generated code found in public GitHub projects.
-We identified 452 code snippets generated by Copilot from GitHub projects and analyzed those snippets for security weaknesses using static analysis tools.
+We identified 733 code snippets generated by Copilot and other tools (i.e., CodeWhisperer and Codeium) from GitHub projects and analyzed those snippets for security weaknesses using static analysis tools.
+We also examined using Copilot Chat to fix security issues in Copilot-generated code.
 This study aims to help developers understand the security risks of weaknesses introduced in the code generated by Copilot (and potentially similar code generation tools).
-Our results show that (1) 29.8% of the 452 Copilot-generated code snippets contain security weaknesses.
-Developers have a high risk of raising security issues when using Copilot, regardless of the programming language, so security checks are necessary.
-(2) The detected security weaknesses are diverse in nature and are associated with 38 different CWEs.
-Developers face a variety of development scenarios and application environments in production and need the appropriate security awareness and skills.
-(3) Among these CWEs, eight appear in the MITRE CWE Top-25 list, and six belong to the Stubborn Weaknesses, demonstrating their severity.
+Our results show that (1) around 30% of the 733 generated code snippets contain security weaknesses.
+Developers have a high risk of raising security weaknesses when using Copilot or other AI code generation tools, regardless of the programming language, and therefore, applying appropriate security checks is necessary.
+(2) The detected security weaknesses are diverse in nature and are associated with 43 different CWEs.
+Developers encounter a variety of development scenarios and production environments, requiring appropriate security awareness and skills.
+(3) Among these CWEs, eight appear in the 2023 CWE Top-25 list, and six belong to the Stubborn Weaknesses, indicating high severity.
+(4) Copilot Chat can help fix many of the security issues in Copilot-generated code, but its success rate varies across different CWEs.
+(5) Providing Copilot Chat with a warning message from the static analysis tool leads to a better fix.
 In the future, we plan to: (1) collect additional code snippets from other open source repositories and industrial projects and code snippets generated by newer releases of Copilot; (2) analyze and summarize the application scenarios of these code snippets, studying how practitioners use Copilot and fix the issues in development; and (3) compare the results with other emerging Generative AI code generation tools such as CodeWhisperer, aiXcoder, and Code Llama.
-Data availability
-We have shared the link to our dataset in the reference (Fu et al., 2024).
 Acknowledgments